### 30 Mar 2026 v1

# MuonEq: Balancing Before Orthogonalization with Lightweight Equilibration

Abstract
Orthogonalized-update optimizers such as Muon improve training of matrix-valued parameters, but existing extensions mostly act either after orthogonalization by rescaling updates or before it with heavier whitening-based preconditioners. We introduce MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon in three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C). These variants rebalance the momentum matrix before finite-step Newton–Schulz using row/column squared-norm statistics and only $\mathcal{O}(m+n)$ auxiliary state. We show that finite-step orthogonalization is governed by input spectral properties, especially stable rank and condition number, and that row/column normalization is a zeroth-order whitening surrogate that removes marginal scale mismatch. For the hidden matrix weights targeted by MuonEq, the row-normalized variant R is the natural default and preserves the $\widetilde{\mathcal{O}}(T^{-1/4})$ stationarity guarantee of Muon-type methods. In LLaMA2 pretraining on C4, the default R variant consistently outperforms Muon on 130M and 350M models, yielding faster convergence and lower validation perplexity.
1 Introduction
As the computational and data costs of large language model pretraining grow, optimization increasingly determines sample efficiency, throughput, and scalability. Beyond coordinate-wise adaptive methods such as AdamW (Loshchilov and Hutter, 2019), Adan (Xie et al., 2024), and MARS (Liu et al., 2025b), recent work has shifted toward a geometric view of matrix-valued updates. Muon (Jordan et al., ), for instance, uses a Newton–Schulz approximation to the polar factor of the momentum matrix to produce spectrally aligned updates (Bernstein and Newhouse, 2024; Chen et al., 2025). It has shown strong empirical scalability in LLM pretraining, often improving the trade-off among data efficiency, compute, and wall-clock time at fixed target loss (Liu et al., 2025a; Shah et al., 2025). Recent theory further characterizes Muon as spectral-norm steepest descent and establishes nonconvex convergence guarantees for both exact and approximate orthogonalization (Chang et al., 2025; Sato et al., 2025; Li and Hong, 2025; Shen et al., 2025; Sfyraki and Wang, 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025).
Follow-up work around Muon mainly proceeds along two directions. (i) Post-orthogonalization rescaling augments an already constructed orthogonal update with additional magnitude correction. Muon+ (Zhang et al., 2026b), NorMuon (Li et al., 2025), and AdaMuon (Si et al., 2025) respectively add row/column normalization, neuron-wise normalization, or element-wise variance adaptation after orthogonalization. These variants show that orthogonal updates remain compatible with later adaptive scaling, but they do not change the matrix that finite-step Newton–Schulz actually sees. (ii) Pre-orthogonalization geometric correction instead modifies the coordinate system before spectral optimization. SOAP (Vyas et al., 2024) connects Shampoo (Gupta et al., 2018) with Adam/Adafactor (Shazeer and Stern, 2018; Kingma and Ba, 2015) through a preconditioned basis, while FISMO (Xu et al., 2026a) and Mousse (Zhang et al., 2026c) further perform spectral optimization in whitened coordinates. These methods suggest that improving geometry before orthogonalization is the more direct lever, but doing so usually requires maintaining left and right preconditioners together with extra matrix operations such as eigendecomposition, inverse power iteration, and whitening/dewhitening maps.
These works motivate a more precise question:
Is there a lightweight design between post-orthogonalization rescaling and full whitening-based preconditioning that directly improves the input geometry for orthogonalization?
Our starting point is that, in Muon, the relevant object is not only the exact polar factor but also the finite-step Newton–Schulz approximation used in practice. Section 3.1 shows that both the convergence speed of Newton–Schulz iterations and the gap between finite-step iterations and exact orthogonalization are governed by input spectral geometry, especially stable rank and condition number. This points to a simple design principle: improve the geometry before orthogonalization, but do so with the smallest possible intervention. A natural formalization is the optimal diagonal preconditioning problem:
$\min_{\mathbf{D}_{1},\mathbf{D}_{2}\succ 0}\kappa\bigl(\mathbf{D}_{1}^{1/2}\mathbf{A}\mathbf{D}_{2}^{-1/2}\bigr).$
Qu et al. (Qu et al., 2025) derive an optimal diagonal preconditioner, but applying it online typically requires an auxiliary iterative solve. Classical row/column equilibration is less exact yet much lighter-weight (Ruiz, 2001); moreover, recent operator-geometry results identify row and column normalization as steepest-descent directions under the $p\to\infty$ and $1\to q$ geometries, respectively (Xu et al., 2026b). This makes diagonal equilibration a natural lightweight way to improve the matrix presented to Muon-style orthogonalization.
Based on this observation, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon. MuonEq has three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C). All three act directly on the momentum matrix before Newton–Schulz using only row/column squared-norm statistics and $\mathcal{O}(m+n)$ auxiliary state. Detailed justification is deferred to Section 3: Theorem 3.1 quantifies the spectral sensitivity of finite-step orthogonalization, Proposition 3.2 shows that row/column normalization is a zeroth-order surrogate for whitening, and Proposition 3.4 together with Theorem 3.5 clarify the trade-off between stronger two-sided equilibration and cleaner one-sided geometry. For the hidden Transformer weight matrices targeted by MuonEq, recent architecture-aware analyses by (Xu et al., 2026b) identify row-sided geometry as the natural one-sided choice, whereas column-sided geometry is more closely associated with embedding matrices; accordingly, we adopt R as the default and treat RC and C as companion forms for analysis and ablation. Compared with Muon+ (Zhang et al., 2026b), which rescales the update after orthogonalization, MuonEq modifies the matrix fed into orthogonalization. Unlike whitening-based methods such as FISMO and Mousse (Xu et al., 2026a; Zhang et al., 2026c), it avoids Kronecker factors, matrix-valued second-order statistics, eigendecomposition, and momentum transport, and requires only $\mathcal{O}(m+n)$ auxiliary state.
Our contributions are threefold.
  1. 1.
    First, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon with three forms: RC, R, and C; these variants rebalance the matrix before Newton–Schulz using only row/column squared-norm statistics, with R as the default for hidden matrix weights.
  2. 2.
    Second, we show that finite-step Newton–Schulz is governed by the input spectrum, especially stable rank and condition number; that row/column normalization acts as a zeroth-order surrogate for whitening; and that RC and R admit a clear trade-off between stronger spectral correction and cleaner one-sided convergence structure, with the default R variant satisfying a $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex guarantee under standard assumptions.
  3. 3.
    Third, empirically, the default R variant consistently outperforms Muon across model sizes and token budgets, while ablations over RC, R, and C show that R is the strongest default form for the hidden-weight setting targeted by MuonEq.
2 Method
Notation and setup. Scalars are denoted by non-bold letters, vectors by bold lowercase letters, and matrices by bold uppercase letters. On $\mathbb{R}^{d}$, we use the Euclidean inner product $\langle\mathbf{x},\mathbf{y}\rangle_{2}:=\mathbf{x}^{\top}\mathbf{y}$ and norm $\|\mathbf{x}\|_{2}$. For matrices, $\langle\mathbf{A},\mathbf{B}\rangle_{F}:=\operatorname{tr}(\mathbf{A}^{\top}\mathbf{B})$ and $\|\mathbf{A}\|_{F}$ denote the Frobenius inner product and norm, and $\|\mathbf{A}\|_{*}$ denotes the nuclear norm. We write $m]:=\{1,2,\ldots,m\}$ and let $\mathbb{N}$ be the set of non-negative integers. The model parameter is a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, and, unless otherwise stated, we assume $m\geq n$. For any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ with rank $r\leq n$, we write its compact singular value decomposition as $\mathbf{A}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})$ and $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$, and define $\operatorname{Orth}(\mathbf{A}):=\mathbf{U}\mathbf{V}^{\top}$. In finite-data form, the training objective can be written as $f(\mathbf{X}):=\frac{1}{N}\sum_{i\in[N]}f_{i}(\mathbf{X})$, where $f_{i}(\mathbf{X})$ is the loss associated with the $i$-th sample $\mathbf{z}_{i}$. Equivalently, we use the stochastic formulation
$\min_{\mathbf{X}\in\mathbb{R}^{m\times n}}f(\mathbf{X}),\qquad f(\mathbf{X})=\mathbb{E}_{\xi\sim\mathcal{D}}[f(\mathbf{X};\xi)],$ (1)
where $f$ may be nonconvex, $\xi$ is independent of $\mathbf{X}$, and $\mathbb{E}_{\xi}[\cdot]$ denotes expectation with respect to $\xi$. For any matrix $\mathbf{A}$ with singular values $\{\sigma_{i}(\mathbf{A})\}_{i=1}^{r}$, we further use $\operatorname{sr}(\mathbf{A}):=\frac{\|\mathbf{A}\|_{F}^{2}}{\|\mathbf{A}\|_{2}^{2}},\kappa_{2}(\mathbf{A}):=\frac{\sigma_{1}(\mathbf{A})}{\sigma_{r}(\mathbf{A})},p_{i}:=\frac{\sigma_{i}(\mathbf{A})^{2}}{\|\mathbf{A}\|_{F}^{2}},H(p):=-\sum_{i=1}^{r}p_{i}\log p_{i}.$
Muon as finite-step orthogonalized momentum. For matrix parameters, Muon constructs the update by orthogonalizing the momentum matrix. Let $\mathbf{M}_{t}$ denote the momentum at iteration $t$. The abstract Muon update is
$\mathbf{U}_{t}=\operatorname{Orth}(\tilde{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}\mathbf{U}_{t},$ (2)
where $\tilde{\mathbf{M}}_{t}$ is either the standard momentum or its Nesterov variant (Jordan et al., ; Liu et al., 2025a; Chang et al., 2025). In practice, $\operatorname{Orth}(\tilde{\mathbf{M}}_{t})$ is implemented by a small fixed number of Newton–Schulz iterations rather than exact polar decomposition (Jordan et al., ). The design question is therefore not only what direction is desirable in the exact limit, but also what matrix should be fed into finite-step Newton–Schulz.
Pre-orthogonalization equilibration family. MuonEq inserts diagonal preconditioning before orthogonalization:
$\hat{\mathbf{M}}_{t}=\textbf{DiagPre}(\tilde{\mathbf{M}}_{t};s),\qquad\mathbf{U}_{t}=\operatorname{Orth}(\hat{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}\mathbf{U}_{t},$ (3)
where the mode $s\in\{\mathrm{RC},\mathrm{R},\mathrm{C}\}$ selects one of three equilibration forms. Let
$\mathbf{D}_{r,t}=\operatorname{diag}\bigl(\operatorname{rowsum}(\tilde{\mathbf{M}}_{t}\odot\tilde{\mathbf{M}}_{t})+\varepsilon\bigr),\qquad\mathbf{D}_{c,t}=\operatorname{diag}\bigl(\operatorname{colsum}(\tilde{\mathbf{M}}_{t}\odot\tilde{\mathbf{M}}_{t})+\varepsilon\bigr).$
We consider the following three maps:
$\textbf{DiagPre}(\tilde{\mathbf{M}}_{t};s)=\begin{cases}\mathbf{D}_{r,t}^{-1/2}\tilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{RC},\\ \mathbf{D}_{r,t}^{-1/2}\tilde{\mathbf{M}}_{t},&s=\mathrm{R},\\ \tilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{C}.\end{cases}$ (4)
Since these maps use only row and column squared norms, the extra state is $\mathcal{O}(m+n)$, in the factorized spirit of Adafactor (Shazeer and Stern, 2018). Unlike methods that rely on matrix-valued or adaptive preconditioners (Gupta et al., 2018; Shazeer and Stern, 2018; Kingma and Ba, 2015; Xu et al., 2026a), MuonEq changes only the matrix fed to Newton–Schulz. Relative to Muon (Jordan et al., ), the orthogonalization routine itself is unchanged; relative to Muon+, NorMuon, and AdaMuon (Zhang et al., 2026b; Li et al., 2025; Si et al., 2025), MuonEq rebalances the input rather than rescales the output.
Why the default variant is R. This choice is block-dependent rather than a universal preference for row normalization. The trade-off is guided by
$\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}\leq$ $\underbrace{\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(S(\mathbf{M}))\right\|_{F}}_{\text{finite-step approximation error}}$ (5)
$+\underbrace{\left\|\mathrm{Orth}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}}_{\text{preconditioning bias}},$
where $S$ is one of RC, R, or C. The first term is the finite-step Newton–Schulz error and depends on the spectrum of $S(\mathbf{M})$; Theorem 3.1 therefore explains why the more aggressive two-sided RC map often gives the largest spectral correction. The second term is the bias induced by preprocessing. Proposition 3.2 shows that row/column normalization is a zeroth-order surrogate for whitening, removing marginal scale mismatch before orthogonalization. Among the one-sided variants, recent operator-geometry analyses identify row-sided geometry as the natural choice for the hidden Transformer weight matrices targeted by MuonEq, whereas column-sided geometry is more closely associated with embedding matrices (Xu et al., 2026b). This makes R the most appropriate default in our setting. Consistent with this picture, Proposition 3.4 characterizes the pure RC regime as a stronger two-sided spectral corrector, whereas Theorem 3.5 gives the cleanest nonconvex guarantee for the row-normalized variant.
[…]
Remark 2.1.
Unless otherwise stated, we use the row-normalized variant $s=\mathrm{R}$. The RC and C modes are kept as equally lightweight companion forms for analysis and ablation.
3 Analysis
The analysis follows the logic of MuonEq: Section 3.1 establishes the spectrum sensitivity of finite-step Newton–Schulz; Section 3.2 interprets row/column normalization as a zeroth-order surrogate for whitening; and Section 3.3 motivates the R rather than RC, with row normalization admitting a clean stochastic nonconvex guarantee. Together, Theorem 3.1, Proposition 3.2, Proposition 3.4, and Theorem 3.5 justify the MuonEq family and the default choice of R.
3.1 Why spectral geometry matters for finite-step Newton–Schulz orthogonalization
For clarity, the theorem below is stated for the wide case $p\leq q$; the tall case follows by transposition.
Theorem 3.1.
Let $\mathbf{G}\in\mathbb{R}^{p\times q}$ have rank $r\geq 1$, with compact SVD $\mathbf{G}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, and suppose $p\leq q$. Fix $\alpha\geq\|\mathbf{G}\|_{2}$, set $\mathbf{X}_{0}=\alpha^{-1}\mathbf{G}$, and define $\mathbf{X}_{k+1}=\left(a\mathbf{I}_{p}+b\mathbf{X}_{k}\mathbf{X}_{k}^{\top}+c(\mathbf{X}_{k}\mathbf{X}_{k}^{\top})^{2}\right)\mathbf{X}_{k}$ for $k\in\mathbb{N}$, where $a=3.4445$, b=4.7750b=-4.7750, and $c=2.0315$ are the coefficients used in Muon’s Newton–Schulz iteration in (Jordan et al., ). Let $\phi(s)=as+bs^{3}+cs^{5}$. Then, for each $k\in\mathbb{N}$, $\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}(s_{1}^{(k)},\ldots,s_{r}^{(k)})\mathbf{V}^{\top},s_{i}^{(0)}=\sigma_{i}/\alpha,s_{i}^{(k+1)}=\phi\bigl(s_{i}^{(k)}\bigr),i\in[r],$ and hence
$\frac{1}{\sqrt{r}}\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\bigl(1-a^{k}\sigma_{i}/\alpha\bigr)_{+}^{2}\right)^{1/2}.$
Here $(x)_{+}:=\max\{x,0\}$.
See Appendix A for details.
Remark 3.1.
Under the normalization $\alpha=\|\mathbf{G}\|_{F}$, which is common in practical Newton–Schulz implementations (Jordan et al., ; Liu et al., 2025a), let $p_{i}:=\frac{\sigma_{i}^{2}}{\|\mathbf{G}\|_{F}^{2}},\tau_{i}:=\log_{a}\frac{\|\mathbf{G}\|_{F}}{\sigma_{i}}=\log_{a}\frac{1}{\sqrt{p_{i}}}.$ Then Theorem 3.1 implies
$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\bigl(1-a^{k-\tau_{i}}\bigr)_{+}^{2}\right)^{1/2}.$
Thus $\tau_{i}$ is precisely the iteration at which the $i$-th singular direction leaves the linear regime. In particular, $\tau_{1}=\log_{a}\sqrt{\operatorname{sr}(\mathbf{G})},\tau_{r}=\log_{a}\bigl(\kappa_{2}(\mathbf{G})\sqrt{\operatorname{sr}(\mathbf{G})}\bigr),\tau_{r}-\tau_{1}=\log_{a}\kappa_{2}(\mathbf{G}).$ Therefore, the stable rank determines when the first singular direction enters the nonlinear regime, whereas the condition number determines the width of the transition window over which the remaining singular directions follow.
Theorem 3.1 shows that broad spectra slow finite-step orthogonalization by delaying contraction along small singular directions. To isolate this effect, Figure 1 uses random Gaussian matrices with controlled shapes and spectral spreads. The top row shows the finite-step error in Theorem 3.1, and the bottom row shows the corresponding condition numbers before and after normalization.
To verify the same mechanism on real training trajectories, Figure 2 uses pre-NS momentum matrices from a 2.6B-token Muon run on LLaMA2-130M Touvron et al. (2023) trained on C4 (Raffel et al., 2020). The top row shows the module-wise median of the error in Theorem 3.1, and the bottom row shows the corresponding mean.
(a) $\times1024$
[…]
(d) $\times4096$
(e) $\times1024$
[…]
(h) $\times4096$
Figure 1: Random Gaussian matrices with controlled shapes and spectral spreads. Top: finite-step relative Frobenius error to the exact polar factor across Newton–Schulz steps. Bottom: raw and post-normalization condition numbers. Two-sided row/column normalization yields the smallest error and the most consistent spectral compression.
(a) Step 1%
(b) Step 10%
(c) Step 50%
(d) Step 100%
(e) Step 1%
(f) Step 10%
(g) Step 50%
(h) Step 100%
Figure 2: Finite-step orthogonalization error across Newton–Schulz steps at 1%, 10%, 50%, and 100% of training. Top: module-wise median; bottom: module-wise mean; shaded bands denote the 25%–75% range. Two-sided row/column normalization decays fastest.
3.2 Row/column normalization as a zeroth-order surrogate for whitening
The next result shows that normalization is not merely heuristic rescaling. It is the leading-order term of exact whitening/orthogonalization when the normalized Gram matrix is close to identity. Consequently, normalization removes marginal scale mismatch first and leaves only the correlation correction to the polar step.
Proposition 3.2.
The following first-order expansions hold.
(i) Column/right whitening. Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full column rank, and set $\mathbf{D}_{c}:=\operatorname{diag}(\|\mathbf{M}_{:,1}\|_{2},\ldots,\|\mathbf{M}_{:,q}\|_{2})\succ 0$, $\mathbf{N}_{c}:=\mathbf{M}\mathbf{D}_{c}^{-1}$, and $\mathbf{C}_{c}:=\mathbf{N}_{c}^{\top}\mathbf{N}_{c}-\mathbf{I}_{q}$. If $\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}$ is sufficiently small, then
$\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{M}(\mathbf{M}^{\top}\mathbf{M})^{-1/2}=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}\bigl(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}\bigr),$
where $\mathbf{L}_{c}$ is the unique solution of $\mathbf{D}_{c}\mathbf{L}_{c}+\mathbf{L}_{c}\mathbf{D}_{c}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$ In particular,
$\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}+\mathcal{O}\bigl(\|\mathbf{C}_{c}\|_{2}\bigr).$
(ii) Row/left whitening. Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full row rank, and set $\mathbf{D}_{r}:=\operatorname{diag}(\|\mathbf{M}_{1,:}\|_{2},\ldots,\|\mathbf{M}_{p,:}\|_{2})\succ 0$, $\mathbf{N}_{r}:=\mathbf{D}_{r}^{-1}\mathbf{M}$, and $\mathbf{C}_{r}:=\mathbf{N}_{r}\mathbf{N}_{r}^{\top}-\mathbf{I}_{p}$. If $\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}$ is sufficiently small, then
$\operatorname{Orth}_{\ell}(\mathbf{M})=(\mathbf{M}\mathbf{M}^{\top})^{-1/2}\mathbf{M}=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}\bigr),$
where $\mathbf{L}_{r}$ is the unique solution of $\mathbf{D}_{r}\mathbf{L}_{r}+\mathbf{L}_{r}\mathbf{D}_{r}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$ In particular,
$\operatorname{Orth}_{\ell}(\mathbf{M})=\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{C}_{r}\|_{2}\bigr).$
See Appendix C for details.
Corollary 3.3.
Under the assumptions of the row/left part of Proposition 3.2, if $\|\mathbf{C}_{r}\|_{2}$ is sufficiently small, then
$\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{p}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\frac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{C}_{r}\|_{2}^{2}\bigr),$
and hence
$\|\operatorname{Orth}_{\ell}(\mathbf{N}_{r})-\mathbf{N}_{r}\|_{2}\leq\frac{1}{2}\|\mathbf{N}_{r}\|_{2}\|\mathbf{C}_{r}\|_{2}+\mathcal{O}\bigl(\|\mathbf{C}_{r}\|_{2}^{2}\bigr).$
In particular, since $\operatorname{diag}(\mathbf{N}_{r}\mathbf{N}_{r}^{\top})=\mathbf{I}_{p}$, row normalization removes the marginal scale mismatch, so the leading whitening correction acts only on the residual off-diagonal Gram error.
Remark 3.2.
By contrast, the corresponding bound for $\operatorname{Orth}_{\ell}(\mathbf{M})-\mathbf{M}$ contains the additional zeroth-order term
$\|\mathbf{D}_{r}-\mathbf{I}_{p}\|_{2}\|\mathbf{N}_{r}\|_{2}.$
Thus row normalization separates marginal rescaling from the genuine whitening step. The column-normalized/right analogue and the corresponding two-sided normalization statements are completely symmetric; see Appendix D.
Together, Proposition 3.2, Corollary 3.3, and Remark 3.2 explain the division of labor in MuonEq: normalization removes marginal-scale mismatch at zeroth order, and the subsequent Newton–Schulz orthogonalization handles the remaining correlation structure. This is exactly why DiagPre Eq. (4) is a meaningful lightweight surrogate for a much more expensive whitening step.
3.3 Convergence analysis of the RC and R variants
To analyze the MuonEq family, we focus on the two variants that are most informative theoretically: the two-sided RC map and the default row-normalized R map. RC is the strongest spectral corrector among the three forms but induces a more involved alignment term, whereas R yields the cleanest one-sided geometry and convergence argument. Accordingly, Proposition 3.4 studies the pure RC variant, and Theorem 3.5 provides the main guarantee for the default R variant. The column-normalized C map is retained as the embedding-aligned companion in the method family and empirical ablations.
Assumption 3.1.
The objective is bounded below: there exists $f^{\star}>$-$\infty$ such that $f(\mathbf{X})\geq f^{\star}$ for all $\mathbf{X}\in\mathbb{R}^{m\times n}$.
Assumption 3.2.
The function $f$ is $L$-smooth, that is, $\|\nabla f(\mathbf{Y})-\nabla f(\mathbf{X})\|_{F}\leq L\|\mathbf{Y}-\mathbf{X}\|_{F}$ for all $\mathbf{X},\mathbf{Y}\in\mathbb{R}^{m\times n}$.
Assumption 3.3.
The stochastic gradient is unbiased with bounded variance. Specifically, there exists $\sigma>0$ such that, for all $\mathbf{X}\in\mathbb{R}^{m\times n}$, $\mathbb{E}[\nabla f(\mathbf{X};\xi)]=\nabla f(\mathbf{X}),\mathbb{E}\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}\leq\sigma^{2}.$
These assumptions are standard in analyses of adaptive and Muon-type methods (Xie et al., 2024; Chen et al., 2025; Li et al., 2023; Wang et al., 2023; Chang and Yuan, 2025; Huang et al., 2025).
Proposition 3.4.
Assume Assumptions 3.1, 3.2, and $3.$ Moreover, assume there exists $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty}$ almost surely. Consider the RC variant (4) with $\lambda=0$ and Nesterov disabled, so that $\hat{\mathbf{M}}_{t}=\mathbf{D}_{r,t}^{-1/2}\mathbf{M}_{t}\mathbf{D}_{c,t}^{-1/2}$ for all $t\in[T]$. Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, and $a:=0.2\sqrt{\max\{m,n\}}$. Suppose further that $\varepsilon\geq\frac{4}{5}G_{\infty}^{2}\max\{m,n\}$, and define $\rho_{r}:=\sqrt{1+\frac{nG_{\infty}^{2}}{\varepsilon}},\rho_{c}:=\sqrt{1+\frac{mG_{\infty}^{2}}{\varepsilon}},\chi_{\varepsilon}:=\frac{(\rho_{r}+1)(\rho_{c}+1)}{4\rho_{r}\rho_{c}}-\frac{3\rho_{r}\rho_{c}-\rho_{r}-\rho_{c}-1}{4}.$ Then $\rho_{r},\rho_{c}\leq 3/2$ and hence $\chi_{\varepsilon}\geq 1/144$. Consequently,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}(1+\ln T)+C_{2}}{a\chi_{\varepsilon}\cdot T^{1/4}},$
where $C_{1}=\frac{a}{L}\bigl(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}\bigr)+aL\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2},C_{2}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}n}{2}.$
Theorem 3.5.
Assume Assumptions 3.1, 3.2, and $3.$ Consider the practical row-normalized R variant (4), with Nesterov disabled and $\lambda=0,\varepsilon=0$. Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, $a:=0.2\sqrt{\max\{m,n\}}$, and suppose the trajectory-wise NS5 polar-approximation error $\varepsilon_{\mathrm{ns}}$ defined in Lemma G.2 satisfies $\varepsilon_{\mathrm{ns}}<1$. Then, for every $T\geq 1$,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}}{\bigl(a(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}\bigr)\cdot T^{1/4}},$
where $C_{1}^{\mathrm{ns}}=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)+aL\Bigl(\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)^{2},C_{2}^{\mathrm{ns}}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}.$
Remark 3.3.
Setting $\varepsilon_{\mathrm{ns}}=0$ recovers the exact-polar theorem for the row-normalized R variant. More generally, compared with the two-sided RC analysis, the descent coefficient is now the fixed positive constant $(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}$, so no auxiliary condition of the form $\chi_{\varepsilon}>0$ is needed.
Remark 3.4.
Equation (5) clarifies the difference between the RC and R variants. The first term is the finite-step orthogonalization error. By Theorem 3.1, it is controlled by the singular spectrum of $S(\mathbf{M})$, so the two-sided RC map can provide a stronger spectral correction. The second term is the bias introduced by preprocessing. Proposition 3.4 shows that the pure RC update preserves descent only under the auxiliary positivity condition $\chi_{\varepsilon}>0$. By contrast, the row-normalized map admits the fixed alignment lower bound
$\left\langle\mathbf{M},\operatorname{Orth}(\mathbf{P}(\mathbf{M})\mathbf{M})\right\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}\|_{F},\qquad\mathbf{P}(\mathbf{M}):=\operatorname{diag}\!\bigl(\operatorname{rowsum}(\mathbf{M}\odot\mathbf{M})\bigr)^{-1/2},$
which is exactly why Theorem 3.5 no longer needs a condition of the form $\chi_{\varepsilon}>0$. Thus RC should be viewed as the stronger two-sided spectral corrector, whereas R is the cleaner one-sided variant and therefore the default choice for the hidden-weight setting considered in this paper. The column-normalized C variant is retained as the embedding-aligned companion and is studied empirically rather than through a separate main theorem.
Remark 3.5.
Theorem 3.5 yields the sublinear stationarity estimate
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}=\mathcal{O}\!\left(\frac{1+\ln T}{T^{1/4}}\right).$
Hence, up to the logarithmic factor and the explicit problem-dependent constants, the default R variant has the same $T^{-1/4}$-type baseline dependence as the currently known convergence theory for standard stochastic nonconvex Muon-type methods (Chen et al., 2025; Chang et al., 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025). Improvements beyond this baseline typically require additional mechanisms such as variance reduction (Liu et al., 2025b; Chang et al., 2025; Huang et al., 2025; Qian et al., 2025; Fang et al., 2018; Zhou et al., 2020; Cutkosky and Orabona, 2019; Huang et al., 2021).
4 Experiments
4.1 Main Result
(a) Training Loss
(b) Validation Loss
(c) Wall-clock time
Figure 3: The training and validation loss curves, plotted against both training tokens and wall-clock time on LLaMA-2 130M.
(a) Training Loss
(b) Validation Loss
(c) Wall-clock time
Figure 4: The training and validation loss curves, plotted against both training tokens and wall-clock time on LLaMA-2 350M.
In this section, we evaluate the performance of the proposed MuonEq optimizers on LLM pretraining. All experiments are conducted using 64 Ascend 910C (64GB) NPUs. Detailed experimental settings are provided in Appendix J.
[…]
Figure 3 reports training loss versus tokens, validation loss versus tokens, and validation loss versus wall-clock time for LLaMA2-130M trained to 10.5B tokens. Figure 4 reports the corresponding validation-loss comparison for LLaMA2-350M trained to 21.0B tokens. Table 1 reports the best validation perplexity, peak memory (GB) usage, and training wall-clock time (Second/Step) for these main comparison runs. Across Figure 3, Figure 4, and Table 1, MuonEq consistently improves over Muon.
Table 1: Best validation perplexity, peak memory usage, and wall-clock training time on LLaMA2/C4.
Optimizer 130M 350M
Perplexity Memory Time Cost Perplexity Memory Time Cost
AdamW Loshchilov and Hutter (2019) 21.35 12.94 0.21 16.25 29.43 0.53
MARS-AdamW Liu et al. (2025b) 20.03 12.99 0.23 15.62 29.47 0.57
Muon Jordan et al. ; Liu et al. (2025a) 20.52 12.52 0.25 15.21 28.18 0.61
Muon-Nes Jordan et al. ; Liu et al. (2025a); Chang et al. (2025) 20.18 12.52 0.25 15.19 28.18 0.61
[…]
Training Tokens 10.5B 21.0B
4.2 Ablation Study
We ablate the three MuonEq variants RC, R, and C to isolate pre-orthogonalization geometry. All variants use the same optimizer states, scalar hyperparameters, and Newton–Schulz budget; only the pre-orthogonalization diagonal map changes. We take R as the default variant. All ablations run on 4 RTX Pro6000 (96GB) GPUs; details are deferred to Appendix J.2.
The results consistently favor R. Among the three static variants, R performs best on both CIFAR-10/ResNet-18 and FineWeb/GPT2-small, reaching 94.53% test accuracy on CIFAR-10 and 24.94 validation perplexity on FineWeb/GPT2-small. RC is competitive but falls short of R, while C is weaker in this hidden-weight setting. Full static-variant comparisons appear in Tables 3 and 3.
This pattern matches the trade-off in Eq. (5) and the analysis in Section 3. RC provides stronger two-sided spectral correction and can reduce the finite-step Newton–Schulz term more aggressively, but also introduces larger preconditioning bias. For the hidden matrix weights targeted by MuonEq, the more relevant one-sided geometry is row-sided, so R offers the best trade-off between spectral correction and geometric alignment. This interpretation is consistent with (Xu et al., 2026b). Comparisons with (Pethick et al., 2025) and (Glentis et al., 2025) require care because their row/column terminology depends on parameter layout and task setting; we discuss this point in the appendix J.3.
Table 2: Ablation on CIFAR-10 with ResNet-18.
Method Component Test Acc. (%)
Row Col
Baseline Experiments
SGD 92.84
AdamW 93.42
Muon-Nes 94.41
Ablation Experiments
MuonEq (RC) 94.23
MuonEq (C) 94.44
[…]
Table 3: Ablation on Fineweb-10.5B tokens with GPT2-small.
Method Component Validation Perplexity
Row Col
Baseline Experiments
AdamW 26.60
Muon-Nes 25.31
Ablation Experiments
MuonEq (RC) 24.98
MuonEq (C) 25.23
[…]
5 Related Work
MuonEq is viewed as a lightweight pre-balanced Muon family rather than as a fully whitened optimizer. It inserts diagonal equilibration before Newton–Schulz orthogonalization and includes three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C), with R as the default choice for the hidden matrix weights considered in this paper. All three use only row/column squared-norm statistics and preserve the $\mathcal{O}(m+n)$ state profile.
Muon and orthogonalized-update optimizers. Muon Jordan et al. introduced the core design of orthogonalizing matrix-valued momentum with a small fixed number of Newton–Schulz iterations. Subsequent work has expanded this idea into a broader family of orthogonalized-update optimizers, including post-orthogonalization normalization and rescaling methods such as Muon+, NorMuon, and AdaMuon Zhang et al. (2026b); Li et al. (2025); Si et al. (2025); adaptive stepsize variants such as AdaGO and NAMO Zhang et al. (2025, 2026a); distributed and communication-aware extensions such as Dion and MuLoCo Ahn et al. (2025); Thérien et al. (2025); reduced-state implementations Gupta et al. (2025); and fine-tuning-oriented variants Page et al. (2025). These methods mainly alter the orthogonalized update itself or the surrounding optimization machinery. By contrast, MuonEq modifies the matrix fed into orthogonalization through lightweight pre-orthogonalization equilibration.
Theory of Muon-style methods. Recent theory interprets Muon as steepest descent under spectral-norm or nuclear-norm geometry Bernstein and Newhouse (2024); Chen et al. (2025), establishes stochastic and nonconvex convergence guarantees Shen et al. (2025); Chang et al. (2025); Nagashima and Iiduka (2026); Zhang and Lin (2026), and analyzes practical ingredients such as weight decay, critical batch size, and inexact orthogonalization Sato et al. (2025); Gyu Yeol Kim (2026); Shulgin et al. (2025). This literature is directly relevant to MuonEq because we retain the same finite-step orthogonalization primitive as Muon and change only the geometry of its input. Our analysis further highlights the role of input spectrum in finite-step Newton–Schulz and clarifies how diagonal equilibration trades stronger spectral correction against preprocessing bias.
Structured preconditioning, factorized statistics, and whitening-inspired methods. A separate line of work studies matrix-aware preconditioning, including Shampoo Gupta et al. (2018), K-FAC Martens and Grosse (2015); Eschenhagen et al. (2023), SOAP Vyas et al. (2024), Sketchy Feinberg et al. (2023), and ASGO An et al. (2025). Adafactor Shazeer and Stern (2018) is especially relevant from a systems perspective because it shows how row- and column-factorized second-moment statistics reduce optimizer state to $\mathcal{O}(m+n)$ for matrix-valued parameters. Recent Muon-adjacent whitening or preconditioned-polar methods, including FISMO and Mousse Xu et al. (2026a); Zhang et al. (2026c), pursue a similar high-level goal: improve the geometry before imposing spectral constraints. The key difference is that these methods rely on richer matrix- or Kronecker-structured preconditioners, whereas MuonEq restricts the intervention to diagonal row/column equilibration. In this sense, MuonEq is closer to low-cost equilibration than to full whitening.
Orthogonalization algorithms. There is also a complementary numerical literature on faster and more accurate polar/orthogonalization routines, including improved Newton–Schulz variants and matrix-sign methods such as CANS and Polar Express Grishina et al. (2025); Amsel et al. (2025). This literature matters because Muon-style optimizers operate in the regime of approximate orthogonalization, where only a small number of iterations is affordable; MuonEq is designed explicitly for this setting.
6 Conclusion
We study how to improve the matrix geometry seen by finite-step orthogonalization in Muon-style optimization without the cost of full whitening. MuonEq does so with a lightweight family of pre-orthogonalization equilibration schemes: two-sided row/column normalization (RC), row normalization (R), and column normalization (C). These variants modify the input to Newton–Schulz orthogonalization while preserving the $\mathcal{O}(m+n)$ memory profile of factorized methods. For the hidden matrix weights studied here, we use R as the default variant.
The analysis shows that finite-step orthogonalization is highly sensitive to the input spectrum and that row/column normalization acts as a zeroth-order whitening surrogate by removing marginal scale mismatch. It also clarifies the trade-off within the MuonEq family: RC gives stronger two-sided spectral correction, whereas R yields cleaner one-sided geometry and the main $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex guarantee. Empirically, the default R variant consistently outperforms Muon in LLaMA2 pretraining on C4 across evaluated model sizes and token budgets. Ablations over RC, R, and C further support R as the strongest default for the hidden-weight setting considered here.
Appendix
Appendix A Proofs of Theorem 3.1
Proof.
The claim for $k=0$ is immediate from
$\mathbf{X}_{0}=\alpha^{-1}\mathbf{G}=\mathbf{U}\operatorname{diag}\Bigl(\frac{\sigma_{1}}{\alpha},\dots,\frac{\sigma_{r}}{\alpha}\Bigr)\mathbf{V}^{\top}.$
Assume that, for some $k\geq 0$,
$\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}\bigl(s_{1}^{(k)},\dots,s_{r}^{(k)}\bigr)\mathbf{V}^{\top}.$
Then
$\mathbf{X}_{k}\mathbf{X}_{k}^{\top}=\mathbf{U}\operatorname{diag}\bigl((s_{1}^{(k)})^{2},\dots,(s_{r}^{(k)})^{2}\bigr)\mathbf{U}^{\top}.$
Substituting this into the iteration and using $\mathbf{U}^{\top}\mathbf{U}=I_{r}$, we obtain
$\mathbf{X}_{k+1}=\mathbf{U}\operatorname{diag}\bigl(\phi(s_{1}^{(k)}),\dots,\phi(s_{r}^{(k)})\bigr)\mathbf{V}^{\top}.$
Hence, by induction,
$\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}\bigl(s_{1}^{(k)},\dots,s_{r}^{(k)}\bigr)\mathbf{V}^{\top},\qquad s_{i}^{(k+1)}=\phi\left(s_{i}^{(k)}\right).$
Moreover, we have
$\operatorname{Orth}(\mathbf{G})=\mathbf{U}\mathbf{V}^{\top}=\mathbf{U}I_{r}\mathbf{V}^{\top},\quad\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})=\mathbf{U}\Bigl(\operatorname{diag}(s_{1}^{(k)},\dots,s_{r}^{(k)})-I_{r}\Bigr)\mathbf{V}^{\top}.$
Since left and right multiplication by matrices with orthonormal columns preserves the Frobenius norm, it follows that
$\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}^{2}=\sum_{i=1}^{r}(s_{i}^{(k)}-1)^{2},$
which proves the exact error representation.
We next prove the lower bound. Define
$q(t):=a+bt+ct^{2},\qquad t\in[0,1].$
Then
$q^{\prime}(t)=b+2ct=-4.7750+4.0630t<0,\qquad t\in[0,1].$
Hence $q$ is strictly decreasing on $[0,1]$, and in particular
$0<q(1)=a+b+c=0.701\leq q(t)\leq q(0)=a,\qquad t\in[0,1].$
Fix any $i\in\{1,\dots,r\}$ and $k\geq 0$. If $a^{k}\sigma_{i}/\alpha\geq 1$, then trivially
$|1-s_{i}^{(k)}|\geq 0=\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}.$
Now suppose that $a^{k}\sigma_{i}/\alpha<1$. We claim that, for every $0\leq j\leq k$,
$0<s_{i}^{(j)}\leq a^{j}\frac{\sigma_{i}}{\alpha}<1.$
The case $j=0$ is immediate. Suppose the claim holds for some $j<k$. Then $s_{i}^{(j)}\in(0,1)$, so $(s_{i}^{(j)})^{2}\in(0,1)$ and thus
$0<q\bigl((s_{i}^{(j)})^{2}\bigr)\leq a.$
Therefore,
$0<s_{i}^{(j+1)}=s_{i}^{(j)}q\bigl((s_{i}^{(j)})^{2}\bigr)\leq as_{i}^{(j)}\leq a^{j+1}\frac{\sigma_{i}}{\alpha}.$
Since $j+1\leq k$, we further have
$a^{j+1}\frac{\sigma_{i}}{\alpha}\leq a^{k}\frac{\sigma_{i}}{\alpha}<1.$
This proves the claim. In particular,
$0<s_{i}^{(k)}\leq a^{k}\frac{\sigma_{i}}{\alpha}<1,$
and hence
$|1-s_{i}^{(k)}|=1-s_{i}^{(k)}\geq 1-a^{k}\frac{\sigma_{i}}{\alpha}=\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}.$
Combining the two cases, we conclude that, for every $i$,
$|1-s_{i}^{(k)}|\geq\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}.$
Squaring both sides, summing over $i$, and invoking the exact error representation established above yields
$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}^{2}\right)^{1/2}.$
This completes the proof. ∎
Remark A.1.
Since
$\phi(1)=a+b+c=0.701\neq 1,$
the scalar map induced by the above Muon coefficients does not have $1$ as a fixed point. Therefore, $T_{\varepsilon}$ in the theorem should be interpreted as the first hitting time of the error threshold $\varepsilon$, rather than as the number of iterations required for asymptotic convergence to $\operatorname{Orth}(\mathbf{G})$.
Appendix B Lemmas for Proposition 3.2
B.1 Lemma B.1
Lemma B.1.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be symmetric positive definite, and let $\mathbf{E}\in\mathbb{R}^{n\times n}$ be symmetric. Define $g(\mathbf{X})=\mathbf{X}^{1/2},f(\mathbf{X})=\mathbf{X}^{-1/2}$. Then $g$ and $f$ are Fréchet differentiable at $\mathbf{A}$. If $\mathbf{S}=\mathbf{A}^{1/2},\mathbf{L}=L_{g}(\mathbf{A};\mathbf{E})$, then $\mathbf{L}$ is the unique symmetric solution of the Sylvester equation
$\mathbf{S}\mathbf{L}+\mathbf{L}\mathbf{S}=\mathbf{E}.$
Moreover,
$L_{f}(\mathbf{A};\mathbf{E})=-\mathbf{A}^{-1/2}\mathbf{L}\mathbf{A}^{-1/2}.$
Consequently, for $\|\mathbf{E}\|_{2}$ sufficiently small,
$(\mathbf{A}+\mathbf{E})^{-1/2}=\mathbf{A}^{-1/2}-\mathbf{A}^{-1/2}\mathbf{L}\mathbf{A}^{-1/2}+\mathcal{O}(\|\mathbf{E}\|_{2}^{2}).$
Proof.
For completeness, differentiate the identity $g(\mathbf{X})^{2}=\mathbf{X}$ at $\mathbf{A}$ in the direction $\mathbf{E}$. Writing $\mathbf{L}=L_{g}(\mathbf{A};\mathbf{E})$ and $\mathbf{S}=\mathbf{A}^{1/2}$ gives
$\mathbf{S}\mathbf{L}+\mathbf{L}\mathbf{S}=\mathbf{E}.$
Since $\mathbf{S}\succ 0$, the Sylvester operator $\mathbf{X}\mapsto\mathbf{S}\mathbf{X}+\mathbf{X}\mathbf{S}$ is invertible, so the solution $\mathbf{L}$ is unique. Next, since $f=\mathrm{inv}\circ g$, the chain rule and the Fréchet derivative of the matrix inverse,
$L_{\mathrm{inv}}(\mathbf{Z};\mathbf{W})=-\mathbf{Z}^{-1}\mathbf{W}\mathbf{Z}^{-1},$
yield
$L_{f}(\mathbf{A};\mathbf{E})=-\mathbf{S}^{-1}\mathbf{L}\mathbf{S}^{-1}=-\mathbf{A}^{-1/2}\mathbf{L}\mathbf{A}^{-1/2}.$
The stated first-order expansion follows from the Fréchet differentiability of $f$ at $\mathbf{A}$. ∎
B.2 Lemma B.2
Lemma B.2.
Let $\mathbf{A}\succ 0$ and set $\mathbf{S}=\mathbf{A}^{1/2}\succ 0$. Consider the Sylvester equation
$\mathbf{SL}+\mathbf{LS}=\mathbf{E}.$
Its unique solution is given by [5]
$\mathbf{L}=\int_{0}^{\infty}e^{-t\mathbf{S}}\mathbf{E}e^{-t\mathbf{S}}dt.$
Moreover, in the operator norm,
$\|\mathbf{L}\|_{2}\leq\frac{1}{2\lambda_{\min}(\mathbf{S})}\|\mathbf{E}\|_{2}=\frac{1}{2\sqrt{\lambda_{\min}(\mathbf{A})}}\|\mathbf{E}\|_{2}.$
The same bound also holds for the Frobenius norm, and more generally for any unitarily invariant norm.
Proof.
Differentiate the integrand:
$\frac{d}{dt}\bigl(e^{-t\mathbf{S}}\mathbf{E}e^{-t\mathbf{S}}\bigr)=-\mathbf{S}e^{-t\mathbf{S}}\mathbf{E}e^{-t\mathbf{S}}-e^{-t\mathbf{S}}\mathbf{E}e^{-t\mathbf{S}}\mathbf{S}.$
Integrating from 0 to $\infty$ yields
$\mathbf{S}\mathbf{L}+\mathbf{L}\mathbf{S}=\mathbf{E},$
which proves the integral representation.
For the norm bound, using submultiplicativity and the fact that
$\|e^{-t\mathbf{S}}\|_{2}=e^{-t\lambda_{\min}(\mathbf{S})},$
we have
$\|\mathbf{L}\|_{2}\leq\int_{0}^{\infty}\|e^{-t\mathbf{S}}\|_{2}^{2}\|\mathbf{E}\|_{2}\,dt=\|\mathbf{E}\|_{2}\int_{0}^{\infty}e^{-2t\lambda_{\min}(\mathbf{S})}dt=\frac{1}{2\lambda_{\min}(\mathbf{S})}\|\mathbf{E}\|_{2}.$
Since $\lambda_{\min}(\mathbf{S})=\sqrt{\lambda_{\min}(\mathbf{A})}$, the stated bound follows. ∎
Appendix C Proofs of Proposition 3.2
Proof.
For the column/right statement, we have
$\mathbf{M}^{\top}\mathbf{M}=\mathbf{D}_{c}(\mathbf{I}_{n}+\mathbf{C}_{c})\mathbf{D}_{c}=\mathbf{D}_{c}^{2}+\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$
Apply Lemma B.1 with $\mathbf{A}=\mathbf{D}_{c}^{2},\mathbf{E}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}$. Then
$(\mathbf{M}^{\top}\mathbf{M})^{-1/2}=\mathbf{D}_{c}^{-1}-\mathbf{D}_{c}^{-1}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}),$
where $\mathbf{L}_{c}$ solves $\mathbf{D}_{c}\mathbf{L}_{c}+\mathbf{L}_{c}\mathbf{D}_{c}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$
Multiplying on the left by $\mathbf{M}$ yields
$\operatorname{Orth}_{r}(\mathbf{M})$ $=\mathbf{M}\mathbf{D}_{c}^{-1}-\mathbf{M}\mathbf{D}_{c}^{-1}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2})$
$=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}).$
Moreover, by Lemma B.2, we have
$\|\mathbf{L}_{c}\|_{2}\leq\frac{1}{2\lambda_{\min}(\mathbf{D}_{c})}\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}\leq\frac{\|\mathbf{D}_{c}\|_{2}^{2}}{2\lambda_{\min}(\mathbf{D}_{c})}\|\mathbf{C}_{c}\|_{2}.$
Hence
$\|\mathbf{L}_{c}\mathbf{D}_{c}^{-1}\|_{2}\leq\|\mathbf{L}_{c}\|_{2}\|\mathbf{D}_{c}^{-1}\|_{2}\leq\frac{\kappa(\mathbf{D}_{c})^{2}}{2}\|\mathbf{C}_{c}\|_{2}.$
Therefore the first-order correction is $\mathcal{O}(\|\mathbf{C}_{c}\|_{2})$, and
$\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}).$
For the row/left statement, we have
$\mathbf{M}\mathbf{M}^{\top}=\mathbf{D}_{r}(\mathbf{I}_{m}+\mathbf{C}_{r})\mathbf{D}_{r}=\mathbf{D}_{r}^{2}+\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$
Applying Lemma B.1 with $\mathbf{A}=\mathbf{D}_{r}^{2},\mathbf{E}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}$ gives
$(\mathbf{M}\mathbf{M}^{\top})^{-1/2}=\mathbf{D}_{r}^{-1}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{D}_{r}^{-1}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}),$
where $\mathbf{L}_{r}$ solves $\mathbf{D}_{r}\mathbf{L}_{r}+\mathbf{L}_{r}\mathbf{D}_{r}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$
Multiplying on the right by $\mathbf{M}$ yields
$\operatorname{Orth}_{\ell}(\mathbf{M})$ $=\mathbf{D}_{r}^{-1}\mathbf{M}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{D}_{r}^{-1}\mathbf{M}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2})$
$=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}).$
Again by Lemma B.2, we have
$\|\mathbf{L}_{r}\|_{2}\leq\frac{1}{2\lambda_{\min}(\mathbf{D}_{r})}\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}\leq\frac{\|\mathbf{D}_{r}\|_{2}^{2}}{2\lambda_{\min}(\mathbf{D}_{r})}\|\mathbf{C}_{r}\|_{2},$
so
$\|\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\|_{2}\leq\|\mathbf{D}_{r}^{-1}\|_{2}\|\mathbf{L}_{r}\|_{2}\leq\frac{\kappa(\mathbf{D}_{r})^{2}}{2}\|\mathbf{C}_{r}\|_{2}.$
Therefore
$\operatorname{Orth}_{\ell}(\mathbf{M})=\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}).$
Appendix D Normalization and whitening
This appendix records the normalization–whitening estimates omitted from the main text. We use the notation of Proposition 3.2. All $\mathcal{O}(\cdot)$ terms are understood in operator norm; the implied constants may depend on the fixed matrices in Proposition 3.2, but not on the perturbation terms.
Proposition D.1.
The following perturbative statements hold.
  1. (a)
    In the column/right setting of Proposition 3.2, if $\|\mathbf{C}_{c}\|_{2}$ is small enough, then $\operatorname{Orth}_{r}(\mathbf{N}_{c})=\mathbf{N}_{c}(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{N}_{c}-\tfrac{1}{2}\mathbf{N}_{c}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$, so $\|\operatorname{Orth}_{r}(\mathbf{N}_{c})-\mathbf{N}_{c}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{c}\|_{2}\|\mathbf{C}_{c}\|_{2}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$. Also,
    $\|\operatorname{Orth}_{r}(\mathbf{M})-\mathbf{M}\|_{2}\leq{}$ $\|\mathbf{N}_{c}\|_{2}\|\mathbf{D}_{c}-\mathbf{I}_{n}\|_{2}$
    $+\frac{\kappa(\mathbf{D}_{c})^{2}}{2}\|\mathbf{N}_{c}\|_{2}\|\mathbf{C}_{c}\|_{2}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}).$
  2. (b)
    In the row/left setting of Proposition 3.2, if $\|\mathbf{C}_{r}\|_{2}$ is small enough, then $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, so $\|\operatorname{Orth}_{\ell}(\mathbf{N}_{r})-\mathbf{N}_{r}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{r}\|_{2}\|\mathbf{C}_{r}\|_{2}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$. Also,
    $\|\operatorname{Orth}_{\ell}(\mathbf{M})-\mathbf{M}\|_{2}\leq{}$ $\|\mathbf{D}_{r}-\mathbf{I}_{m}\|_{2}\|\mathbf{N}_{r}\|_{2}$
    $+\frac{\kappa(\mathbf{D}_{r})^{2}}{2}\|\mathbf{N}_{r}\|_{2}\|\mathbf{C}_{r}\|_{2}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}).$
Proof.
For (a), Lemma B.1 with $\mathbf{A}=\mathbf{I}_{n}$ and $\mathbf{E}=\mathbf{C}_{c}$ gives the Sylvester equation $\mathbf{L}+\mathbf{L}=\mathbf{C}_{c}$, hence $\mathbf{L}=\tfrac{1}{2}\mathbf{C}_{c}$ and $(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{I}_{n}-\tfrac{1}{2}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$. Left multiplication by $\mathbf{N}_{c}$ gives the expansion of $\operatorname{Orth}_{r}(\mathbf{N}_{c})$, and the norm bound follows from submultiplicativity.
For the comparison with $\mathbf{M}$, Proposition 3.2 gives $\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2})$. Since $\mathbf{M}=\mathbf{N}_{c}\mathbf{D}_{c}$,
$\operatorname{Orth}_{r}(\mathbf{M})-\mathbf{M}=\mathbf{N}_{c}(\mathbf{I}_{n}-\mathbf{D}_{c})-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}).$
Lemma B.2 yields $\|\mathbf{L}_{c}\mathbf{D}_{c}^{-1}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{c})^{2}\|\mathbf{C}_{c}\|_{2}$, which gives the estimate.
The proof of (b) is the same. Lemma B.1 with $\mathbf{A}=\mathbf{I}_{m}$ and $\mathbf{E}=\mathbf{C}_{r}$ gives $(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}=\mathbf{I}_{m}-\tfrac{1}{2}\mathbf{C}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, hence $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$. The norm bound again follows from submultiplicativity. Proposition 3.2 also gives $\operatorname{Orth}_{\ell}(\mathbf{M})=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2})$, so
$\operatorname{Orth}_{\ell}(\mathbf{M})-\mathbf{M}=(\mathbf{I}_{m}-\mathbf{D}_{r})\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}).$
Using $\|\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{r})^{2}\|\mathbf{C}_{r}\|_{2}$ completes the proof. ∎
Remark D.1.
Proposition D.1 shows that one-sided normalization removes the zeroth-order marginal scale terms $\|\mathbf{N}_{c}\|_{2}\|\mathbf{D}_{c}-\mathbf{I}_{n}\|_{2}$ and $\|\mathbf{D}_{r}-\mathbf{I}_{m}\|_{2}\|\mathbf{N}_{r}\|_{2}$. Whitening then acts only on the first-order Gram error. Since $\operatorname{diag}(\mathbf{N}_{c}^{\top}\mathbf{N}_{c})=\mathbf{I}_{n}$ and $\operatorname{diag}(\mathbf{N}_{r}\mathbf{N}_{r}^{\top})=\mathbf{I}_{m}$, this leading correction is purely off-diagonal.
Corollary D.2.
Assume that every row and every column of $\mathbf{M}$ is nonzero, and set $\hat{\mathbf{M}}:=\mathbf{D}_{r}^{-1/2}\mathbf{M}\mathbf{D}_{c}^{-1/2}$.
  1. (a)
    If $\mathbf{M}$ has full column rank and $\hat{\mathbf{C}}_{c}:=\hat{\mathbf{M}}^{\top}\hat{\mathbf{M}}-\mathbf{I}_{n}$, then for $\|\hat{\mathbf{C}}_{c}\|_{2}$ small enough, $\operatorname{Orth}_{r}(\hat{\mathbf{M}})=\hat{\mathbf{M}}(\mathbf{I}_{n}+\hat{\mathbf{C}}_{c})^{-1/2}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{M}}\hat{\mathbf{C}}_{c}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$, so $\|\operatorname{Orth}_{r}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{c}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$.
  2. (b)
    If $\mathbf{M}$ has full row rank and $\hat{\mathbf{C}}_{r}:=\hat{\mathbf{M}}\hat{\mathbf{M}}^{\top}-\mathbf{I}_{m}$, then for $\|\hat{\mathbf{C}}_{r}\|_{2}$ small enough, $\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})=(\mathbf{I}_{m}+\hat{\mathbf{C}}_{r})^{-1/2}\hat{\mathbf{M}}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{C}}_{r}\hat{\mathbf{M}}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$, so $\|\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{r}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$.
Proof.
The same expansion applies to the two-sided normalized matrix $\hat{\mathbf{M}}$.
[…]
Remark D.2.
After two-sided normalization, whitening is centered at the identity Gram matrix of $\hat{\mathbf{M}}$. It no longer corrects marginal row or column scales; it only removes the residual mismatch left after diagonal equilibration.
Appendix E Lemmas for Proposition 3.4
E.1 Lemma E.1
Lemma E.1.
Let the momentum iterate satisfy $\mathbf{M}_{t}=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{G}_{t},\qquad\beta\in[0,1)$, and assume there exists a constant $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty},\|\mathbf{M}_{0}\|_{\infty}\leq G_{\infty}$. For $\varepsilon>0$, define $\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr),\mathbf{D}_{c,t}:=\operatorname{diag}\bigl(\operatorname{colsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr)$, and $\mathbf{P}_{t}:=\mathbf{D}_{r,t}^{-1/2},\mathbf{Q}_{t}:=\mathbf{D}_{c,t}^{-1/2}$. Then, for all $t$,
$(nG_{\infty}^{2}+\varepsilon)^{-1/2}\mathbf{I}_{m}\preceq\mathbf{P}_{t}\preceq\varepsilon^{-1/2}\mathbf{I}_{m},\quad(mG_{\infty}^{2}+\varepsilon)^{-1/2}\mathbf{I}_{n}\preceq\mathbf{Q}_{t}\preceq\varepsilon^{-1/2}\mathbf{I}_{n}.$
Moreover,
$\sqrt{\varepsilon}\mathbf{I}_{m}\preceq\mathbf{P}_{t}^{-1}\preceq\sqrt{nG_{\infty}^{2}+\varepsilon}\mathbf{I}_{m},\quad\sqrt{\varepsilon}\mathbf{I}_{n}\preceq\mathbf{Q}_{t}^{-1}\preceq\sqrt{mG_{\infty}^{2}+\varepsilon}\mathbf{I}_{n}.$
Proof.
We first show that the momentum iterate $\mathbf{M}_{t}$ is uniformly bounded entrywise.
Since
$\mathbf{M}_{t}=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{G}_{t},$
we have
$\|\mathbf{M}_{t}\|_{\infty}\leq\beta\|\mathbf{M}_{t-1}\|_{\infty}+(1-\beta)\|\mathbf{G}_{t}\|_{\infty}\leq\beta\|\mathbf{M}_{t-1}\|_{\infty}+(1-\beta)G_{\infty}.$
Using $\|\mathbf{M}_{0}\|_{\infty}=0\leq G_{\infty}$, an induction on $t$ yields
$\|\mathbf{M}_{t}\|_{\infty}\leq G_{\infty},\qquad\forall t.$
Therefore, every entry of $\mathbf{M}_{t}$ satisfies
$|(\mathbf{M}_{t})_{ij}|\leq G_{\infty},\qquad\forall i,j,t.$
Now fix any row $i$. Since that row has $n$ entries,
$\sum_{j=1}^{n}(\mathbf{M}_{t})_{ij}^{2}\leq nG_{\infty}^{2}.$
Hence each diagonal entry of $\mathbf{D}_{r,t}$ satisfies
$\varepsilon\leq(\mathbf{D}_{r,t})_{ii}=\sum_{j=1}^{n}(\mathbf{M}_{t})_{ij}^{2}+\varepsilon\leq nG_{\infty}^{2}+\varepsilon.$
Similarly, for any column $j$,
$\sum_{i=1}^{m}(\mathbf{M}_{t})_{ij}^{2}\leq mG_{\infty}^{2},$
so each diagonal entry of $\mathbf{D}_{c,t}$ satisfies
$\varepsilon\leq(\mathbf{D}_{c,t})_{jj}=\sum_{i=1}^{m}(\mathbf{M}_{t})_{ij}^{2}+\varepsilon\leq mG_{\infty}^{2}+\varepsilon.$
Since $\mathbf{P}_{t}=\mathbf{D}_{r,t}^{-1/2}$ and $\mathbf{Q}_{t}=\mathbf{D}_{c,t}^{-1/2}$ are diagonal matrices, their diagonal entries are
$(\mathbf{P}_{t})_{ii}=\left(\sum_{j=1}^{n}(\mathbf{M}_{t})_{ij}^{2}+\varepsilon\right)^{-1/2},\quad(\mathbf{Q}_{t})_{jj}=\left(\sum_{i=1}^{m}(\mathbf{M}_{t})_{ij}^{2}+\varepsilon\right)^{-1/2}.$
Using the bounds above, we obtain
$(nG_{\infty}^{2}+\varepsilon)^{-1/2}\leq(\mathbf{P}_{t})_{ii}\leq\varepsilon^{-1/2},\quad(mG_{\infty}^{2}+\varepsilon)^{-1/2}\leq(\mathbf{Q}_{t})_{jj}\leq\varepsilon^{-1/2}.$
Since positive diagonal matrices are ordered entrywise in the Loewner sense, it follows that
$(nG_{\infty}^{2}+\varepsilon)^{-1/2}\mathbf{I}_{m}\preceq\mathbf{P}_{t}\preceq\varepsilon^{-1/2}\mathbf{I}_{m},\quad(mG_{\infty}^{2}+\varepsilon)^{-1/2}\mathbf{I}_{n}\preceq\mathbf{Q}_{t}\preceq\varepsilon^{-1/2}\mathbf{I}_{n}.$
Because $\mathbf{P}_{t}$ and $\mathbf{Q}_{t}$ are symmetric positive definite diagonal matrices, their singular values are exactly their diagonal entries. Therefore,
$(nG_{\infty}^{2}+\varepsilon)^{-1/2}\leq\sigma_{\min}(\mathbf{P}_{t})\leq\|\mathbf{P}_{t}\|_{2}\leq\varepsilon^{-1/2},\quad(mG_{\infty}^{2}+\varepsilon)^{-1/2}\leq\sigma_{\min}(\mathbf{Q}_{t})\leq\|\mathbf{Q}_{t}\|_{2}\leq\varepsilon^{-1/2}.$
Finally, inverting the above matrix inequalities gives
$\sqrt{\varepsilon}\mathbf{I}_{m}\preceq\mathbf{P}_{t}^{-1}\preceq\sqrt{nG_{\infty}^{2}+\varepsilon}\mathbf{I}_{m},\quad\sqrt{\varepsilon}\mathbf{I}_{n}\preceq\mathbf{Q}_{t}^{-1}\preceq\sqrt{mG_{\infty}^{2}+\varepsilon}\mathbf{I}_{n}.$
This completes the proof. ∎
E.2 Lemma E.2
Lemma E.2.
Suppose that $\{E_{t},A_{t}\}$ are nonnegative sequences. Assume $E_{t+1}\leq(1-\alpha_{t+1})E_{t}+A_{t+1}$ where $\alpha_{t}=t^{-p}$, $p\in(0,1]$. Then $\alpha_{t}E_{t}\leq 2(E_{t}-E_{t+1}+A_{t+1}).$
Proof.
See [6], Lemma A.3. ∎
E.3 Lemma E.3
Lemma E.3.
For Algorithm 1, the accumulated error between the momentum term and the true gradient is bounded for $t\geq 1$:
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
Proof.
First, we have
$\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$=\|\beta_{t+1}\mathbf{M}_{t}+(1-\beta_{t+1})\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$=\|\beta_{t+1}(\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}))+(1-\beta_{t+1})(\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1}))$
$\quad+\beta_{t+1}(\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}))\|_{F}^{2}$
$=\beta_{t+1}^{2}\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+\beta_{t+1}^{2}\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$\quad+(1-\beta_{t+1})^{2}\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$
$\quad+2\beta_{t+1}^{2}\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}$
$\quad+2\beta_{t+1}(1-\beta_{t+1})\langle\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1}),\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}.$
According to Assumption 3.3. Taking the expectation of its squared norm, and using the unbiasedness and independence of the stochastic gradient, we obtain:
$\mathbb{E}[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}]=$ $\beta_{t+1}^{2}\mathbb{E}[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}]$
$\quad+(1-\beta_{t+1})^{2}\mathbb{E}[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}]$
$+\beta_{t+1}^{2}\mathbb{E}[\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}]$
$+2\beta_{t+1}^{2}\mathbb{E}[\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle].$
Applying Young’s inequality with a parameter ($ab\leq\frac{\epsilon}{2}a^{2}+\frac{1}{2\epsilon}b^{2}$), we have
$\langle\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t}),\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\rangle_{F}\leq\frac{\epsilon}{2}\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}+\frac{1}{2\epsilon}\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}.$
Thus, we have:
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$ $\leq\beta_{t+1}^{2}(1+\epsilon)\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]$
$\quad+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]$
$\quad+(1-\beta_{t+1})^{2}\mathbb{E}\left[\|\nabla f(\mathbf{X}_{t+1};\xi_{t+1})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right].$
According to Assumption 3.2,
$\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$ $\leq L^{2}\|\mathbf{X}_{t}-\mathbf{X}_{t+1}\|_{F}^{2}$
$=L^{2}\eta_{t}^{2}\|\mathbf{O}_{t}\|_{F}^{2}$
$\leq L^{2}\eta_{t}^{2}a^{2}n.$
Therefore:
$\mathbb{E}\left[\left\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\right\|_{F}^{2}\right]$ $\leq\beta_{t+1}^{2}(1+\epsilon)\mathbb{E}\left[\left\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\right\|_{F}^{2}\right]$
$+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
Then, by letting $\epsilon:=\frac{1-\beta_{t+1}}{\beta_{t+1}},t\geq 1$, we have
$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$ (6)
Appendix F Proofs of Proposition 3.4
Proof.
By Assumption 3.2 and the descent lemma,
$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})+\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$=f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{La^{2}\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{La^{2}n}{2}\eta_{t}^{2},$
where the last step uses $\|\mathbf{O}_{t}\|_{F}^{2}\leq n$.
Let $\mathbf{S}_{t}:=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},r_{\varepsilon}:=\sqrt{\varepsilon},R_{\varepsilon}:=\sqrt{nG_{\infty}^{2}+\varepsilon},C_{\varepsilon}:=\sqrt{mG_{\infty}^{2}+\varepsilon}.$
Define the midpoint-rescaled matrices
$\widetilde{\mathbf{P}}_{t}:=\frac{R_{\varepsilon}+r_{\varepsilon}}{2}\mathbf{P}_{t},\widetilde{\mathbf{Q}}_{t}:=\frac{C_{\varepsilon}+r_{\varepsilon}}{2}\mathbf{Q}_{t},\widetilde{\mathbf{A}}_{t}:=\widetilde{\mathbf{P}}_{t}^{-1},\widetilde{\mathbf{B}}_{t}:=\widetilde{\mathbf{Q}}_{t}^{-1}.$
Since the rescaling is by positive scalars,
$\operatorname{polar}(\widetilde{\mathbf{P}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{Q}}_{t})=\operatorname{polar}(\mathbf{P}_{t}\mathbf{M}_{t}\mathbf{Q}_{t})=\mathbf{O}_{t}.$
Hence
$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle=\left\langle\widetilde{\mathbf{A}}_{t}\nabla f(\mathbf{X}_{t})\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle.$
By Lemma E.1, we have
$R_{\varepsilon}^{-1}\mathbf{I}_{m}\preceq\mathbf{P}_{t}\preceq r_{\varepsilon}^{-1}\mathbf{I}_{m},\qquad C_{\varepsilon}^{-1}\mathbf{I}_{n}\preceq\mathbf{Q}_{t}\preceq r_{\varepsilon}^{-1}\mathbf{I}_{n}.$
Since $R_{\varepsilon}=r_{\varepsilon}\rho_{r},C_{\varepsilon}=r_{\varepsilon}\rho_{c}$, we obtain
$\frac{\rho_{r}+1}{2\rho_{r}}\mathbf{I}_{m}\preceq\widetilde{\mathbf{P}}_{t}\preceq\frac{\rho_{r}+1}{2}\mathbf{I}_{m},\qquad\frac{\rho_{c}+1}{2\rho_{c}}\mathbf{I}_{n}\preceq\widetilde{\mathbf{Q}}_{t}\preceq\frac{\rho_{c}+1}{2}\mathbf{I}_{n}.$
Therefore, we have
$\|\widetilde{\mathbf{A}}_{t}\|_{2}\leq\frac{2\rho_{r}}{\rho_{r}+1},\quad\|\widetilde{\mathbf{B}}_{t}\|_{2}\leq\frac{2\rho_{c}}{\rho_{c}+1},\quad\|\widetilde{\mathbf{A}}_{t}-\mathbf{I}_{m}\|_{2}\leq\frac{\rho_{r}-1}{\rho_{r}+1},\quad\|\widetilde{\mathbf{B}}_{t}-\mathbf{I}_{n}\|_{2}\leq\frac{\rho_{c}-1}{\rho_{c}+1}.$
Using $\nabla f(\mathbf{X}_{t})=\mathbf{M}_{t}+\mathbf{S}_{t}$, we split
$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle$ $=\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle+\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{S}_{t}\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle$
$=\|\widetilde{\mathbf{P}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{Q}}_{t}\|_{*}+\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{S}_{t}\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle+\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t}-\mathbf{M}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle.$
First, by the singular-value inequality
$\sigma_{i}(\mathbf{A}\mathbf{X}\mathbf{B})\geq\sigma_{\min}(\mathbf{A})\sigma_{i}(\mathbf{X})\sigma_{\min}(\mathbf{B}),$
we have
$\|\widetilde{\mathbf{P}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{Q}}_{t}\|_{*}\geq\frac{(\rho_{r}+1)(\rho_{c}+1)}{4\rho_{r}\rho_{c}}\|\mathbf{M}_{t}\|_{*}.$
Second, using Hölder’s inequality, $\|\mathbf{O}_{t}\|_{2}=1$, and the bounds above,
$\left|\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{S}_{t}\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle\right|$ $\leq\|\widetilde{\mathbf{A}}_{t}\mathbf{S}_{t}\widetilde{\mathbf{B}}_{t}\|_{*}\|\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\|_{2}$
$\leq\|\widetilde{\mathbf{A}}_{t}\|_{2}\|\widetilde{\mathbf{B}}_{t}\|_{2}\|\mathbf{S}_{t}\|_{*}\|\widetilde{\mathbf{P}}_{t}\|_{2}\|\widetilde{\mathbf{Q}}_{t}\|_{2}$
$\leq\rho_{r}\rho_{c}\|\mathbf{S}_{t}\|_{*}$
$\leq\rho_{r}\rho_{c}\sqrt{n}\|\mathbf{S}_{t}\|_{F}.$
Third, we use the symmetric decomposition
$\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t}-\mathbf{M}_{t}=(\widetilde{\mathbf{A}}_{t}-\mathbf{I}_{m})\mathbf{M}_{t}+\mathbf{M}_{t}(\widetilde{\mathbf{B}}_{t}-\mathbf{I}_{n})+(\widetilde{\mathbf{A}}_{t}-\mathbf{I}_{m})\mathbf{M}_{t}(\widetilde{\mathbf{B}}_{t}-\mathbf{I}_{n}).$
Hence
$\|\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t}-\mathbf{M}_{t}\|_{*}$ $\leq\left(\frac{\rho_{r}-1}{\rho_{r}+1}+\frac{\rho_{c}-1}{\rho_{c}+1}+\frac{(\rho_{r}-1)(\rho_{c}-1)}{(\rho_{r}+1)(\rho_{c}+1)}\right)\|\mathbf{M}_{t}\|_{*}.$
Therefore, we have
$\left|\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t}-\mathbf{M}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle\right|$ $\leq\|\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t}-\mathbf{M}_{t}\|_{*}\|\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\|_{2}$
$\leq\frac{3\rho_{r}\rho_{c}-\rho_{r}-\rho_{c}-1}{4}\|\mathbf{M}_{t}\|_{*}.$
Combining the above bounds, we obtain
$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle\geq\chi_{\varepsilon}\|\mathbf{M}_{t}\|_{*}-\rho_{r}\rho_{c}\sqrt{n}\|\mathbf{S}_{t}\|_{F}.$
Under the setting $\varepsilon\geq\frac{4}{5}G_{\infty}^{2}\max\{m,n\}$, we have $\rho_{r}=\sqrt{1+\frac{nG_{\infty}^{2}}{\varepsilon}}\leq\sqrt{1+\frac{5n}{4\max\{m,n\}}}\leq\frac{3}{2},\rho_{c}=\sqrt{1+\frac{mG_{\infty}^{2}}{\varepsilon}}\leq\sqrt{1+\frac{5m}{4\max\{m,n\}}}\leq\frac{3}{2}$.
Let $\psi(x,y):=\frac{(x+1)(y+1)}{4xy}-\frac{3xy-x-y-1}{4},x,y\geq 1.$ A direct calculation gives $\partial_{x}\psi(x,y)=\frac{1-x^{-2}-x^{-2}y^{-1}-3y}{4}<0,\partial_{y}\psi(x,y)=\frac{1-y^{-2}-x^{-1}y^{-2}-3x}{4}<0$, for all $x,y\geq 1$. Hence $\psi$ is decreasing in each argument on $1,\infty)^{2}$. Therefore, we have $\chi_{\varepsilon}=\psi(\rho_{r},\rho_{c})\geq\psi\left(\frac{3}{2},\frac{3}{2}\right)=\frac{1}{144}>0$. Since $\chi_{\varepsilon}>0$, $\|\mathbf{M}_{t}\|_{*}\geq\|\mathbf{M}_{t}\|_{F}$, and
$\|\mathbf{M}_{t}\|_{F}\geq\|\nabla f(\mathbf{X}_{t})\|_{F}-\|\mathbf{S}_{t}\|_{F},$
it follows that
$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle\geq\chi_{\varepsilon}\|\nabla f(\mathbf{X}_{t})\|_{F}-\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)\|\mathbf{S}_{t}\|_{F}.$
Substituting the above inequality into the descent bound yields
$f(\mathbf{X}_{t+1})\leq$ $f(\mathbf{X}_{t})-a\chi_{\varepsilon}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}$
$\quad+a\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)\eta_{t}\|\mathbf{S}_{t}\|_{F}+\frac{La^{2}n}{2}\eta_{t}^{2}.$
Apply Young’s inequality with parameter $h_{t+1}/L$:
$a\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)\eta_{t}\|\mathbf{S}_{t}\|_{F}\leq\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{aL}{2}\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}.$
Hence
$f(\mathbf{X}_{t+1})\leq$ $f(\mathbf{X}_{t})-a\chi_{\varepsilon}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}$
$+\frac{aL}{2}\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}n}{2}\eta_{t}^{2}.$
Taking expectations and summing the above inequality over $t=1,\dots,T$, then using $f(\mathbf{X}_{T+1})\geq f^{\star}$, we have
$a\chi_{\varepsilon}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $\leq f(\mathbf{X}_{1})-f^{\star}+\frac{a}{2L}\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
$+\frac{aL}{2}\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}n}{2}\sum_{t=1}^{T}\eta_{t}^{2}.$
By Lemma E.3,
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{(1-h_{t+1})^{2}}{h_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+h_{t+1}^{2}\sigma^{2}.$
Since $(1-h_{t+1})^{2}\leq 1$ and
$\frac{\eta_{t}^{2}}{h_{t+1}}=\frac{t^{-3/2}}{(t+1)^{-1/2}}=\frac{\sqrt{t+1}}{t^{3/2}}\leq\frac{2\sqrt{2}}{t+1}=2\sqrt{2}h_{t+1}^{2},$
we have
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+h_{t+1}^{2}(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}).$
According to Lemma E.2, by letting $E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},A_{t+1}=h_{t+1}^{2}(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})$, we have
$h_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\left(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}\right).$
Moreover, since $\beta_{1}=0$,
$\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}$ $=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\mathbf{M}_{1}\|_{F}^{2}$
$=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}.$
It follows that
$\sum_{t=1}^{T}h_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq 2\sum_{t=1}^{T}\left(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}\right)$
$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})\sum_{t=1}^{T}\frac{1}{t+1}$
$\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})(1+\ln T).$
Hence, since $h_{t+1}\leq h_{t}$,
$\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\sigma^{2}+2(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})(1+\ln T).$
Also,
$\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}=\sum_{t=1}^{T}\frac{t^{-3/2}}{(t+1)^{-1/2}}\leq 2(1+\ln T),\qquad\sum_{t=1}^{T}\eta_{t}^{2}=\sum_{t=1}^{T}t^{-3/2}\leq 3.$
Substituting these bounds into the previous inequality, we get
$a\chi_{\varepsilon}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq f(\mathbf{X}_{1})-f^{\star}+C_{1}(1+\ln T)+C_{2},$
where
$C_{1}=\frac{a}{L}\bigl(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}\bigr)+aL\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2},\quad C_{2}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}n}{2}.$
Finally, since $\eta_{t}=t^{-3/4}\geq T^{-3/4}$ for all $1\leq t\leq T$,
$T^{-3/4}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}.$
Therefore, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}(1+\ln T)+C_{2}}{a\chi_{\varepsilon}\cdot T^{1/4}}.$
This completes the proof. ∎
Appendix G Lemmas for Theorem 3.5
G.1 Lemma G.1
Lemma G.1.
Let $\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\mathbf{R}_{t}:=\mathbf{D}_{r,t}^{-1/2}\mathbf{M}_{t},\mathbf{O}_{t}:=\operatorname{Orth}(\mathbf{R}_{t})$. Then, for all $t$,
$\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F},\qquad\|\mathbf{O}_{t}\|_{F}^{2}=\operatorname{rank}(\mathbf{R}_{t})\leq n.$
Proof.
If $\mathbf{M}_{t}=\mathbf{0}$, then $\mathbf{R}_{t}=\mathbf{O}_{t}=\mathbf{0}$ and the claim is immediate. Assume $\mathbf{M}_{t}\neq\mathbf{0}$. Write
$d_{i,t}:=\|(\mathbf{M}_{t})_{i,:}\|_{2},\qquad\mathbf{D}_{t}:=\operatorname{diag}(d_{1,t},\ldots,d_{m,t}),$
so that $\mathbf{D}_{t}=\mathbf{D}_{r,t}^{1/2}$ and $\mathbf{M}_{t}=\mathbf{D}_{t}\mathbf{R}_{t}$. Let $\mathbf{H}_{t}:=(\mathbf{R}_{t}\mathbf{R}_{t}^{\top})^{1/2}$. If $\mathbf{R}_{t}=\mathbf{U}_{t}\boldsymbol{\Sigma}_{t}\mathbf{V}_{t}^{\top}$ is a compact SVD, then $\mathbf{O}_{t}=\mathbf{U}_{t}\mathbf{V}_{t}^{\top}$ and therefore $\mathbf{O}_{t}\mathbf{R}_{t}^{\top}=\mathbf{H}_{t}$. Hence
$\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=\operatorname{tr}(\mathbf{O}_{t}^{\top}\mathbf{D}_{t}\mathbf{R}_{t})=\operatorname{tr}(\mathbf{D}_{t}\mathbf{O}_{t}\mathbf{R}_{t}^{\top})=\operatorname{tr}(\mathbf{D}_{t}\mathbf{H}_{t})=\sum_{i=1}^{m}d_{i,t}(\mathbf{H}_{t})_{ii}.$
If $d_{i,t}>0$, then the $i$th row of $\mathbf{R}_{t}$ has unit Euclidean norm, so
$1=\mathbf{e}_{i}^{\top}\mathbf{R}_{t}\mathbf{R}_{t}^{\top}\mathbf{e}_{i}=\mathbf{e}_{i}^{\top}\mathbf{H}_{t}^{2}\mathbf{e}_{i}\leq\|\mathbf{H}_{t}\|_{2}\mathbf{e}_{i}^{\top}\mathbf{H}_{t}\mathbf{e}_{i}=\|\mathbf{R}_{t}\|_{2}(\mathbf{H}_{t})_{ii}.$
Thus $(\mathbf{H}_{t})_{ii}\geq\|\mathbf{R}_{t}\|_{2}^{-1}$ whenever $d_{i,t}>0$. If $d_{i,t}=0$, the corresponding term vanishes. Therefore,
$\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1}{\|\mathbf{R}_{t}\|_{2}}\sum_{i=1}^{m}d_{i,t}.$
Now each nonzero row of $\mathbf{R}_{t}$ has norm $1$, so $\|\mathbf{R}_{t}\|_{F}^{2}\leq m$, hence $\|\mathbf{R}_{t}\|_{2}\leq\sqrt{m}$. Also, $\sum_{i=1}^{m}d_{i,t}\geq\|\mathbf{M}_{t}\|_{F}$. It follows that
$\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}.$
Finally, if $r_{t}:=\operatorname{rank}(\mathbf{R}_{t})$, then the nonzero singular values of $\mathbf{O}_{t}$ are all equal to $1$, so $\|\mathbf{O}_{t}\|_{F}^{2}=r_{t}\leq n$. ∎
G.2 Lemma G.2
Consider the practical row-normalized R variant (4), with Nesterov disabled and $\lambda=0,\varepsilon=0$. Then the update is
$\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\hat{\mathbf{M}}_{t}:=(\mathbf{D}_{r,t}^{1/2})^{\dagger}\mathbf{M}_{t},\mathbf{O}_{t}:=\mathrm{NS5}(\hat{\mathbf{M}}_{t}),\mathbf{X}_{t+1}=\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t},$
where $a:=0.2\sqrt{\max\{m,n\}}$. Here denotes the Moore–Penrose pseudoinverse, so a zero row remains zero after row normalization, and we adopt the convention $\operatorname{Orth}(\mathbf{0})=\mathbf{0}$.
Remark G.1.
Following the practical Muon implementation [20] and Theorem 3.1, we write $\mathrm{NS5}$ as the five-step pre-scaled Newton–Schulz map. For any $\mathbf{A}\in\mathbb{R}^{m\times n}$, define $\alpha(\mathbf{A}):=\max\{1,\|\mathbf{A}\|_{F}\},\mathbf{Y}^{(0)}(\mathbf{A}):=\mathbf{A}/\alpha(\mathbf{A}),$ and, for $k=0,1,2,3,4$,
$\mathbf{Y}^{(k+1)}(\mathbf{A})=p_{\mathrm{ns}}\bigl(\mathbf{Y}^{(k)}(\mathbf{A})\mathbf{Y}^{(k)}(\mathbf{A})^{\top}\bigr)\mathbf{Y}^{(k)}(\mathbf{A}),\qquad p_{\mathrm{ns}}(\mathbf{Z}):=a_{\mathrm{ns}}\mathbf{I}_{m}+b_{\mathrm{ns}}\mathbf{Z}+c_{\mathrm{ns}}\mathbf{Z}^{2}.$
We then define
$\mathrm{NS5}(\mathbf{A}):=\mathbf{Y}^{(5)}(\mathbf{A}).$
Remark G.2.
Following the trajectory-wise error formalism in [17, Definition 3 and Theorem 2], we work with the NS5 polar-approximation error along the algorithmic trajectory. We define
$\varepsilon_{\mathrm{ns}}:=\sup_{t\geq 1}\bigl\|\mathrm{NS5}(\hat{\mathbf{M}}_{t})-\operatorname{Orth}(\hat{\mathbf{M}}_{t})\bigr\|_{2}.$
In subsequent derivations, we assume $\varepsilon_{\mathrm{ns}}<1$. The quantity $\varepsilon_{\mathrm{ns}}$ plays the same role as the trajectory-wise polar-approximation error $\varepsilon_{q}$ in [17, Definition 3 and Theorem 2].
Lemma G.2.
Let $\mathbf{A}\in\mathbb{R}^{m\times n}$. If $\mathbf{A}=\mathbf{0}$, define $\operatorname{Orth}(\mathbf{A})=\mathbf{0}$ and $\mathrm{NS5}(\mathbf{A})=\mathbf{0}$. Otherwise, let $\mathbf{A}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$ be the compact SVD, and set $\mathbf{Q}:=\operatorname{Orth}(\mathbf{A})=\mathbf{U}\mathbf{V}^{\top}$ and $\mathbf{O}:=\mathrm{NS5}(\mathbf{A})$. Then there exists a diagonal matrix $\widetilde{\boldsymbol{\Sigma}}$ such that $\mathbf{O}=\mathbf{U}\widetilde{\boldsymbol{\Sigma}}\mathbf{V}^{\top}.$ Moreover, for every row-normalized input $\mathbf{A}=\hat{\mathbf{M}}_{t}$ along the algorithmic trajectory,
$\|\mathrm{NS5}(\mathbf{A})-\operatorname{Orth}(\mathbf{A})\|_{2}\leq\varepsilon_{\mathrm{ns}}.$
Consequently, every diagonal entry $\tilde{\sigma}_{i}$ of $\widetilde{\boldsymbol{\Sigma}}$ satisfies
$1-\varepsilon_{\mathrm{ns}}\leq\tilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}},\qquad i\in[\operatorname{rank}(\mathbf{A})].$
In particular, whenever $\mathbf{A}=\hat{\mathbf{M}}_{t}\neq\mathbf{0}$, we have $\operatorname{Orth}(\mathrm{NS5}(\mathbf{A}))=\operatorname{Orth}(\mathbf{A}).$
Proof.
If $\mathbf{A}=\mathbf{0}$, the conclusion is immediate by definition. Assume $\mathbf{A}\neq\mathbf{0}$.
The internal scaling multiplies $\mathbf{A}$ by a positive scalar and hence does not change its left or right singular vectors. Thus
$\mathbf{Y}^{(0)}(\mathbf{A})=\mathbf{U}\boldsymbol{\Sigma}^{(0)}\mathbf{V}^{\top}$
for some diagonal matrix $\boldsymbol{\Sigma}^{(0)}$. Assume inductively that
$\mathbf{Y}^{(k)}(\mathbf{A})=\mathbf{U}\boldsymbol{\Sigma}^{(k)}\mathbf{V}^{\top}$
with diagonal $\boldsymbol{\Sigma}^{(k)}$. Then
$\mathbf{Y}^{(k)}(\mathbf{A})\mathbf{Y}^{(k)}(\mathbf{A})^{\top}=\mathbf{U}\bigl(\boldsymbol{\Sigma}^{(k)}\bigr)^{2}\mathbf{U}^{\top},$
and therefore
$\mathbf{Y}^{(k+1)}(\mathbf{A})=\mathbf{U}p_{\mathrm{ns}}\bigl((\boldsymbol{\Sigma}^{(k)})^{2}\bigr)\boldsymbol{\Sigma}^{(k)}\mathbf{V}^{\top}.$
Hence every iterate has the same left and right singular vectors as $\mathbf{A}$, and in particular
$\mathrm{NS5}(\mathbf{A})=\mathbf{U}\widetilde{\boldsymbol{\Sigma}}\mathbf{V}^{\top}$
for some diagonal matrix $\widetilde{\boldsymbol{\Sigma}}$.
Now, if $\mathbf{A}=\hat{\mathbf{M}}_{t}$ is a row-normalized input on the algorithmic trajectory, then by the definition of $\varepsilon_{\mathrm{ns}}$,
$\|\mathrm{NS5}(\mathbf{A})-\operatorname{Orth}(\mathbf{A})\|_{2}=\|\mathbf{U}(\widetilde{\boldsymbol{\Sigma}}-\mathbf{I})\mathbf{V}^{\top}\|_{2}=\|\widetilde{\boldsymbol{\Sigma}}-\mathbf{I}\|_{2}\leq\varepsilon_{\mathrm{ns}}.$
Hence
$\max_{i}|\tilde{\sigma}_{i}-1|\leq\varepsilon_{\mathrm{ns}},$
which is equivalent to
$1-\varepsilon_{\mathrm{ns}}\leq\tilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}}\qquad\text{for all }i.$
Since $\varepsilon_{\mathrm{ns}}<1$, these diagonal entries are strictly positive. Therefore
$\operatorname{Orth}(\mathrm{NS5}(\mathbf{A}))=\operatorname{Orth}(\mathbf{A})$
whenever $\mathbf{A}=\hat{\mathbf{M}}_{t}\neq\mathbf{0}$. ∎
G.3 Lemma G.3
Lemma G.3.
Under Lemma G.2, for all $t$,
$\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F},\qquad\|\mathbf{O}_{t}\|_{F}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n.$
Proof.
If $\mathbf{M}_{t}=\mathbf{0}$, then $\hat{\mathbf{M}}_{t}=\mathbf{O}_{t}=\mathbf{0}$ and the claim is immediate. Assume $\mathbf{M}_{t}\neq\mathbf{0}$. We have
$d_{i,t}:=\|(\mathbf{M}_{t})_{i,:}\|_{2},\qquad\mathbf{D}_{t}:=\mathbf{D}_{r,t}^{1/2}=\operatorname{diag}(d_{1,t},\dots,d_{m,t}),$
so that
$\mathbf{M}_{t}=\mathbf{D}_{t}\hat{\mathbf{M}}_{t}.$
Let
$\hat{\mathbf{M}}_{t}=\mathbf{U}_{t}\boldsymbol{\Sigma}_{t}\mathbf{V}_{t}^{\top},\qquad\mathbf{H}_{t}:=(\hat{\mathbf{M}}_{t}\hat{\mathbf{M}}_{t}^{\top})^{1/2}=\mathbf{U}_{t}\boldsymbol{\Sigma}_{t}\mathbf{U}_{t}^{\top}.$
By Lemma G.2,
$\mathbf{O}_{t}=\mathbf{U}_{t}\widetilde{\boldsymbol{\Sigma}}_{t}\mathbf{V}_{t}^{\top}$
with diagonal entries of $\widetilde{\boldsymbol{\Sigma}}_{t}$ lying in $1-\varepsilon_{\mathrm{ns}},1+\varepsilon_{\mathrm{ns}}]$. Therefore,
$\mathbf{O}_{t}\hat{\mathbf{M}}_{t}^{\top}=\mathbf{U}_{t}\widetilde{\boldsymbol{\Sigma}}_{t}\boldsymbol{\Sigma}_{t}\mathbf{U}_{t}^{\top}\succeq(1-\varepsilon_{\mathrm{ns}})\mathbf{U}_{t}\boldsymbol{\Sigma}_{t}\mathbf{U}_{t}^{\top}=(1-\varepsilon_{\mathrm{ns}})\mathbf{H}_{t}.$
Hence
$\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle=\operatorname{tr}(\mathbf{O}_{t}^{\top}\mathbf{D}_{t}\hat{\mathbf{M}}_{t})=\operatorname{tr}(\mathbf{D}_{t}\mathbf{O}_{t}\hat{\mathbf{M}}_{t}^{\top})\geq(1-\varepsilon_{\mathrm{ns}})\operatorname{tr}(\mathbf{D}_{t}\mathbf{H}_{t})=(1-\varepsilon_{\mathrm{ns}})\sum_{i=1}^{m}d_{i,t}(\mathbf{H}_{t})_{ii}.$
If $d_{i,t}>0$, then the $i$-th row of $\hat{\mathbf{M}}_{t}$ has unit Euclidean norm, so
$1=\mathbf{e}_{i}^{\top}\hat{\mathbf{M}}_{t}\hat{\mathbf{M}}_{t}^{\top}\mathbf{e}_{i}=\mathbf{e}_{i}^{\top}\mathbf{H}_{t}^{2}\mathbf{e}_{i}\leq\|\mathbf{H}_{t}\|_{2}\mathbf{e}_{i}^{\top}\mathbf{H}_{t}\mathbf{e}_{i}=\|\hat{\mathbf{M}}_{t}\|_{2}(\mathbf{H}_{t})_{ii}.$
Thus $(\mathbf{H}_{t})_{ii}\geq\|\hat{\mathbf{M}}_{t}\|_{2}^{-1}$ whenever $d_{i,t}>0$. If $d_{i,t}=0$, the corresponding term vanishes. Consequently,
$\operatorname{tr}(\mathbf{D}_{t}\mathbf{H}_{t})\geq\|\hat{\mathbf{M}}_{t}\|_{2}^{-1}\sum_{i=1}^{m}d_{i,t}.$
Now each nonzero row of $\hat{\mathbf{M}}_{t}$ has norm $1$, so $\|\hat{\mathbf{M}}_{t}\|_{F}^{2}\leq m$, hence $\|\hat{\mathbf{M}}_{t}\|_{2}\leq\sqrt{m}$. Also,
$\sum_{i=1}^{m}d_{i,t}\geq\Bigl(\sum_{i=1}^{m}d_{i,t}^{2}\Bigr)^{1/2}=\|\mathbf{M}_{t}\|_{F}.$
Combining the above bounds gives
$\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}.$
For the norm bound, let $r_{t}:=\operatorname{rank}(\hat{\mathbf{M}}_{t})$. Again by Lemma G.2, every nonzero singular value of $\mathbf{O}_{t}$ is at most $1+\varepsilon_{\mathrm{ns}}$. Therefore
$\|\mathbf{O}_{t}\|_{F}^{2}=\sum_{i=1}^{r_{t}}\tilde{\sigma}_{i,t}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}r_{t}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n.$
This completes the proof. ∎
Appendix H Proofs of Theorem 3.5
Proof.
Let $\mathbf{S}_{t}:=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\gamma_{\mathrm{ns}}:=1-\varepsilon_{\mathrm{ns}},c_{m,n}^{\mathrm{ns}}:=\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}$. By Assumption 3.2 and the descent lemma, we have
$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})+\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
$=f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{La^{2}\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
$\leq f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2},$ (7)
where the last step uses Lemma G.3.
Again by Lemma G.3,
$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle$ $=\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\langle\mathbf{S}_{t},\mathbf{O}_{t}\rangle$
$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-\|\mathbf{S}_{t}\|_{F}\|\mathbf{O}_{t}\|_{F}$
$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\|\mathbf{S}_{t}\|_{F}$
$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\nabla f(\mathbf{X}_{t})\|_{F}-\Bigl(\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)\|\mathbf{S}_{t}\|_{F}.$ (8)
Substituting (H) into (H),
$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+ac_{m,n}^{\mathrm{ns}}\eta_{t}\|\mathbf{S}_{t}\|_{F}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2}.$ (9)
Apply Young’s inequality with parameter $h_{t+1}/L$:
$ac_{m,n}^{\mathrm{ns}}\eta_{t}\|\mathbf{S}_{t}\|_{F}\leq\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}.$
Hence
$f(\mathbf{X}_{t+1})\leq$ $\ f(\mathbf{X}_{t})-\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}$
$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2}.$
Taking expectations and summing from $t=1$ to $T$, then using $f(\mathbf{X}_{T+1})\geq f^{\star}$, we obtain
$\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $\leq f(\mathbf{X}_{1})-f^{\star}+\frac{a}{2L}\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\sum_{t=1}^{T}\eta_{t}^{2}.$ (10)
A direct repetition of the proof of Lemma E.3, using $\|\mathbf{O}_{t}\|_{F}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n$ from Lemma G.3, gives
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{(1-h_{t+1})^{2}}{h_{t+1}}L^{2}\eta_{t}^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+h_{t+1}^{2}\sigma^{2}.$
Since $(1-h_{t+1})^{2}\leq 1$ and
$\frac{\eta_{t}^{2}}{h_{t+1}}=\frac{t^{-3/2}}{(t+1)^{-1/2}}=\frac{\sqrt{t+1}}{t^{3/2}}\leq\frac{2\sqrt{2}}{t+1}=2\sqrt{2}h_{t+1}^{2},$
we have
$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr).$
Applying Lemma E.2 with
$E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},\qquad A_{t+1}=h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr),$
we get
$h_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\Bigl(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}\Bigr).$
Moreover, since $\beta_{1}=0$,
$\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\mathbf{M}_{1}\|_{F}^{2}=\mathbb{E}\|\nabla f(\mathbf{X}_{1})-\nabla f(\mathbf{X}_{1};\xi_{1})\|_{F}^{2}\leq\sigma^{2}.$
Thus,
$\sum_{t=1}^{T}h_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq 2\sum_{t=1}^{T}\Bigl(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}\Bigr)$
$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)\sum_{t=1}^{T}\frac{1}{t+1}$
$\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)(1+\ln T).$
Hence, since $h_{t+1}\leq h_{t}$,
$\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)(1+\ln T).$
Also,
$\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}=\sum_{t=1}^{T}\frac{t^{-3/2}}{(t+1)^{-1/2}}\leq 2(1+\ln T),\qquad\sum_{t=1}^{T}\eta_{t}^{2}=\sum_{t=1}^{T}t^{-3/2}\leq 3.$
Substituting these bounds into (H),
$\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}.$
Finally, since $\eta_{t}=t^{-3/4}\geq T^{-3/4}$ for all $1\leq t\leq T$,
$T^{-3/4}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}.$
Therefore,
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}}{\bigl(a\gamma_{\mathrm{ns}}/\sqrt{m}\bigr)T^{1/4}}.$
This completes the proof. ∎
Appendix I More Results
[…]
(a) 130M
(b) 350M
Figure 5: Learning-rate sweeps for LLaMA2-130M and LLaMA2-350M trained on C4 for 2.6B and 7.5B tokens, respectively.
(a) Step 1%
(b) Step 10%
(c) Step 50%
(d) Step 100%
Figure 6: Singular-value entropy and stable rank of Muon momentum matrices under different normalization schemes during LLaMA2-130M training on C4 for 2.6B tokens. RowColNorm consistently improves both metrics over Direct, ColNorm, and RowNorm.
(a) Step 1%
(b) Step 10%
(c) Step 50%
(d) Step 100%
Figure 7: Layerwise Muon NS5 bias decomposition under different normalization schemes during LLaMA2-130M training on C4 for 2.6B tokens. The top panel reports $\|\mathrm{NS}_{5}(S(\mathbf{M}))-\mathrm{Orth}(S(\mathbf{M}))\|_{F}$, while the bottom panel reports $\|\mathrm{Orth}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\|_{F}$, both evaluated across layers at the same training step. These plots reveal how the choice of normalization shifts the error budget between the finite-step approximation error term and the preconditioning bias term.
Appendix J Experimental Details
J.1 Pretraining on C4
Table 4: Hyperparameters used for training LLaMA2-130M on C4
Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq
Max Learning Rate 5e-3 3e-3 3e-3 3e-3 3e-3
Warmup Steps 500
Batch Size 128
Maximum Length 4096
Weight Decay 0.1
$(\beta_{1},\beta_{2})$ (0.9,0.95)
Muon-Momentum 0.95
Gamma 0.025 0.05
Table 5: Hyperparameters used for training LLaMA2-350M on C4
Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq
Max Learning Rate 1.5e-3 2e-3 1.5e-3 2e-3 2e-3
Warmup Steps 500
Batch Size 256
Maximum Length 4096
Weight Decay 0.1
$(\beta_{1},\beta_{2})$ (0.9,0.95)
Muon-Momentum 0.95
Gamma 0.025 0.05
[…]
J.2 Ablation Study Details
Table 6: Hyperparameters used for training ResNet-18 on CIFAR10
Hyper-parameter SGD AdamW Muon-Nes MuonEq (RC) MuonEq (C) MuonEq (R)
Max Learning Rate 5e-2 1e-3 5e-2 5e-2 5e-2 5e-2
Warmup Steps 0
Epochs 100
Batch Size 128
$(\beta_{1},\beta_{2})$ (0.9,0.999) (0.9,0.95)
Muon-Momentum 0.9
Table 7: Hyperparameters used for training GPT2-small on FineWeb
Hyper-parameter AdamW Muon-Nes MuonEq (RC) MuonEq (C) MuonEq (R)
Max Learning Rate 2e-3 2e-3 2e-3 2e-3 2e-3
Warmup Steps 1000
Total Steps 20000
Batch Size 128
Maximum Length 4096
Weight Decay 0.1
$(\beta_{1},\beta_{2})$ (0.9,0.95)
Muon-Momentum 0.95
Gamma 0.05
[…]
We compare the three MuonEq variants RC, R, and C. RC applies two-sided row/column normalization throughout training, R applies row normalization throughout training, and C applies column normalization throughout training. We take R as the default variant. In all cases, the orthogonalization routine, optimizer states, and scalar hyperparameters are unchanged.
The purpose of this ablation is to isolate the effect of pre-orthogonalization geometry rather than to redesign the optimizer for each variant. We therefore use the same training budget and evaluation protocol across all runs. On CIFAR-10 we report final test accuracy. On FineWeb/GPT2-small we report validation perplexity at the end of the 10.5B-token budget.
This ablation is directly connected to Eq. (5). RC provides the strongest two-sided spectral correction and mainly targets the finite-step Newton–Schulz term. R and C test the two one-sided geometries under the same optimizer configuration. Since MuonEq targets hidden matrix weights, R is the more relevant default in our setting, while C is included as the column-sided companion form.
J.3 A Note on Row/Column Terminology in Prior Work
A careful comparison to prior row/column normalization papers requires fixing the matrix convention. [13] write linear weights as $\theta\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ and define column-wise normalization along the output dimension. Under the standard deep-learning storage layout $\mathbf{W}=\theta^{\top}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$, this corresponds to row normalization of the stored hidden-weight matrix. Their empirical observation that row-wise normalization can be unstable is also traced largely to the LM-head, where row-wise normalization produces extreme gradient values. This is not the same setting as MuonEq, which is aimed at hidden matrix weights.
By contrast, [31] use the standard convention $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$, so their ColNorm is genuinely column-wise in the same orientation as ours. Their main recommendation for language models is first-layer ColNorm, hidden-layer Spectral, and last-layer Sign. We therefore interpret our row-sided default choice as directly aligned with [49], compatible with [13] after accounting for layout conventions, and not in tension with [31] once first-layer/input geometry is separated from hidden-layer geometry.

### 10 May 2026 v2

@@
 Abstract
-Orthogonalized-update optimizers such as Muon improve training of matrix-valued parameters, but existing extensions mostly act either after orthogonalization by rescaling updates or before it with heavier whitening-based preconditioners.
-We introduce MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon in three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C).
-These variants rebalance the momentum matrix before finite-step Newton–Schulz using row/column squared-norm statistics and only $\mathcal{O}(m+n)$ auxiliary state.
-We show that finite-step orthogonalization is governed by input spectral properties, especially stable rank and condition number, and that row/column normalization is a zeroth-order whitening surrogate that removes marginal scale mismatch.
-For the hidden matrix weights targeted by MuonEq, the row-normalized variant R is the natural default and preserves the $\widetilde{\mathcal{O}}(T^{-1/4})$ stationarity guarantee of Muon-type methods.
-In LLaMA2 pretraining on C4, the default R variant consistently outperforms Muon on 130M and 350M models, yielding faster convergence and lower validation perplexity.
+Orthogonalized-update optimizers such as Muon improve training of matrix-valued parameters, but existing extensions typically either rescale updates after orthogonalization or use heavier whitening-based preconditioners before it.
+We introduce MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon with three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C).
+By rebalancing the momentum matrix before finite-step Newton–Schulz orthogonalization, MuonEq improves the geometry seen by orthogonalization.
+We show that finite-step orthogonalization is governed by the input spectrum, especially stable rank and condition number, and that row/column normalization acts as a zeroth-order surrogate for whitening.
+For hidden matrix weights, R is the default variant.
+Theoretically, MuonEq (R) retains the standard $\widetilde{\mathcal{O}}(T^{-1/4})$ Muon-type nonconvex stationarity guarantee with decoupled weight decay and a horizon-free diminishing learning-rate schedule, and extends it to finite-step NS5 up to an explicit inexactness constant.
+In LLaMA2 pretraining on C4, MuonEq (R) consistently outperforms Muon on 130M, 350M, and 1B models, with faster convergence and lower validation perplexity.
+The code is available at the MuonEq codebase.
 1 Introduction
 As the computational and data costs of large language model pretraining grow, optimization increasingly determines sample efficiency, throughput, and scalability.
-Beyond coordinate-wise adaptive methods such as AdamW (Loshchilov and Hutter, 2019), Adan (Xie et al., 2024), and MARS (Liu et al., 2025b), recent work has shifted toward a geometric view of matrix-valued updates.
+Beyond coordinate-wise adaptive methods such as AdamW (Loshchilov and Hutter, 2019), Adan (Xie et al., 2024), and MARS (Yuan et al., 2024), recent work has shifted toward a geometric view of matrix-valued updates.
 Muon (Jordan et al., ), for instance, uses a Newton–Schulz approximation to the polar factor of the momentum matrix to produce spectrally aligned updates (Bernstein and Newhouse, 2024; Chen et al., 2025).
-It has shown strong empirical scalability in LLM pretraining, often improving the trade-off among data efficiency, compute, and wall-clock time at fixed target loss (Liu et al., 2025a; Shah et al., 2025).
+It has shown strong empirical scalability in LLM pretraining, often improving the trade-off among data efficiency, compute, and wall-clock time at fixed target loss (Liu et al., 2025a; Shah et al., 2025; DeepSeek-AI, 2026).
 Recent theory further characterizes Muon as spectral-norm steepest descent and establishes nonconvex convergence guarantees for both exact and approximate orthogonalization (Chang et al., 2025; Sato et al., 2025; Li and Hong, 2025; Shen et al., 2025; Sfyraki and Wang, 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025).
@@
 SOAP (Vyas et al., 2024) connects Shampoo (Gupta et al., 2018) with Adam/Adafactor (Shazeer and Stern, 2018; Kingma and Ba, 2015) through a preconditioned basis, while FISMO (Xu et al., 2026a) and Mousse (Zhang et al., 2026c) further perform spectral optimization in whitened coordinates.
-These methods suggest that improving geometry before orthogonalization is the more direct lever, but doing so usually requires maintaining left and right preconditioners together with extra matrix operations such as eigendecomposition, inverse power iteration, and whitening/dewhitening maps.
+These methods motivate studying geometry correction before orthogonalization as a promising lever, but existing approaches typically incur substantially higher state and computational overhead.
 These works motivate a more precise question:
@@
 This makes diagonal equilibration a natural lightweight way to improve the matrix presented to Muon-style orthogonalization.
-Based on this observation, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon.
-MuonEq has three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C).
-All three act directly on the momentum matrix before Newton–Schulz using only row/column squared-norm statistics and $\mathcal{O}(m+n)$ auxiliary state.
-Detailed justification is deferred to Section 3: Theorem 3.1 quantifies the spectral sensitivity of finite-step orthogonalization, Proposition 3.2 shows that row/column normalization is a zeroth-order surrogate for whitening, and Proposition 3.4 together with Theorem 3.5 clarify the trade-off between stronger two-sided equilibration and cleaner one-sided geometry.
-For the hidden Transformer weight matrices targeted by MuonEq, recent architecture-aware analyses by (Xu et al., 2026b) identify row-sided geometry as the natural one-sided choice, whereas column-sided geometry is more closely associated with embedding matrices; accordingly, we adopt R as the default and treat RC and C as companion forms for analysis and ablation.
-Compared with Muon+ (Zhang et al., 2026b), which rescales the update after orthogonalization, MuonEq modifies the matrix fed into orthogonalization.
-Unlike whitening-based methods such as FISMO and Mousse (Xu et al., 2026a; Zhang et al., 2026c), it avoids Kronecker factors, matrix-valued second-order statistics, eigendecomposition, and momentum transport, and requires only $\mathcal{O}(m+n)$ auxiliary state.
+Motivated by this observation, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon, instantiated as two-sided row/column normalization (RC), row normalization (R), and column normalization (C).
+All variants rescale the momentum matrix before Newton–Schulz, using row/column squared-norm reductions from the current momentum.
+Section 3 provides the theory: Theorem 3.1 quantifies the spectral sensitivity of finite-step orthogonalization; Proposition 3.2 shows that row/column normalization is a zeroth-order whitening surrogate; and Proposition 3.4, Theorem 3.5, and Corollary 3.6 formalize the RC/R trade-off.
+In particular, the default R variant satisfies the standard $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex stationarity guarantee while explicitly accounting for decoupled weight decay, horizon-free learning-rate schedules, and finite-step NS5.
+For the hidden Transformer weight matrices targeted by MuonEq, recent architecture-aware analyses (Xu et al., 2026b) identify row-sided geometry as the natural one-sided choice, while column-sided geometry is more aligned with embeddings; hence we use R by default and keep RC/C for analysis and ablation.
+Compared with Muon+ (Zhang et al., 2026b), which rescales updates after orthogonalization, MuonEq changes the input to orthogonalization.
+Unlike whitening-based methods such as FISMO and Mousse (Xu et al., 2026a; Zhang et al., 2026c), MuonEq avoids Kronecker factors, matrix-valued second-order statistics, eigendecomposition, and momentum transport; its diagonal rescaling is computed on the fly from row/column squared norms, with no persistent optimizer state.
 Our contributions are threefold.
-  1. 1.
-    First, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon with three forms: RC, R, and C; these variants rebalance the matrix before Newton–Schulz using only row/column squared-norm statistics, with R as the default for hidden matrix weights.
-  2. 2.
-    Second, we show that finite-step Newton–Schulz is governed by the input spectrum, especially stable rank and condition number; that row/column normalization acts as a zeroth-order surrogate for whitening; and that RC and R admit a clear trade-off between stronger spectral correction and cleaner one-sided convergence structure, with the default R variant satisfying a $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex guarantee under standard assumptions.
-  3. 3.
-    Third, empirically, the default R variant consistently outperforms Muon across model sizes and token budgets, while ablations over RC, R, and C show that R is the strongest default form for the hidden-weight setting targeted by MuonEq.
+(i) First, we propose MuonEq, a lightweight family of pre-orthogonalization equilibration schemes for Muon with three forms: RC, R, and C; these variants rebalance the current momentum matrix before Newton–Schulz using row/column squared norms computed on the fly, with R as the default for hidden matrix weights.
+(ii) Second, we analyze how pre-orthogonalization geometry affects finite-step Newton–Schulz: Theorem 3.1 links its behavior to stable rank and condition number, Proposition 3.2 identifies row/column normalization as a zeroth-order whitening surrogate, and Proposition 3.4 together with Theorem 3.5 and Corollary 3.6 formalize the trade-off between stronger two-sided spectral correction and cleaner one-sided geometry.
+In particular, MuonEq (R) satisfies the standard $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex stationarity guarantee while accounting for decoupled weight decay, horizon-free learning-rate schedules, and finite-step NS5.
+(iii) Third, empirically, the default R variant consistently outperforms Muon across model sizes and token budgets, while ablations over RC, R, and C show that R is the default form for the hidden-weight setting targeted by MuonEq.
 2 Method
@@
 The model parameter is a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, and, unless otherwise stated, we assume $m\geq n$.
-For any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ with rank $r\leq n$, we write its compact singular value decomposition as $\mathbf{A}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})$ and $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$, and define $\operatorname{Orth}(\mathbf{A}):=\mathbf{U}\mathbf{V}^{\top}$.
+For any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ with rank $r\leq n$, we write its compact singular value decomposition as $\mathbf{A}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r})$ and $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$, and define the polar factor, which we also call the orthogonalized form, as $\mathrm{Orth}(\mathbf{A}):=\mathbf{U}\mathbf{V}^{\top}$.
+Here, orthogonalization and polar step mean applying or approximating this map, whereas whitening denotes full Gram inverse-square-root preconditioning.
+The row and column rescalings introduced below are diagonal equilibration operations.
 In finite-data form, the training objective can be written as $f(\mathbf{X}):=\frac{1}{N}\sum_{i\in[N]}f_{i}(\mathbf{X})$, where $f_{i}(\mathbf{X})$ is the loss associated with the $i$-th sample $\mathbf{z}_{i}$.
@@
 where $f$ may be nonconvex, $\xi$ is independent of $\mathbf{X}$, and $\mathbb{E}_{\xi}[\cdot]$ denotes expectation with respect to $\xi$.
-For any matrix $\mathbf{A}$ with singular values $\{\sigma_{i}(\mathbf{A})\}_{i=1}^{r}$, we further use $\operatorname{sr}(\mathbf{A}):=\frac{\|\mathbf{A}\|_{F}^{2}}{\|\mathbf{A}\|_{2}^{2}},\kappa_{2}(\mathbf{A}):=\frac{\sigma_{1}(\mathbf{A})}{\sigma_{r}(\mathbf{A})},p_{i}:=\frac{\sigma_{i}(\mathbf{A})^{2}}{\|\mathbf{A}\|_{F}^{2}},H(p):=-\sum_{i=1}^{r}p_{i}\log p_{i}.$
+For any matrix $\mathbf{A}$ with singular values $\{\sigma_{i}(\mathbf{A})\}_{i=1}^{r}$, we further use $\operatorname{sr}(\mathbf{A}):=\frac{\|\mathbf{A}\|_{F}^{2}}{\|\mathbf{A}\|_{2}^{2}},\kappa(\mathbf{A}):=\frac{\sigma_{1}(\mathbf{A})}{\sigma_{r}(\mathbf{A})},\kappa_{i}(\mathbf{A}):=\frac{\sigma_{1}(\mathbf{A})}{\sigma_{i}(\mathbf{A})},p_{i}:=\frac{\sigma_{i}(\mathbf{A})^{2}}{\|\mathbf{A}\|_{F}^{2}},H(p):=-\sum_{i=1}^{r}p_{i}\log p_{i}.$
 Muon as finite-step orthogonalized momentum.
@@
 The abstract Muon update is
-$\mathbf{U}_{t}=\operatorname{Orth}(\tilde{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}\mathbf{U}_{t},$ (2)
-where $\tilde{\mathbf{M}}_{t}$ is either the standard momentum or its Nesterov variant (Jordan et al., ; Liu et al., 2025a; Chang et al., 2025).
-In practice, $\operatorname{Orth}(\tilde{\mathbf{M}}_{t})$ is implemented by a small fixed number of Newton–Schulz iterations rather than exact polar decomposition (Jordan et al., ).
-The design question is therefore not only what direction is desirable in the exact limit, but also what matrix should be fed into finite-step Newton–Schulz.
+$\mathbf{O}_{t}=\operatorname{Orth}(\widetilde{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t},$ (2)
+where $\widetilde{\mathbf{M}}_{t}$ is the standard momentum or its Nesterov variant (Jordan et al., ; Liu et al., 2025a; Chang et al., 2025); following LLM Muon practice (Liu et al., 2025a), we set $a=0.2\sqrt{\max\{m,n\}}$.
+In practice, $\operatorname{Orth}(\widetilde{\mathbf{M}}_{t})$ is implemented by a fixed number of Newton–Schulz iterations rather than exact polar decomposition (Jordan et al., ).
 Pre-orthogonalization equilibration family. MuonEq inserts diagonal preconditioning before orthogonalization:
-$\hat{\mathbf{M}}_{t}=\textbf{DiagPre}(\tilde{\mathbf{M}}_{t};s),\qquad\mathbf{U}_{t}=\operatorname{Orth}(\hat{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}\mathbf{U}_{t},$ (3)
+$\hat{\mathbf{M}}_{t}=\textbf{DiagPre}(\widetilde{\mathbf{M}}_{t},s,\varepsilon),\qquad\mathbf{O}_{t}=\operatorname{Orth}(\hat{\mathbf{M}}_{t}),\qquad\mathbf{X}_{t+1}=\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t},$ (3)
 where the mode $s\in\{\mathrm{RC},\mathrm{R},\mathrm{C}\}$ selects one of three equilibration forms. Let
-$\mathbf{D}_{r,t}=\operatorname{diag}\bigl(\operatorname{rowsum}(\tilde{\mathbf{M}}_{t}\odot\tilde{\mathbf{M}}_{t})+\varepsilon\bigr),\qquad\mathbf{D}_{c,t}=\operatorname{diag}\bigl(\operatorname{colsum}(\tilde{\mathbf{M}}_{t}\odot\tilde{\mathbf{M}}_{t})+\varepsilon\bigr).$
+$\mathbf{D}_{r,t}=\operatorname{diag}\bigl(\operatorname{rowsum}(\widetilde{\mathbf{M}}_{t}\odot\widetilde{\mathbf{M}}_{t})+\varepsilon\bigr),\qquad\mathbf{D}_{c,t}=\operatorname{diag}\bigl(\operatorname{colsum}(\widetilde{\mathbf{M}}_{t}\odot\widetilde{\mathbf{M}}_{t})+\varepsilon\bigr).$
 We consider the following three maps:
-$\textbf{DiagPre}(\tilde{\mathbf{M}}_{t};s)=\begin{cases}\mathbf{D}_{r,t}^{-1/2}\tilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{RC},\\ \mathbf{D}_{r,t}^{-1/2}\tilde{\mathbf{M}}_{t},&s=\mathrm{R},\\ \tilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{C}.\end{cases}$ (4)
-Since these maps use only row and column squared norms, the extra state is $\mathcal{O}(m+n)$, in the factorized spirit of Adafactor (Shazeer and Stern, 2018).
-Unlike methods that rely on matrix-valued or adaptive preconditioners (Gupta et al., 2018; Shazeer and Stern, 2018; Kingma and Ba, 2015; Xu et al., 2026a), MuonEq changes only the matrix fed to Newton–Schulz.
-Relative to Muon (Jordan et al., ), the orthogonalization routine itself is unchanged; relative to Muon+, NorMuon, and AdaMuon (Zhang et al., 2026b; Li et al., 2025; Si et al., 2025), MuonEq rebalances the input rather than rescales the output.
+$\textbf{DiagPre}(\widetilde{\mathbf{M}}_{t},s,\varepsilon)=\begin{cases}\mathbf{D}_{r,t}^{-1/2}\widetilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{RC},\\ \mathbf{D}_{r,t}^{-1/2}\widetilde{\mathbf{M}}_{t},&s=\mathrm{R},\\ \widetilde{\mathbf{M}}_{t}\mathbf{D}_{c,t}^{-1/2},&s=\mathrm{C}.\end{cases}$ (4)
+Unlike methods that rely on matrix-valued or adaptive preconditioners (Gupta et al., 2018; Shazeer and Stern, 2018; Kingma and Ba, 2015; Xu et al., 2026a), MuonEq changes only the transient input fed to Newton–Schulz.
+Relative to Muon (Jordan et al., ), the orthogonalization routine and persistent optimizer state are unchanged; relative to Muon+, NorMuon, and AdaMuon (Zhang et al., 2026b; Li et al., 2025; Si et al., 2025), MuonEq rebalances the input rather than rescales the output.
 Why the default variant is R. This choice is block-dependent rather than a universal preference for row normalization. The trade-off is guided by
-$\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}\leq$ $\underbrace{\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(S(\mathbf{M}))\right\|_{F}}_{\text{finite-step approximation error}}$ (5)
-$+\underbrace{\left\|\mathrm{Orth}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}}_{\text{preconditioning bias}},$
+$\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}\leq\underbrace{\left\|\mathrm{NS}_{K}(S(\mathbf{M}))-\mathrm{Orth}(S(\mathbf{M}))\right\|_{F}}_{\text{finite-step approximation error}}+\underbrace{\left\|\mathrm{Orth}(S(\mathbf{M}))-\mathrm{Orth}(\mathbf{M})\right\|_{F}}_{\text{preconditioning bias}},$ (5)
 where $S$ is one of RC, R, or C.
@@
 Remark 2.1.
-Unless otherwise stated, we use the row-normalized variant $s=\mathrm{R}$. The RC and C modes are kept as equally lightweight companion forms for analysis and ablation.
+Unless otherwise stated, we use $s=\mathrm{R}$; $\mathrm{RC}$ and $\mathrm{C}$ are retained for analysis and ablation.
+For varying $\beta_{t}$, the Nesterov line follows the EMA-buffer convention; for $\beta_{t}\equiv\beta$, it reduces to the Nesterov form Yuan et al.
+(2024); Chang et al.
+(2025).
 3 Analysis
 The analysis follows the logic of MuonEq: Section 3.1 establishes the spectrum sensitivity of finite-step Newton–Schulz; Section 3.2 interprets row/column normalization as a zeroth-order surrogate for whitening; and Section 3.3 motivates the R rather than RC, with row normalization admitting a clean stochastic nonconvex guarantee.
-Together, Theorem 3.1, Proposition 3.2, Proposition 3.4, and Theorem 3.5 justify the MuonEq family and the default choice of R.
+Together, these results motivate the MuonEq family and clarify the trade-off between stronger spectral correction and preprocessing bias.
+Within the hidden-weight setting studied here, they support using R as our default working variant.
 3.1 Why spectral geometry matters for finite-step Newton–Schulz orthogonalization
@@
 Theorem 3.1.
-Let $\mathbf{G}\in\mathbb{R}^{p\times q}$ have rank $r\geq 1$, with compact SVD $\mathbf{G}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, and suppose $p\leq q$.
-Fix $\alpha\geq\|\mathbf{G}\|_{2}$, set $\mathbf{X}_{0}=\alpha^{-1}\mathbf{G}$, and define $\mathbf{X}_{k+1}=\left(a\mathbf{I}_{p}+b\mathbf{X}_{k}\mathbf{X}_{k}^{\top}+c(\mathbf{X}_{k}\mathbf{X}_{k}^{\top})^{2}\right)\mathbf{X}_{k}$ for $k\in\mathbb{N}$, where $a=3.4445$, b=4.7750b=-4.7750, and $c=2.0315$ are the coefficients used in Muon’s Newton–Schulz iteration in (Jordan et al., ).
-Let $\phi(s)=as+bs^{3}+cs^{5}$.
-Then, for each $k\in\mathbb{N}$, $\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}(s_{1}^{(k)},\ldots,s_{r}^{(k)})\mathbf{V}^{\top},s_{i}^{(0)}=\sigma_{i}/\alpha,s_{i}^{(k+1)}=\phi\bigl(s_{i}^{(k)}\bigr),i\in[r],$ and hence
-$\frac{1}{\sqrt{r}}\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\bigl(1-a^{k}\sigma_{i}/\alpha\bigr)_{+}^{2}\right)^{1/2}.$
-Here $(x)_{+}:=\max\{x,0\}$.
-See Appendix A for details.
+Let $\mathbf{G}\in\mathbb{R}^{p\times q}$ have rank $r\geq 1$ and compact SVD $\mathbf{G}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ with $\sigma_{1}\geq\cdots\geq\sigma_{r}>0$ and $p\leq q$.
+Consider the Newton–Schulz iteration
+$\mathbf{X}_{0}=\frac{\mathbf{G}}{\|\mathbf{G}\|_{F}},\quad\mathbf{X}_{k+1}=\left(a\mathbf{I}_{p}+b\mathbf{X}_{k}\mathbf{X}_{k}^{\top}+c(\mathbf{X}_{k}\mathbf{X}_{k}^{\top})^{2}\right)\mathbf{X}_{k},$
+where $\phi(s):=as+bs^{3}+cs^{5}$ and $q(t):=a+bt+ct^{2}$ satisfy $a>1,0<q(t)\leq a,t\in[0,1].$ Such coefficient choices include those used in (Jordan et al., ; Gyu Yeol Kim, 2026).
+Then, for each $k\geq 0$,
+$\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}\bigl(s_{1}^{(k)},\ldots,s_{r}^{(k)}\bigr)\mathbf{V}^{\top},\qquad s_{i}^{(0)}=\frac{\sigma_{i}}{\|\mathbf{G}\|_{F}},\qquad s_{i}^{(k+1)}=\phi\bigl(s_{i}^{(k)}\bigr).$
+Moreover,
+$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\left(1-\frac{a^{k}}{\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}}\right)_{+}^{2}\right)^{1/2}.$
+Equivalently, the $i$-th singular direction leaves the linear regime at $\tau_{i}=\log_{a}\left(\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}\right)$ .
+In particular, $\tau_{1}=\log_{a}\sqrt{\operatorname{sr}(\mathbf{G})},\tau_{r}=\log_{a}\left(\kappa(\mathbf{G})\sqrt{\operatorname{sr}(\mathbf{G})}\right),\tau_{r}-\tau_{1}=\log_{a}\kappa(\mathbf{G})$ .
+See Appendix B for details.
 Remark 3.1.
-Under the normalization $\alpha=\|\mathbf{G}\|_{F}$, which is common in practical Newton–Schulz implementations (Jordan et al., ; Liu et al., 2025a), let $p_{i}:=\frac{\sigma_{i}^{2}}{\|\mathbf{G}\|_{F}^{2}},\tau_{i}:=\log_{a}\frac{\|\mathbf{G}\|_{F}}{\sigma_{i}}=\log_{a}\frac{1}{\sqrt{p_{i}}}.$ Then Theorem 3.1 implies
-$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\bigl(1-a^{k-\tau_{i}}\bigr)_{+}^{2}\right)^{1/2}.$
-Thus $\tau_{i}$ is precisely the iteration at which the $i$-th singular direction leaves the linear regime.
-In particular, $\tau_{1}=\log_{a}\sqrt{\operatorname{sr}(\mathbf{G})},\tau_{r}=\log_{a}\bigl(\kappa_{2}(\mathbf{G})\sqrt{\operatorname{sr}(\mathbf{G})}\bigr),\tau_{r}-\tau_{1}=\log_{a}\kappa_{2}(\mathbf{G}).$ Therefore, the stable rank determines when the first singular direction enters the nonlinear regime, whereas the condition number determines the width of the transition window over which the remaining singular directions follow.
-Theorem 3.1 shows that broad spectra slow finite-step orthogonalization by delaying contraction along small singular directions.
+Theorem 3.1 shows that the stable rank controls the onset of the linear to nonlinear transition, while the condition number controls its width across singular directions.
+Refer to captionRefer to caption
+((a))
+Refer to captionRefer to caption
+((b))
+Refer to captionRefer to caption
+((c))
+Refer to captionRefer to caption
+((d))
+Figure 1: Random Gaussian matrices with controlled shapes and spectral spreads.
+Top: finite-step relative Frobenius error to the exact polar factor across Newton–Schulz steps.
+Bottom: raw and post-normalization condition numbers.
+Two-sided row/column normalization yields the smallest error and the most consistent spectral compression.
 To isolate this effect, Figure 1 uses random Gaussian matrices with controlled shapes and spectral spreads.
@@
 The top row shows the module-wise median of the error in Theorem 3.1, and the bottom row shows the corresponding mean.
-(a) $\times1024$
-[…]
-[…]
-(d) $\times4096$
-(e) $\times1024$
-[…]
-[…]
-(h) $\times4096$
-Figure 1: Random Gaussian matrices with controlled shapes and spectral spreads.
-Top: finite-step relative Frobenius error to the exact polar factor across Newton–Schulz steps.
-Bottom: raw and post-normalization condition numbers.
-Two-sided row/column normalization yields the smallest error and the most consistent spectral compression.
-(a) Step 1%
-(b) Step 10%
-(c) Step 50%
-(d) Step 100%
-(e) Step 1%
-(f) Step 10%
-(g) Step 50%
-(h) Step 100%
+Appendix J provides a more detailed module-wise view of the same phenomenon: Figure 6 reports singular-value entropy and stable-rank diagnostics for the pre-NS momentum matrices, showing that equilibration reshapes the spectrum seen by finite-step Newton–Schulz.
+Refer to captionRefer to caption
+((a))
+Refer to captionRefer to caption
+((b))
+Refer to captionRefer to caption
+((c))
+Refer to captionRefer to caption
+((d))
 Figure 2: Finite-step orthogonalization error across Newton–Schulz steps at 1%, 10%, 50%, and 100% of training.
-Top: module-wise median; bottom: module-wise mean; shaded bands denote the 25%–75% range.
+In each column, the top panel shows the module-wise median and the bottom panel shows the module-wise mean; shaded bands denote the 25%–75% range.
 Two-sided row/column normalization decays fastest.
@@
 Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full column rank, and set $\mathbf{D}_{c}:=\operatorname{diag}(\|\mathbf{M}_{:,1}\|_{2},\ldots,\|\mathbf{M}_{:,q}\|_{2})\succ 0$, $\mathbf{N}_{c}:=\mathbf{M}\mathbf{D}_{c}^{-1}$, and $\mathbf{C}_{c}:=\mathbf{N}_{c}^{\top}\mathbf{N}_{c}-\mathbf{I}_{q}$.
-If $\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}$ is sufficiently small, then
-$\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{M}(\mathbf{M}^{\top}\mathbf{M})^{-1/2}=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}\bigl(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}\bigr),$
-where $\mathbf{L}_{c}$ is the unique solution of $\mathbf{D}_{c}\mathbf{L}_{c}+\mathbf{L}_{c}\mathbf{D}_{c}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$ In particular,
+If $\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{r}(\mathbf{M}):=\mathbf{M}(\mathbf{M}^{\top}\mathbf{M})^{-1/2}=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}\bigl(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}\bigr),$ where $\mathbf{L}_{c}$ is the unique solution of $\mathbf{D}_{c}\mathbf{L}_{c}+\mathbf{L}_{c}\mathbf{D}_{c}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$ In particular,
 $\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}+\mathcal{O}\bigl(\|\mathbf{C}_{c}\|_{2}\bigr).$
@@
 Let $\mathbf{M}\in\mathbb{R}^{p\times q}$ have full row rank, and set $\mathbf{D}_{r}:=\operatorname{diag}(\|\mathbf{M}_{1,:}\|_{2},\ldots,\|\mathbf{M}_{p,:}\|_{2})\succ 0$, $\mathbf{N}_{r}:=\mathbf{D}_{r}^{-1}\mathbf{M}$, and $\mathbf{C}_{r}:=\mathbf{N}_{r}\mathbf{N}_{r}^{\top}-\mathbf{I}_{p}$.
-If $\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}$ is sufficiently small, then
-$\operatorname{Orth}_{\ell}(\mathbf{M})=(\mathbf{M}\mathbf{M}^{\top})^{-1/2}\mathbf{M}=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}\bigr),$
-where $\mathbf{L}_{r}$ is the unique solution of $\mathbf{D}_{r}\mathbf{L}_{r}+\mathbf{L}_{r}\mathbf{D}_{r}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$ In particular,
+If $\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{\ell}(\mathbf{M}):=(\mathbf{M}\mathbf{M}^{\top})^{-1/2}\mathbf{M}=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}\bigr),$ where $\mathbf{L}_{r}$ is the unique solution of $\mathbf{D}_{r}\mathbf{L}_{r}+\mathbf{L}_{r}\mathbf{D}_{r}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$ In particular,
 $\operatorname{Orth}_{\ell}(\mathbf{M})=\mathbf{N}_{r}+\mathcal{O}\bigl(\|\mathbf{C}_{r}\|_{2}\bigr).$
-See Appendix C for details.
+See Appendix D for details.
 Corollary 3.3.
@@
 Thus row normalization separates marginal rescaling from the genuine whitening step.
-The column-normalized/right analogue and the corresponding two-sided normalization statements are completely symmetric; see Appendix D.
+The column-normalized/right analogue and the corresponding two-sided normalization statements are completely symmetric; see Appendix E.
 Together, Proposition 3.2, Corollary 3.3, and Remark 3.2 explain the division of labor in MuonEq: normalization removes marginal-scale mismatch at zeroth order, and the subsequent Newton–Schulz orthogonalization handles the remaining correlation structure.
@@
 3.3 Convergence analysis of the RC and R variants
-To analyze the MuonEq family, we focus on the two variants that are most informative theoretically: the two-sided RC map and the default row-normalized R map.
-RC is the strongest spectral corrector among the three forms but induces a more involved alignment term, whereas R yields the cleanest one-sided geometry and convergence argument.
-Accordingly, Proposition 3.4 studies the pure RC variant, and Theorem 3.5 provides the main guarantee for the default R variant.
-The column-normalized C map is retained as the embedding-aligned companion in the method family and empirical ablations.
+To analyze the MuonEq family, we focus on the two variants that are most informative theoretically: RC and the default row-normalized R.
+RC gives stronger two-sided spectral correction but leads to a more involved alignment term, while R has a cleaner one-sided geometry.
+Accordingly, Proposition 3.4 studies RC.
+For R, Theorem 3.5 gives the main guarantee for the exact-polar update, and Corollary 3.6 extends it to the finite-step NS5 implementation.
+Compared with recent Muon analyses Chang et al.
+(2025); Sato et al.
+(2025); Shen et al.
+(2025); Gyu Yeol Kim (2026); Shulgin et al.
+(2025); Nagashima and Iiduka (2026), our contribution is not to improve the asymptotic rate, but to establish a guarantee closer to the optimizer setting used in practice: the same $\widetilde{\mathcal{O}}(T^{-1/4})$ stationarity rate is proved while explicitly covering decoupled weight decay, a learning-rate schedule independent of the training horizon, and finite-step NS5 through the inexactness constant $\varepsilon_{\mathrm{ns}}$.
 Assumption 3.1.
@@
 Specifically, there exists $\sigma>0$ such that, for all $\mathbf{X}\in\mathbb{R}^{m\times n}$, $\mathbb{E}[\nabla f(\mathbf{X};\xi)]=\nabla f(\mathbf{X}),\mathbb{E}\|\nabla f(\mathbf{X};\xi)-\nabla f(\mathbf{X})\|_{F}^{2}\leq\sigma^{2}.$
-These assumptions are standard in analyses of adaptive and Muon-type methods (Xie et al., 2024; Chen et al., 2025; Li et al., 2023; Wang et al., 2023; Chang and Yuan, 2025; Huang et al., 2025).
+These assumptions are standard in analyses of adaptive and Muon-type methods (Xie et al., 2024; Chen et al., 2025; Li et al., 2023; Wang et al., 2023; Chang and Yuan, 2026; Li et al., 2026; Huang et al., 2025).
 Proposition 3.4.
-Assume Assumptions 3.1, 3.2, and $3.$ Moreover, assume there exists $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty}$ almost surely.
-Consider the RC variant (4) with $\lambda=0$ and Nesterov disabled, so that $\hat{\mathbf{M}}_{t}=\mathbf{D}_{r,t}^{-1/2}\mathbf{M}_{t}\mathbf{D}_{c,t}^{-1/2}$ for all $t\in[T]$.
 Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, and $a:=0.2\sqrt{\max\{m,n\}}$.
+Assume Assumptions 3.1, 3.2, and 3.3.
+Moreover, assume there exists $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty}$ almost surely.
+For Algorithm 1, consider the exact-polar RC variant in (4), with Nesterov momentum disabled and $\lambda_{t}=0$.
 Suppose further that $\varepsilon\geq\frac{4}{5}G_{\infty}^{2}\max\{m,n\}$, and define $\rho_{r}:=\sqrt{1+\frac{nG_{\infty}^{2}}{\varepsilon}},\rho_{c}:=\sqrt{1+\frac{mG_{\infty}^{2}}{\varepsilon}},\chi_{\varepsilon}:=\frac{(\rho_{r}+1)(\rho_{c}+1)}{4\rho_{r}\rho_{c}}-\frac{3\rho_{r}\rho_{c}-\rho_{r}-\rho_{c}-1}{4}.$ Then $\rho_{r},\rho_{c}\leq 3/2$ and hence $\chi_{\varepsilon}\geq 1/144$.
@@
 $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}(1+\ln T)+C_{2}}{a\chi_{\varepsilon}\cdot T^{1/4}},$
-where $C_{1}=\frac{a}{L}\bigl(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}\bigr)+aL\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2},C_{2}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}n}{2}.$
+where $C_{1}=\frac{a}{L}\bigl(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}\bigr)+aL\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2},C_{2}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}n}{2}.$ See Appendix G for details.
 Theorem 3.5.
-Assume Assumptions 3.1, 3.2, and $3.$ Consider the practical row-normalized R variant (4), with Nesterov disabled and $\lambda=0,\varepsilon=0$.
-Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, $a:=0.2\sqrt{\max\{m,n\}}$, and suppose the trajectory-wise NS5 polar-approximation error $\varepsilon_{\mathrm{ns}}$ defined in Lemma G.2 satisfies $\varepsilon_{\mathrm{ns}}<1$.
+Let $\eta_{t}=t^{-3/4}$, $\beta_{t}=1-t^{-1/2}$, $h_{t}:=1-\beta_{t}=t^{-1/2}$, $\varepsilon=0$ and $a:=0.2\sqrt{\max\{m,n\}}$.
+Assume Assumptions 3.1, 3.2, and 3.3.
+For Algorithm 1, consider the exact-polar R variant in (4), with Nesterov momentum disabled and $0\leq\rho<\frac{a}{\sqrt{m}},0\leq\lambda_{t}\leq\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a\sqrt{n}}\,t^{-1/4}$.
 Then, for every $T\geq 1$,
-$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}}{\bigl(a(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}\bigr)\cdot T^{1/4}},$
-where $C_{1}^{\mathrm{ns}}=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)+aL\Bigl(\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)^{2},C_{2}^{\mathrm{ns}}=\frac{a\sigma^{2}}{L}+\frac{3La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}.$
+$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ep}}(1+\ln T)+C_{2,\rho}^{\mathrm{ep}}}{\bigl(a/\sqrt{m}-\rho\bigr)\cdot T^{1/4}},$
+where $C_{1,\rho}^{\mathrm{ep}}=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}(a\sqrt{n}+\rho)^{2}+\sigma^{2}\Bigr)+aL\Bigl(\frac{1}{\sqrt{m}}+\sqrt{n}\Bigr)^{2},C_{2,\rho}^{\mathrm{ep}}=\frac{a\sigma^{2}}{L}+\frac{3L}{2}(a\sqrt{n}+\rho)^{2}.$
+Corollary 3.6.
+Under the assumptions of Theorem 3.5, replace the exact-polar step by the finite-step NS5 step $\mathbf{O}_{t}:=\mathrm{NS5}(\hat{\mathbf{M}}_{t}),$ where $\mathrm{NS5}$ is the five-step pre-scaled Newton–Schulz map in Lemma H.2.
+For the fixed horizon $T$, we define
+$\varepsilon_{\mathrm{ns}}:=\varepsilon_{\mathrm{ns},T}=\max_{1\leq t\leq T}\bigl\|\mathrm{NS5}(\hat{\mathbf{M}}_{t})-\operatorname{Orth}(\hat{\mathbf{M}}_{t})\bigr\|_{2}.$
+By Lemma H.2, $\varepsilon_{\mathrm{ns}}<1$.
+Replace the decoupled weight decay condition by $0\leq\rho<\frac{a(1-\varepsilon_{\mathrm{ns}})}{\sqrt{m}},0\leq\lambda_{t}\leq\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}}\,t^{-1/4}.$ Then, for every $T\geq 1$,
+$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ns}}(1+\ln T)+C_{2,\rho}^{\mathrm{ns}}}{\bigl(a(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}-\rho\bigr)\cdot T^{1/4}},$
+where $C_{1,\rho}^{\mathrm{ns}}=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}+\sigma^{2}\Bigr)+aL\Bigl(\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)^{2}$ , $C_{2,\rho}^{\mathrm{ns}}=\frac{a\sigma^{2}}{L}+\frac{3L}{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}$.
+See Appendix I for details.
 Remark 3.3.
-Setting $\varepsilon_{\mathrm{ns}}=0$ recovers the exact-polar theorem for the row-normalized R variant.
-More generally, compared with the two-sided RC analysis, the descent coefficient is now the fixed positive constant $(1-\varepsilon_{\mathrm{ns}})/\sqrt{m}$, so no auxiliary condition of the form $\chi_{\varepsilon}>0$ is needed.
+When $\varepsilon=0$, $\mathbf{D}_{r,t}^{-1/2}$ is interpreted as $(\mathbf{D}_{r,t}^{1/2})^{\dagger}$, leaving zero rows unchanged.
+Theorem 3.5 is the main exact-polar guarantee for the default R variant.
+Corollary 3.6 transfers the same $T^{-1/4}$ dependence to the NS5 implementation, up to the explicit $(1\pm\varepsilon_{\mathrm{ns}})$ constants.
+In Algorithm 1, enabling Nesterov momentum simply amounts to replacing $\mathbf{M}_{t}$ with $\widetilde{\mathbf{M}}t$ in both $\mathbf{D}{r,t}$ and $\hat{\mathbf{M}}_{t}$; this should not change the stated $T^{-1/4}$ stationarity result, consistent with Chen et al.
+(2025); Chang et al.
+(2025).
 Remark 3.4.
-Equation (5) clarifies the difference between the RC and R variants.
-The first term is the finite-step orthogonalization error.
-By Theorem 3.1, it is controlled by the singular spectrum of $S(\mathbf{M})$, so the two-sided RC map can provide a stronger spectral correction.
-The second term is the bias introduced by preprocessing.
-Proposition 3.4 shows that the pure RC update preserves descent only under the auxiliary positivity condition $\chi_{\varepsilon}>0$.
-By contrast, the row-normalized map admits the fixed alignment lower bound
-$\left\langle\mathbf{M},\operatorname{Orth}(\mathbf{P}(\mathbf{M})\mathbf{M})\right\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}\|_{F},\qquad\mathbf{P}(\mathbf{M}):=\operatorname{diag}\!\bigl(\operatorname{rowsum}(\mathbf{M}\odot\mathbf{M})\bigr)^{-1/2},$
-which is exactly why Theorem 3.5 no longer needs a condition of the form $\chi_{\varepsilon}>0$.
-Thus RC should be viewed as the stronger two-sided spectral corrector, whereas R is the cleaner one-sided variant and therefore the default choice for the hidden-weight setting considered in this paper.
-The column-normalized C variant is retained as the embedding-aligned companion and is studied empirically rather than through a separate main theorem.
-Remark 3.5.
-Theorem 3.5 yields the sublinear stationarity estimate
-$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}=\mathcal{O}\!\left(\frac{1+\ln T}{T^{1/4}}\right).$
-Hence, up to the logarithmic factor and the explicit problem-dependent constants, the default R variant has the same $T^{-1/4}$-type baseline dependence as the currently known convergence theory for standard stochastic nonconvex Muon-type methods (Chen et al., 2025; Chang et al., 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025).
-Improvements beyond this baseline typically require additional mechanisms such as variance reduction (Liu et al., 2025b; Chang et al., 2025; Huang et al., 2025; Qian et al., 2025; Fang et al., 2018; Zhou et al., 2020; Cutkosky and Orabona, 2019; Huang et al., 2021).
+Equation (5) separates finite-step orthogonalization error from preprocessing bias.
+RC is the stronger two-sided spectral corrector, but Proposition 3.4 requires the auxiliary condition $\chi_{\varepsilon}>0$.
+By contrast, the row-normalized map satisfies
+$\left\langle\mathbf{M},\operatorname{Orth}(\mathbf{P}(\mathbf{M})\mathbf{M})\right\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}\|_{F},\qquad\mathbf{P}(\mathbf{M}):=\bigl(\operatorname{diag}(\operatorname{rowsum}(\mathbf{M}\odot\mathbf{M}))^{1/2}\bigr)^{\dagger},$
+so Theorem 3.5 needs no $\chi_{\varepsilon}$-type condition, and Corollary 3.6 only degrades this coefficient by the explicit factor $1-\varepsilon_{\mathrm{ns}}$. Hence the default R variant obeys
+$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}=\mathcal{O}\left(\frac{1+\ln T}{T^{1/4}}\right),$
+matching the standard Muon-type baseline up to logarithmic (Chen et al., 2025; Chang et al., 2025; Gyu Yeol Kim, 2026; Shulgin et al., 2025) and constant factors, while C is retained as the embedding-aligned companion and studied empirically (Xu et al., 2026b).
+Improvements beyond this baseline typically require additional mechanisms such as variance reduction (Yuan et al., 2024; Chang et al., 2025; Liu et al., 2025b; Qian et al., 2025; Fang et al., 2018; Zhou et al., 2020; Cutkosky and Orabona, 2019; Huang et al., 2021).
 4 Experiments
-4.1 Main Result
-(a) Training Loss
-(b) Validation Loss
-(c) Wall-clock time
-Figure 3: The training and validation loss curves, plotted against both training tokens and wall-clock time on LLaMA-2 130M.
-(a) Training Loss
-(b) Validation Loss
-(c) Wall-clock time
-Figure 4: The training and validation loss curves, plotted against both training tokens and wall-clock time on LLaMA-2 350M.
+4.1 Main Results
 In this section, we evaluate the performance of the proposed MuonEq optimizers on LLM pretraining.
-All experiments are conducted using 64 Ascend 910C (64GB) NPUs.
-Detailed experimental settings are provided in Appendix J.
+We use the same 8-node Ascend 910C cluster for all LLaMA experiments; 130M uses 32 NPUs, whereas 350M and 1B use 64 NPUs.
+Detailed experimental settings are provided in Appendix K.
+We will refer to the MuonEq with Nesterov acceleration enabled as MuonEq-Nes.
+Refer to captionRefer to captionRefer to caption
+((a))
+Refer to captionRefer to captionRefer to caption
+((b))
+Refer to captionRefer to captionRefer to caption
+((c))
+Figure 3: Comparison of LLaMA2 130M, 350M, and 1B from three perspectives.
+In each column, the top panel corresponds to the 130M model, the middle panel corresponds to the 350M model, and the bottom panel corresponds to the 1B model.
+From left to right, the panels show training loss versus tokens, validation loss versus tokens, and validation loss versus wall-clock time.
 […]
-Figure 3 reports training loss versus tokens, validation loss versus tokens, and validation loss versus wall-clock time for LLaMA2-130M trained to 10.5B tokens.
-Figure 4 reports the corresponding validation-loss comparison for LLaMA2-350M trained to 21.0B tokens.
-Table 1 reports the best validation perplexity, peak memory (GB) usage, and training wall-clock time (Second/Step) for these main comparison runs.
-Across Figure 3, Figure 4, and Table 1, MuonEq consistently improves over Muon.
 Table 1: Best validation perplexity, peak memory usage, and wall-clock training time on LLaMA2/C4.
-Optimizer 130M 350M
-Perplexity Memory Time Cost Perplexity Memory Time Cost
-AdamW Loshchilov and Hutter (2019) 21.35 12.94 0.21 16.25 29.43 0.53
-MARS-AdamW Liu et al. (2025b) 20.03 12.99 0.23 15.62 29.47 0.57
-Muon Jordan et al. ; Liu et al. (2025a) 20.52 12.52 0.25 15.21 28.18 0.61
-Muon-Nes Jordan et al. ; Liu et al. (2025a); Chang et al. (2025) 20.18 12.52 0.25 15.19 28.18 0.61
+Optimizer 130M 350M 1B
+Perplexity Memory Time Perplexity Memory Time Perplexity Memory Time
+AdamW Loshchilov and Hutter (2019) 21.35 12.94 0.21 16.25 29.43 0.53 14.71 39.56 2.47
+MARS-AdamW Yuan et al. (2024) 20.03 12.99 0.23 15.62 29.47 0.57 13.22 39.57 2.50
+Muon Jordan et al. ; Liu et al. (2025a) 20.52 12.52 0.25 15.21 28.18 0.61 13.06 34.81 2.76
+Muon-Nes Jordan et al. ; Liu et al. (2025a); Chang et al. (2025) 20.18 12.52 0.25 15.19 28.18 0.61 13.24 34.81 2.76
 […]
-Training Tokens 10.5B 21.0B
+[…]
+Training Tokens 10.5B 21.0B 21.0B
 4.2 Ablation Study
-We ablate the three MuonEq variants RC, R, and C to isolate pre-orthogonalization geometry.
-All variants use the same optimizer states, scalar hyperparameters, and Newton–Schulz budget; only the pre-orthogonalization diagonal map changes.
-We take R as the default variant.
-All ablations run on 4 RTX Pro6000 (96GB) GPUs; details are deferred to Appendix J.2.
-The results consistently favor R.
-Among the three static variants, R performs best on both CIFAR-10/ResNet-18 and FineWeb/GPT2-small, reaching 94.53% test accuracy on CIFAR-10 and 24.94 validation perplexity on FineWeb/GPT2-small.
-RC is competitive but falls short of R, while C is weaker in this hidden-weight setting.
-Full static-variant comparisons appear in Tables 3 and 3.
-This pattern matches the trade-off in Eq.
-(5) and the analysis in Section 3.
-RC provides stronger two-sided spectral correction and can reduce the finite-step Newton–Schulz term more aggressively, but also introduces larger preconditioning bias.
-For the hidden matrix weights targeted by MuonEq, the more relevant one-sided geometry is row-sided, so R offers the best trade-off between spectral correction and geometric alignment.
-This interpretation is consistent with (Xu et al., 2026b).
-Comparisons with (Pethick et al., 2025) and (Glentis et al., 2025) require care because their row/column terminology depends on parameter layout and task setting; we discuss this point in the appendix J.3.
+We ablate the three static MuonEq-Nes variants, RC, R, and C, to isolate the effect of pre-orthogonalization geometry.
+All ablations run on 4 RTX Pro6000 (96GB) GPUs.
+Detailed experimental settings are provided in Appendix K.2.
+Tables 3 and 3 show that, with results averaged over three random seeds, R is consistently the strongest static variant across both benchmarks, achieving higher test accuracy than Muon-Nes on CIFAR-10 with ResNet-18 and lower validation perplexity on FineWeb-10.5B tokens with GPT2-small.
+((a))
+((b))
+((c))
+((d))
+Figure 4: (a,b) Validation perplexity and step time vs.
+Newton–Schulz iterations $K$ on GPT2-small/FineWeb (10.5B tokens); (c,d) validation perplexity heatmaps over learning rate and momentum at $K=5$ (2.6B tokens).
 Table 2: Ablation on CIFAR-10 with ResNet-18.
@@
 Baseline Experiments
-SGD 92.84
-AdamW 93.42
-Muon-Nes 94.41
+SGD 92.82±0.04
+AdamW 93.36±0.10
+Muon 94.39±0.04
+Muon-Nes 94.44±0.05
 Ablation Experiments
-MuonEq (RC) 94.23
-MuonEq (C) 94.44
+MuonEq-Nes (RC) 94.29±0.10
+MuonEq-Nes (C) 94.45±0.02
 […]
-Table 3: Ablation on Fineweb-10.5B tokens with GPT2-small.
+Table 3: Ablation on FineWeb with GPT2-small.
 Method Component Validation Perplexity
@@
 Baseline Experiments
-AdamW 26.60
-Muon-Nes 25.31
+AdamW 26.48±0.12
+Muon-Nes 25.23±0.08
+AdaMuon 24.99±0.13
+Muon+ 25.12±0.08
+Mousse 25.16±0.10
 Ablation Experiments
-MuonEq (RC) 24.98
-MuonEq (C) 25.23
+MuonEq-Nes (RC) 24.94±0.04
+MuonEq-Nes (C) 25.23±0.06
 […]
-5 Related Work
+Figure LABEL:fig:muon-val and LABEL:fig:muon-speed show that the default R variant of MuonEq-Nes is markedly less sensitive to the NS iteration count $K$ than Muon-Nes: its validation perplexity remains stable over a wider range of $K$, while step time stays comparable throughout the sweep.
+Complementary robustness evidence is given in Figure LABEL:fig:heatmap-muon-k5 and LABEL:fig:heatmap-muoneq-k5, where MuonEq-Nes (R) exhibits a broader low-perplexity region over learning-rate and momentum sweeps than Muon-Nes, especially at $K=5$.
+This pattern is consistent with Eq.
+(5) and Section 3.
+RC applies stronger two-sided spectral correction and can reduce finite-step NS error more aggressively, but at the cost of larger preconditioning bias.
+For the hidden-weight settings we tested, R provides a favorable trade-off between spectral correction and preprocessing bias, consistent with (Xu et al., 2026b).
+Comparisons with (Pethick et al., 2025) and (Glentis et al., 2025) require caution, as row/column terms vary by layout and task; see Appendix K.3.
+5 Conclusion
+We improve the matrix geometry encountered by finite-step orthogonalization in Muon-style optimization without resorting to full whitening.
+MuonEq reshapes the input to Newton–Schulz using lightweight pre-orthogonalization equilibration in three forms, RC, R, and C; R is used by default for hidden weights.
+Theoretically, these schemes serve as zeroth-order whitening surrogates: RC provides stronger two-sided correction, whereas R yields cleaner one-sided geometry and supports the main $\widetilde{\mathcal{O}}(T^{-1/4})$ guarantee.
+Empirically, R consistently outperforms Muon in LLaMA2 pretraining on C4 across all tested scales and budgets.
+Appendix
+Appendix A Related Work
 MuonEq is viewed as a lightweight pre-balanced Muon family rather than as a fully whitened optimizer.
 It inserts diagonal equilibration before Newton–Schulz orthogonalization and includes three forms: two-sided row/column normalization (RC), row normalization (R), and column normalization (C), with R as the default choice for the hidden matrix weights considered in this paper.
-All three use only row/column squared-norm statistics and preserve the $\mathcal{O}(m+n)$ state profile.
+All three compute row/column squared norms from the current momentum and add no persistent optimizer state beyond Muon.
 Muon and orthogonalized-update optimizers.
-Muon Jordan et al.
-introduced the core design of orthogonalizing matrix-valued momentum with a small fixed number of Newton–Schulz iterations.
-Subsequent work has expanded this idea into a broader family of orthogonalized-update optimizers, including post-orthogonalization normalization and rescaling methods such as Muon+, NorMuon, and AdaMuon Zhang et al.
-(2026b); Li et al.
-(2025); Si et al.
-(2025); adaptive stepsize variants such as AdaGO and NAMO Zhang et al.
-(2025, 2026a); distributed and communication-aware extensions such as Dion and MuLoCo Ahn et al.
-(2025); Thérien et al.
-(2025); reduced-state implementations Gupta et al.
-(2025); and fine-tuning-oriented variants Page et al.
-(2025).
+Muon [22] introduced the core design of orthogonalizing matrix-valued momentum with a small fixed number of Newton–Schulz iterations.
+Subsequent work has expanded this idea into a broader family of orthogonalized-update optimizers, including post-orthogonalization normalization and rescaling methods such as Muon+, NorMuon, and AdaMuon [59, 28, 48]; adaptive stepsize variants such as AdaGO and NAMO [57, 58]; distributed and communication-aware extensions such as Dion and MuLoCo [1, 49]; reduced-state implementations [16]; and fine-tuning-oriented variants [34].
 These methods mainly alter the orthogonalized update itself or the surrounding optimization machinery.
@@
 Theory of Muon-style methods.
-Recent theory interprets Muon as steepest descent under spectral-norm or nuclear-norm geometry Bernstein and Newhouse (2024); Chen et al.
-(2025), establishes stochastic and nonconvex convergence guarantees Shen et al.
-(2025); Chang et al.
-(2025); Nagashima and Iiduka (2026); Zhang and Lin (2026), and analyzes practical ingredients such as weight decay, critical batch size, and inexact orthogonalization Sato et al.
-(2025); Gyu Yeol Kim (2026); Shulgin et al.
-(2025).
+Recent theory interprets Muon as steepest descent under spectral-norm or nuclear-norm geometry [4, 8], establishes stochastic and nonconvex convergence guarantees [46, 6, 33, 61], and analyzes practical ingredients such as weight decay, critical batch size, and inexact orthogonalization [42, 18, 47].
 This literature is directly relevant to MuonEq because we retain the same finite-step orthogonalization primitive as Muon and change only the geometry of its input.
@@
 Structured preconditioning, factorized statistics, and whitening-inspired methods.
-A separate line of work studies matrix-aware preconditioning, including Shampoo Gupta et al.
-(2018), K-FAC Martens and Grosse (2015); Eschenhagen et al.
-(2023), SOAP Vyas et al.
-(2024), Sketchy Feinberg et al.
-(2023), and ASGO An et al.
-(2025).
-Adafactor Shazeer and Stern (2018) is especially relevant from a systems perspective because it shows how row- and column-factorized second-moment statistics reduce optimizer state to $\mathcal{O}(m+n)$ for matrix-valued parameters.
-Recent Muon-adjacent whitening or preconditioned-polar methods, including FISMO and Mousse Xu et al.
-(2026a); Zhang et al.
-(2026c), pursue a similar high-level goal: improve the geometry before imposing spectral constraints.
+A separate line of work studies matrix-aware preconditioning, including Shampoo [17], K-FAC [32, 11], SOAP [51], Sketchy [13], and ASGO [3].
+Adafactor [45] is especially relevant from a systems perspective because it shows how row- and column-factorized second-moment statistics reduce optimizer state to $\mathcal{O}(m+n)$ for matrix-valued parameters.
+Recent Muon-adjacent whitening or preconditioned-polar methods, including FISMO and Mousse [54, 60], pursue a similar high-level goal: improve the geometry before imposing spectral constraints.
 The key difference is that these methods rely on richer matrix- or Kronecker-structured preconditioners, whereas MuonEq restricts the intervention to diagonal row/column equilibration.
@@
 Orthogonalization algorithms.
-There is also a complementary numerical literature on faster and more accurate polar/orthogonalization routines, including improved Newton–Schulz variants and matrix-sign methods such as CANS and Polar Express Grishina et al.
-(2025); Amsel et al.
-(2025).
+There is also a complementary numerical literature on faster and more accurate polar/orthogonalization routines, including improved Newton–Schulz variants and matrix-sign methods such as CANS and Polar Express [15, 2].
 This literature matters because Muon-style optimizers operate in the regime of approximate orthogonalization, where only a small number of iterations is affordable; MuonEq is designed explicitly for this setting.
-6 Conclusion
-We study how to improve the matrix geometry seen by finite-step orthogonalization in Muon-style optimization without the cost of full whitening.
-MuonEq does so with a lightweight family of pre-orthogonalization equilibration schemes: two-sided row/column normalization (RC), row normalization (R), and column normalization (C).
-These variants modify the input to Newton–Schulz orthogonalization while preserving the $\mathcal{O}(m+n)$ memory profile of factorized methods.
-For the hidden matrix weights studied here, we use R as the default variant.
-The analysis shows that finite-step orthogonalization is highly sensitive to the input spectrum and that row/column normalization acts as a zeroth-order whitening surrogate by removing marginal scale mismatch.
-It also clarifies the trade-off within the MuonEq family: RC gives stronger two-sided spectral correction, whereas R yields cleaner one-sided geometry and the main $\widetilde{\mathcal{O}}(T^{-1/4})$ nonconvex guarantee.
-Empirically, the default R variant consistently outperforms Muon in LLaMA2 pretraining on C4 across evaluated model sizes and token budgets.
-Ablations over RC, R, and C further support R as the strongest default for the hidden-weight setting considered here.
-Appendix
-Appendix A Proofs of Theorem 3.1
+Appendix B Proofs of Theorem 3.1
 Proof.
-The claim for $k=0$ is immediate from
+Let $\alpha>0$. For $k=0$, the desired form follows directly from
 $\mathbf{X}_{0}=\alpha^{-1}\mathbf{G}=\mathbf{U}\operatorname{diag}\Bigl(\frac{\sigma_{1}}{\alpha},\dots,\frac{\sigma_{r}}{\alpha}\Bigr)\mathbf{V}^{\top}.$
@@
 $\mathbf{X}_{k}\mathbf{X}_{k}^{\top}=\mathbf{U}\operatorname{diag}\bigl((s_{1}^{(k)})^{2},\dots,(s_{r}^{(k)})^{2}\bigr)\mathbf{U}^{\top}.$
-Substituting this into the iteration and using $\mathbf{U}^{\top}\mathbf{U}=I_{r}$, we obtain
+Inserting this expression into the iteration and using $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{r}$ gives
 $\mathbf{X}_{k+1}=\mathbf{U}\operatorname{diag}\bigl(\phi(s_{1}^{(k)}),\dots,\phi(s_{r}^{(k)})\bigr)\mathbf{V}^{\top}.$
-Hence, by induction,
+By induction,
 $\mathbf{X}_{k}=\mathbf{U}\operatorname{diag}\bigl(s_{1}^{(k)},\dots,s_{r}^{(k)}\bigr)\mathbf{V}^{\top},\qquad s_{i}^{(k+1)}=\phi\left(s_{i}^{(k)}\right).$
-Moreover, we have
-$\operatorname{Orth}(\mathbf{G})=\mathbf{U}\mathbf{V}^{\top}=\mathbf{U}I_{r}\mathbf{V}^{\top},\quad\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})=\mathbf{U}\Bigl(\operatorname{diag}(s_{1}^{(k)},\dots,s_{r}^{(k)})-I_{r}\Bigr)\mathbf{V}^{\top}.$
-Since left and right multiplication by matrices with orthonormal columns preserves the Frobenius norm, it follows that
-$\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}^{2}=\sum_{i=1}^{r}(s_{i}^{(k)}-1)^{2},$
-which proves the exact error representation.
-We next prove the lower bound. Define
-$q(t):=a+bt+ct^{2},\qquad t\in[0,1].$
-Then
-$q^{\prime}(t)=b+2ct=-4.7750+4.0630t<0,\qquad t\in[0,1].$
-Hence $q$ is strictly decreasing on $[0,1]$, and in particular
-$0<q(1)=a+b+c=0.701\leq q(t)\leq q(0)=a,\qquad t\in[0,1].$
-Fix any $i\in\{1,\dots,r\}$ and $k\geq 0$. If $a^{k}\sigma_{i}/\alpha\geq 1$, then trivially
-$|1-s_{i}^{(k)}|\geq 0=\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}.$
+We also have
+$\operatorname{Orth}(\mathbf{G})=\mathbf{U}\mathbf{V}^{\top}=\mathbf{U}\mathbf{I}_{r}\mathbf{V}^{\top},\quad\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})=\mathbf{U}\Bigl(\operatorname{diag}(s_{1}^{(k)},\dots,s_{r}^{(k)})-\mathbf{I}_{r}\Bigr)\mathbf{V}^{\top}.$
+Left and right multiplication by matrices with orthonormal columns preserves the Frobenius norm, so
+$\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}^{2}=\sum_{i=1}^{r}(s_{i}^{(k)}-1)^{2}.$
+This proves the exact error representation.
+We now establish the lower bound. Recall that
+$q(t):=a+bt+ct^{2}.$
+By assumption,
+$0<q(t)\leq a,\qquad t\in[0,1].$
+Fix $i\in[r]$ and $k\geq 0$. If $a^{k}\frac{\sigma_{i}}{\alpha}\geq 1,$ then
+$\bigl|1-s_{i}^{(k)}\bigr|\geq 0=\left(1-a^{k}\frac{\sigma_{i}}{\alpha}\right)_{+}.$
 Now suppose that $a^{k}\sigma_{i}/\alpha<1$. We claim that, for every $0\leq j\leq k$,
 $0<s_{i}^{(j)}\leq a^{j}\frac{\sigma_{i}}{\alpha}<1.$
-The case $j=0$ is immediate. Suppose the claim holds for some $j<k$. Then $s_{i}^{(j)}\in(0,1)$, so $(s_{i}^{(j)})^{2}\in(0,1)$ and thus
+The statement is immediate for $j=0$. Suppose it holds for some $j<k$. Then $s_{i}^{(j)}\in(0,1)$, and hence $(s_{i}^{(j)})^{2}\in(0,1)$. Therefore,
 $0<q\bigl((s_{i}^{(j)})^{2}\bigr)\leq a.$
-Therefore,
+It follows that
 $0<s_{i}^{(j+1)}=s_{i}^{(j)}q\bigl((s_{i}^{(j)})^{2}\bigr)\leq as_{i}^{(j)}\leq a^{j+1}\frac{\sigma_{i}}{\alpha}.$
-Since $j+1\leq k$, we further have
+Because $j+1\leq k$,
 $a^{j+1}\frac{\sigma_{i}}{\alpha}\leq a^{k}\frac{\sigma_{i}}{\alpha}<1.$
@@
 $0<s_{i}^{(k)}\leq a^{k}\frac{\sigma_{i}}{\alpha}<1,$
-and hence
+and thus
 $|1-s_{i}^{(k)}|=1-s_{i}^{(k)}\geq 1-a^{k}\frac{\sigma_{i}}{\alpha}=\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}.$
-Combining the two cases, we conclude that, for every $i$,
+Combining the two cases gives, for every $i$,
 $|1-s_{i}^{(k)}|\geq\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}.$
-Squaring both sides, summing over $i$, and invoking the exact error representation established above yields
-$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\Bigl(1-a^{k}\frac{\sigma_{i}}{\alpha}\Bigr)_{+}^{2}\right)^{1/2}.$
+Finally, set $\alpha=\|\mathbf{G}\|_{F}$. Squaring both sides, summing over $i$, and using the exact error representation together with
+$\frac{\sigma_{i}}{\|\mathbf{G}\|_{F}}=\frac{\sigma_{i}/\sigma_{1}}{\|\mathbf{G}\|_{F}/\sigma_{1}}=\frac{1}{\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}},$
+we obtain
+$\frac{1}{\sqrt{r}}\bigl\|\mathbf{X}_{k}-\operatorname{Orth}(\mathbf{G})\bigr\|_{F}\geq\frac{1}{\sqrt{r}}\left(\sum_{i=1}^{r}\left(1-\frac{a^{k}}{\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}}\right)_{+}^{2}\right)^{1/2}.$
+Moreover,
+$\tau_{i}:=\log_{a}\frac{\|\mathbf{G}\|_{F}}{\sigma_{i}}=\log_{a}\left(\kappa_{i}\sqrt{\operatorname{sr}(\mathbf{G})}\right),$
+so
+$\tau_{1}=\log_{a}\sqrt{\operatorname{sr}(\mathbf{G})},\qquad\tau_{r}=\log_{a}\left(\kappa(\mathbf{G})\sqrt{\operatorname{sr}(\mathbf{G})}\right),\qquad\tau_{r}-\tau_{1}=\log_{a}\kappa(\mathbf{G}).$
 This completes the proof. ∎
-Remark A.1.
-Since
-$\phi(1)=a+b+c=0.701\neq 1,$
-the scalar map induced by the above Muon coefficients does not have $1$ as a fixed point.
-Therefore, $T_{\varepsilon}$ in the theorem should be interpreted as the first hitting time of the error threshold $\varepsilon$, rather than as the number of iterations required for asymptotic convergence to $\operatorname{Orth}(\mathbf{G})$.
-Appendix B Lemmas for Proposition 3.2
-B.1 Lemma B.1
-Lemma B.1.
+Appendix C Lemmas for Proposition 3.2
+C.1 Lemma C.1
+Lemma C.1.
 Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be symmetric positive definite, and let $\mathbf{E}\in\mathbb{R}^{n\times n}$ be symmetric.
@@
 Then $g$ and $f$ are Fréchet differentiable at $\mathbf{A}$.
+We use $L_{h}(\mathbf{A};\mathbf{E})$ to denote the Fréchet derivative of a matrix function $h$ at $\mathbf{A}$ applied to $\mathbf{E}$, i.e., $h(\mathbf{A}+\mathbf{E})=h(\mathbf{A})+L_{h}(\mathbf{A};\mathbf{E})+o(\|\mathbf{E}\|_{2})$.
 If $\mathbf{S}=\mathbf{A}^{1/2},\mathbf{L}=L_{g}(\mathbf{A};\mathbf{E})$, then $\mathbf{L}$ is the unique symmetric solution of the Sylvester equation
@@
 The stated first-order expansion follows from the Fréchet differentiability of $f$ at $\mathbf{A}$. ∎
-B.2 Lemma B.2
-Lemma B.2.
+C.2 Lemma C.2
+Lemma C.2.
 Let $\mathbf{A}\succ 0$ and set $\mathbf{S}=\mathbf{A}^{1/2}\succ 0$. Consider the Sylvester equation
@@
 Since $\lambda_{\min}(\mathbf{S})=\sqrt{\lambda_{\min}(\mathbf{A})}$, the stated bound follows. ∎
-Appendix C Proofs of Proposition 3.2
+Appendix D Proofs of Proposition 3.2
 Proof.
@@
 $\mathbf{M}^{\top}\mathbf{M}=\mathbf{D}_{c}(\mathbf{I}_{n}+\mathbf{C}_{c})\mathbf{D}_{c}=\mathbf{D}_{c}^{2}+\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}.$
-Apply Lemma B.1 with $\mathbf{A}=\mathbf{D}_{c}^{2},\mathbf{E}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}$. Then
+Apply Lemma C.1 with $\mathbf{A}=\mathbf{D}_{c}^{2},\mathbf{E}=\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}$. Then
 $(\mathbf{M}^{\top}\mathbf{M})^{-1/2}=\mathbf{D}_{c}^{-1}-\mathbf{D}_{c}^{-1}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}),$
@@
 $=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}).$
-Moreover, by Lemma B.2, we have
+Moreover, by Lemma C.2, we have
 $\|\mathbf{L}_{c}\|_{2}\leq\frac{1}{2\lambda_{\min}(\mathbf{D}_{c})}\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}\leq\frac{\|\mathbf{D}_{c}\|_{2}^{2}}{2\lambda_{\min}(\mathbf{D}_{c})}\|\mathbf{C}_{c}\|_{2}.$
@@
 $\mathbf{M}\mathbf{M}^{\top}=\mathbf{D}_{r}(\mathbf{I}_{m}+\mathbf{C}_{r})\mathbf{D}_{r}=\mathbf{D}_{r}^{2}+\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}.$
-Applying Lemma B.1 with $\mathbf{A}=\mathbf{D}_{r}^{2},\mathbf{E}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}$ gives
+Applying Lemma C.1 with $\mathbf{A}=\mathbf{D}_{r}^{2},\mathbf{E}=\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}$ gives
 $(\mathbf{M}\mathbf{M}^{\top})^{-1/2}=\mathbf{D}_{r}^{-1}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{D}_{r}^{-1}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}),$
@@
 $=\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}).$
-Again by Lemma B.2, we have
+Again by Lemma C.2, we have
 $\|\mathbf{L}_{r}\|_{2}\leq\frac{1}{2\lambda_{\min}(\mathbf{D}_{r})}\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}\leq\frac{\|\mathbf{D}_{r}\|_{2}^{2}}{2\lambda_{\min}(\mathbf{D}_{r})}\|\mathbf{C}_{r}\|_{2},$
@@
 $\operatorname{Orth}_{\ell}(\mathbf{M})=\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}).$
-Appendix D Normalization and whitening
-This appendix records the normalization–whitening estimates omitted from the main text.
+Appendix E Normalization and whitening
+This appendix collects the normalization–whitening estimates that are omitted from the main text.
 We use the notation of Proposition 3.2.
-All $\mathcal{O}(\cdot)$ terms are understood in operator norm; the implied constants may depend on the fixed matrices in Proposition 3.2, but not on the perturbation terms.
-Proposition D.1.
-The following perturbative statements hold.
-  1. (a)
-    In the column/right setting of Proposition 3.2, if $\|\mathbf{C}_{c}\|_{2}$ is small enough, then $\operatorname{Orth}_{r}(\mathbf{N}_{c})=\mathbf{N}_{c}(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{N}_{c}-\tfrac{1}{2}\mathbf{N}_{c}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$, so $\|\operatorname{Orth}_{r}(\mathbf{N}_{c})-\mathbf{N}_{c}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{c}\|_{2}\|\mathbf{C}_{c}\|_{2}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$.
-Also,
+All $\mathcal{O}(\cdot)$ terms are taken in operator norm; the implied constants may depend on the fixed matrices in Proposition 3.2, but not on the perturbation terms.
+Proposition E.1.
+The following estimates hold.
+In the column/right setting of Proposition 3.2, if $\|\mathbf{C}_{c}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{r}(\mathbf{N}_{c})=\mathbf{N}_{c}(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{N}_{c}-\tfrac{1}{2}\mathbf{N}_{c}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{r}(\mathbf{N}_{c})-\mathbf{N}_{c}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{c}\|_{2}\|\mathbf{C}_{c}\|_{2}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$.
+Moreover,
 $\|\operatorname{Orth}_{r}(\mathbf{M})-\mathbf{M}\|_{2}\leq{}$ $\|\mathbf{N}_{c}\|_{2}\|\mathbf{D}_{c}-\mathbf{I}_{n}\|_{2}$
 $+\frac{\kappa(\mathbf{D}_{c})^{2}}{2}\|\mathbf{N}_{c}\|_{2}\|\mathbf{C}_{c}\|_{2}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}).$
-  2. (b)
-    In the row/left setting of Proposition 3.2, if $\|\mathbf{C}_{r}\|_{2}$ is small enough, then $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, so $\|\operatorname{Orth}_{\ell}(\mathbf{N}_{r})-\mathbf{N}_{r}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{r}\|_{2}\|\mathbf{C}_{r}\|_{2}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$.
-Also,
+In the row/left setting of Proposition 3.2, if $\|\mathbf{C}_{r}\|_{2}$ is sufficiently small, then $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{\ell}(\mathbf{N}_{r})-\mathbf{N}_{r}\|_{2}\leq\tfrac{1}{2}\|\mathbf{N}_{r}\|_{2}\|\mathbf{C}_{r}\|_{2}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$.
+Moreover,
 $\|\operatorname{Orth}_{\ell}(\mathbf{M})-\mathbf{M}\|_{2}\leq{}$ $\|\mathbf{D}_{r}-\mathbf{I}_{m}\|_{2}\|\mathbf{N}_{r}\|_{2}$
@@
 Proof.
-For (a), Lemma B.1 with $\mathbf{A}=\mathbf{I}_{n}$ and $\mathbf{E}=\mathbf{C}_{c}$ gives the Sylvester equation $\mathbf{L}+\mathbf{L}=\mathbf{C}_{c}$, hence $\mathbf{L}=\tfrac{1}{2}\mathbf{C}_{c}$ and $(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{I}_{n}-\tfrac{1}{2}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$.
-Left multiplication by $\mathbf{N}_{c}$ gives the expansion of $\operatorname{Orth}_{r}(\mathbf{N}_{c})$, and the norm bound follows from submultiplicativity.
-For the comparison with $\mathbf{M}$, Proposition 3.2 gives $\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2})$.
+Consider first the column/right case.
+Applying Lemma C.1 with $\mathbf{A}=\mathbf{I}_{n}$ and $\mathbf{E}=\mathbf{C}_{c}$ gives the Sylvester equation $\mathbf{L}+\mathbf{L}=\mathbf{C}_{c}$.
+Thus $\mathbf{L}=\tfrac{1}{2}\mathbf{C}_{c}$, and $(\mathbf{I}_{n}+\mathbf{C}_{c})^{-1/2}=\mathbf{I}_{n}-\tfrac{1}{2}\mathbf{C}_{c}+\mathcal{O}(\|\mathbf{C}_{c}\|_{2}^{2})$.
+Multiplying on the left by $\mathbf{N}_{c}$ gives the expansion of $\operatorname{Orth}_{r}(\mathbf{N}_{c})$, and the norm bound follows from submultiplicativity.
+To compare with $\mathbf{M}$, Proposition 3.2 gives $\operatorname{Orth}_{r}(\mathbf{M})=\mathbf{N}_{c}-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2})$.
 Since $\mathbf{M}=\mathbf{N}_{c}\mathbf{D}_{c}$,
 $\operatorname{Orth}_{r}(\mathbf{M})-\mathbf{M}=\mathbf{N}_{c}(\mathbf{I}_{n}-\mathbf{D}_{c})-\mathbf{N}_{c}\mathbf{L}_{c}\mathbf{D}_{c}^{-1}+\mathcal{O}(\|\mathbf{D}_{c}\mathbf{C}_{c}\mathbf{D}_{c}\|_{2}^{2}).$
-Lemma B.2 yields $\|\mathbf{L}_{c}\mathbf{D}_{c}^{-1}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{c})^{2}\|\mathbf{C}_{c}\|_{2}$, which gives the estimate.
-The proof of (b) is the same.
-Lemma B.1 with $\mathbf{A}=\mathbf{I}_{m}$ and $\mathbf{E}=\mathbf{C}_{r}$ gives $(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}=\mathbf{I}_{m}-\tfrac{1}{2}\mathbf{C}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, hence $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$.
+Lemma C.2 yields $\|\mathbf{L}_{c}\mathbf{D}_{c}^{-1}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{c})^{2}\|\mathbf{C}_{c}\|_{2}$, which gives the stated estimate.
+The row/left case follows by the same argument.
+Applying Lemma C.1 with $\mathbf{A}=\mathbf{I}_{m}$ and $\mathbf{E}=\mathbf{C}_{r}$ gives $(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}=\mathbf{I}_{m}-\tfrac{1}{2}\mathbf{C}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$, and hence $\operatorname{Orth}_{\ell}(\mathbf{N}_{r})=(\mathbf{I}_{m}+\mathbf{C}_{r})^{-1/2}\mathbf{N}_{r}=\mathbf{N}_{r}-\tfrac{1}{2}\mathbf{C}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{C}_{r}\|_{2}^{2})$.
 The norm bound again follows from submultiplicativity.
@@
 $\operatorname{Orth}_{\ell}(\mathbf{M})-\mathbf{M}=(\mathbf{I}_{m}-\mathbf{D}_{r})\mathbf{N}_{r}-\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\mathbf{N}_{r}+\mathcal{O}(\|\mathbf{D}_{r}\mathbf{C}_{r}\mathbf{D}_{r}\|_{2}^{2}).$
-Using $\|\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{r})^{2}\|\mathbf{C}_{r}\|_{2}$ completes the proof. ∎
-Remark D.1.
-Proposition D.1 shows that one-sided normalization removes the zeroth-order marginal scale terms $\|\mathbf{N}_{c}\|_{2}\|\mathbf{D}_{c}-\mathbf{I}_{n}\|_{2}$ and $\|\mathbf{D}_{r}-\mathbf{I}_{m}\|_{2}\|\mathbf{N}_{r}\|_{2}$.
+The bound $\|\mathbf{D}_{r}^{-1}\mathbf{L}_{r}\|_{2}\leq\tfrac{1}{2}\kappa(\mathbf{D}_{r})^{2}\|\mathbf{C}_{r}\|_{2}$ completes the proof. ∎
+Remark E.1.
+Proposition E.1 shows that one-sided normalization removes the zeroth-order marginal scale terms $\|\mathbf{N}_{c}\|_{2}\|\mathbf{D}_{c}-\mathbf{I}_{n}\|_{2}$ and $\|\mathbf{D}_{r}-\mathbf{I}_{m}\|_{2}\|\mathbf{N}_{r}\|_{2}$.
 Whitening then acts only on the first-order Gram error.
-Since $\operatorname{diag}(\mathbf{N}_{c}^{\top}\mathbf{N}_{c})=\mathbf{I}_{n}$ and $\operatorname{diag}(\mathbf{N}_{r}\mathbf{N}_{r}^{\top})=\mathbf{I}_{m}$, this leading correction is purely off-diagonal.
-Corollary D.2.
+Since $\operatorname{diag}(\mathbf{N}_{c}^{\top}\mathbf{N}_{c})=\mathbf{I}_{n}$ and $\operatorname{diag}(\mathbf{N}_{r}\mathbf{N}_{r}^{\top})=\mathbf{I}_{m}$, the leading correction is purely off-diagonal.
+Corollary E.2.
 Assume that every row and every column of $\mathbf{M}$ is nonzero, and set $\hat{\mathbf{M}}:=\mathbf{D}_{r}^{-1/2}\mathbf{M}\mathbf{D}_{c}^{-1/2}$.
-  1. (a)
-    If $\mathbf{M}$ has full column rank and $\hat{\mathbf{C}}_{c}:=\hat{\mathbf{M}}^{\top}\hat{\mathbf{M}}-\mathbf{I}_{n}$, then for $\|\hat{\mathbf{C}}_{c}\|_{2}$ small enough, $\operatorname{Orth}_{r}(\hat{\mathbf{M}})=\hat{\mathbf{M}}(\mathbf{I}_{n}+\hat{\mathbf{C}}_{c})^{-1/2}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{M}}\hat{\mathbf{C}}_{c}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$, so $\|\operatorname{Orth}_{r}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{c}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$.
-  2. (b)
-    If $\mathbf{M}$ has full row rank and $\hat{\mathbf{C}}_{r}:=\hat{\mathbf{M}}\hat{\mathbf{M}}^{\top}-\mathbf{I}_{m}$, then for $\|\hat{\mathbf{C}}_{r}\|_{2}$ small enough, $\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})=(\mathbf{I}_{m}+\hat{\mathbf{C}}_{r})^{-1/2}\hat{\mathbf{M}}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{C}}_{r}\hat{\mathbf{M}}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$, so $\|\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{r}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$.
+If $\mathbf{M}$ has full column rank and $\hat{\mathbf{C}}_{c}:=\hat{\mathbf{M}}^{\top}\hat{\mathbf{M}}-\mathbf{I}_{n}$, then for $\|\hat{\mathbf{C}}_{c}\|_{2}$ sufficiently small, $\operatorname{Orth}_{r}(\hat{\mathbf{M}})=\hat{\mathbf{M}}(\mathbf{I}_{n}+\hat{\mathbf{C}}_{c})^{-1/2}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{M}}\hat{\mathbf{C}}_{c}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{r}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{c}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{c}\|_{2}^{2})$.
+If $\mathbf{M}$ has full row rank and $\hat{\mathbf{C}}_{r}:=\hat{\mathbf{M}}\hat{\mathbf{M}}^{\top}-\mathbf{I}_{m}$, then for $\|\hat{\mathbf{C}}_{r}\|_{2}$ sufficiently small, $\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})=(\mathbf{I}_{m}+\hat{\mathbf{C}}_{r})^{-1/2}\hat{\mathbf{M}}=\hat{\mathbf{M}}-\tfrac{1}{2}\hat{\mathbf{C}}_{r}\hat{\mathbf{M}}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$, and therefore $\|\operatorname{Orth}_{\ell}(\hat{\mathbf{M}})-\hat{\mathbf{M}}\|_{2}\leq\tfrac{1}{2}\|\hat{\mathbf{M}}\|_{2}\|\hat{\mathbf{C}}_{r}\|_{2}+\mathcal{O}(\|\hat{\mathbf{C}}_{r}\|_{2}^{2})$.
 Proof.
@@
 […]
-Remark D.2.
+Remark E.2.
 After two-sided normalization, whitening is centered at the identity Gram matrix of $\hat{\mathbf{M}}$.
 It no longer corrects marginal row or column scales; it only removes the residual mismatch left after diagonal equilibration.
-Appendix E Lemmas for Proposition 3.4
-E.1 Lemma E.1
-Lemma E.1.
-Let the momentum iterate satisfy $\mathbf{M}_{t}=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{G}_{t},\qquad\beta\in[0,1)$, and assume there exists a constant $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty},\|\mathbf{M}_{0}\|_{\infty}\leq G_{\infty}$.
+Appendix F Lemmas for Proposition 3.4
+F.1 Lemma F.1
+Lemma F.1.
+Let the momentum iterate satisfy $\mathbf{M}_{t}=\beta\mathbf{M}_{t-1}+(1-\beta)\mathbf{G}_{t},\beta\in[0,1)$, and assume there exists a constant $G_{\infty}>0$ such that $\|\mathbf{G}_{t}\|_{\infty}\leq G_{\infty},\|\mathbf{M}_{0}\|_{\infty}\leq G_{\infty}$.
 For $\varepsilon>0$, define $\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr),\mathbf{D}_{c,t}:=\operatorname{diag}\bigl(\operatorname{colsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})+\varepsilon\bigr)$, and $\mathbf{P}_{t}:=\mathbf{D}_{r,t}^{-1/2},\mathbf{Q}_{t}:=\mathbf{D}_{c,t}^{-1/2}$.
@@
 This completes the proof. ∎
-E.2 Lemma E.2
-Lemma E.2.
+F.2 Lemma F.2
+Lemma F.2.
 Suppose that $\{E_{t},A_{t}\}$ are nonnegative sequences. Assume $E_{t+1}\leq(1-\alpha_{t+1})E_{t}+A_{t+1}$ where $\alpha_{t}=t^{-p}$, $p\in(0,1]$. Then $\alpha_{t}E_{t}\leq 2(E_{t}-E_{t+1}+A_{t+1}).$
@@
 See [6], Lemma A.3. ∎
-E.3 Lemma E.3
-Lemma E.3.
+F.3 Lemma F.3
+Lemma F.3.
 For Algorithm 1, the accumulated error between the momentum term and the true gradient is bounded for $t\geq 1$:
-$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
+$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq$ $\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]$
+$+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}+(1-\beta_{t+1})^{2}\sigma^{2}.$
 Proof.
@@
 $\|\nabla f(\mathbf{X}_{t})-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}$ $\leq L^{2}\|\mathbf{X}_{t}-\mathbf{X}_{t+1}\|_{F}^{2}$
-$=L^{2}\eta_{t}^{2}\|\mathbf{O}_{t}\|_{F}^{2}$
-$\leq L^{2}\eta_{t}^{2}a^{2}n.$
+$=L^{2}\eta_{t}^{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}$
 Therefore:
 $\mathbb{E}\left[\left\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\right\|_{F}^{2}\right]$ $\leq\beta_{t+1}^{2}(1+\epsilon)\mathbb{E}\left[\left\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\right\|_{F}^{2}\right]$
-$+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$
+$+\beta_{t+1}^{2}\left(1+\frac{1}{\epsilon}\right)L^{2}\eta_{t}^{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}+(1-\beta_{t+1})^{2}\sigma^{2}.$
 Then, by letting $\epsilon:=\frac{1-\beta_{t+1}}{\beta_{t+1}},t\geq 1$, we have
-$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+(1-\beta_{t+1})^{2}\sigma^{2}.$ (6)
-Appendix F Proofs of Proposition 3.4
+$\mathbb{E}\left[\|\mathbf{M}_{t+1}-\nabla f(\mathbf{X}_{t+1})\|_{F}^{2}\right]\leq$ $\beta_{t+1}\mathbb{E}\left[\|\mathbf{M}_{t}-\nabla f(\mathbf{X}_{t})\|_{F}^{2}\right]$ (6)
+$+\frac{\beta_{t+1}^{2}}{1-\beta_{t+1}}L^{2}\eta_{t}^{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}+(1-\beta_{t+1})^{2}\sigma^{2}.$
+Appendix G Proofs of Proposition 3.4
 Proof.
@@
 $\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle=\left\langle\widetilde{\mathbf{A}}_{t}\nabla f(\mathbf{X}_{t})\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle.$
-By Lemma E.1, we have
+By Lemma F.1, we have
 $R_{\varepsilon}^{-1}\mathbf{I}_{m}\preceq\mathbf{P}_{t}\preceq r_{\varepsilon}^{-1}\mathbf{I}_{m},\qquad C_{\varepsilon}^{-1}\mathbf{I}_{n}\preceq\mathbf{Q}_{t}\preceq r_{\varepsilon}^{-1}\mathbf{I}_{n}.$
@@
 $=\|\widetilde{\mathbf{P}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{Q}}_{t}\|_{*}+\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{S}_{t}\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle+\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t}-\mathbf{M}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle.$
-First, by the singular-value inequality
+(i) First, by the singular-value inequality
 $\sigma_{i}(\mathbf{A}\mathbf{X}\mathbf{B})\geq\sigma_{\min}(\mathbf{A})\sigma_{i}(\mathbf{X})\sigma_{\min}(\mathbf{B}),$
@@
 $\|\widetilde{\mathbf{P}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{Q}}_{t}\|_{*}\geq\frac{(\rho_{r}+1)(\rho_{c}+1)}{4\rho_{r}\rho_{c}}\|\mathbf{M}_{t}\|_{*}.$
-Second, using Hölder’s inequality, $\|\mathbf{O}_{t}\|_{2}=1$, and the bounds above,
+(ii) Second, using Hölder’s inequality, $\|\mathbf{O}_{t}\|_{2}=1$, and the bounds above,
 $\left|\left\langle\widetilde{\mathbf{A}}_{t}\mathbf{S}_{t}\widetilde{\mathbf{B}}_{t},\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\right\rangle\right|$ $\leq\|\widetilde{\mathbf{A}}_{t}\mathbf{S}_{t}\widetilde{\mathbf{B}}_{t}\|_{*}\|\widetilde{\mathbf{P}}_{t}\mathbf{O}_{t}\widetilde{\mathbf{Q}}_{t}\|_{2}$
@@
 $\leq\rho_{r}\rho_{c}\sqrt{n}\|\mathbf{S}_{t}\|_{F}.$
-Third, we use the symmetric decomposition
+(iii) Third, we use the symmetric decomposition
 $\widetilde{\mathbf{A}}_{t}\mathbf{M}_{t}\widetilde{\mathbf{B}}_{t}-\mathbf{M}_{t}=(\widetilde{\mathbf{A}}_{t}-\mathbf{I}_{m})\mathbf{M}_{t}+\mathbf{M}_{t}(\widetilde{\mathbf{B}}_{t}-\mathbf{I}_{n})+(\widetilde{\mathbf{A}}_{t}-\mathbf{I}_{m})\mathbf{M}_{t}(\widetilde{\mathbf{B}}_{t}-\mathbf{I}_{n}).$
@@
 $+\frac{aL}{2}\bigl(\chi_{\varepsilon}+\rho_{r}\rho_{c}\sqrt{n}\bigr)^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}n}{2}\sum_{t=1}^{T}\eta_{t}^{2}.$
-By Lemma E.3,
+By Lemma F.3 with $\lambda=0$,
 $\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{(1-h_{t+1})^{2}}{h_{t+1}}L^{2}\eta_{t}^{2}a^{2}n+h_{t+1}^{2}\sigma^{2}.$
@@
 $\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+h_{t+1}^{2}(2\sqrt{2}L^{2}a^{2}n+\sigma^{2}).$
-According to Lemma E.2, by letting $E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},A_{t+1}=h_{t+1}^{2}(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})$, we have
+According to Lemma F.2, by letting $E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},A_{t+1}=h_{t+1}^{2}(2\sqrt{2}L^{2}a^{2}n+\sigma^{2})$, we have
 $h_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\left(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}\right).$
@@
 This completes the proof. ∎
-Appendix G Lemmas for Theorem 3.5
-G.1 Lemma G.1
-Lemma G.1.
-Let $\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\mathbf{R}_{t}:=\mathbf{D}_{r,t}^{-1/2}\mathbf{M}_{t},\mathbf{O}_{t}:=\operatorname{Orth}(\mathbf{R}_{t})$.
-Then, for all $t$,
+Appendix H Lemmas for Theorem 3.5 and Corollary 3.6
+H.1 Lemma H.1
+Lemma H.1.
+Let
+$\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\quad\mathbf{P}_{r,t}:=(\mathbf{D}_{r,t}^{1/2})^{\dagger},\quad\mathbf{R}_{t}:=\mathbf{P}_{r,t}\mathbf{M}_{t},\quad\mathbf{O}_{t}:=\operatorname{Orth}(\mathbf{R}_{t}),$
+where denotes the Moore–Penrose pseudoinverse, so zero rows of $\mathbf{M}_{t}$ remain zero after row normalization, and $\operatorname{Orth}(\mathbf{0}):=\mathbf{0}$. Then, for all $t$,
 $\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F},\qquad\|\mathbf{O}_{t}\|_{F}^{2}=\operatorname{rank}(\mathbf{R}_{t})\leq n.$
 Proof.
-If $\mathbf{M}_{t}=\mathbf{0}$, then $\mathbf{R}_{t}=\mathbf{O}_{t}=\mathbf{0}$ and the claim is immediate. Assume $\mathbf{M}_{t}\neq\mathbf{0}$. Write
+(i) If $\mathbf{M}_{t}=\mathbf{0}$, then $\mathbf{R}_{t}=\mathbf{O}_{t}=\mathbf{0}$ and the claim is immediate.
+(ii) Assume $\mathbf{M}_{t}\neq\mathbf{0}$. Let
 $d_{i,t}:=\|(\mathbf{M}_{t})_{i,:}\|_{2},\qquad\mathbf{D}_{t}:=\operatorname{diag}(d_{1,t},\ldots,d_{m,t}),$
@@
 Finally, if $r_{t}:=\operatorname{rank}(\mathbf{R}_{t})$, then the nonzero singular values of $\mathbf{O}_{t}$ are all equal to $1$, so $\|\mathbf{O}_{t}\|_{F}^{2}=r_{t}\leq n$. ∎
-G.2 Lemma G.2
-Consider the practical row-normalized R variant (4), with Nesterov disabled and $\lambda=0,\varepsilon=0$. Then the update is
-$\mathbf{D}_{r,t}:=\operatorname{diag}\bigl(\operatorname{rowsum}(\mathbf{M}_{t}\odot\mathbf{M}_{t})\bigr),\hat{\mathbf{M}}_{t}:=(\mathbf{D}_{r,t}^{1/2})^{\dagger}\mathbf{M}_{t},\mathbf{O}_{t}:=\mathrm{NS5}(\hat{\mathbf{M}}_{t}),\mathbf{X}_{t+1}=\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t},$
-where $a:=0.2\sqrt{\max\{m,n\}}$.
-Here denotes the Moore–Penrose pseudoinverse, so a zero row remains zero after row normalization, and we adopt the convention $\operatorname{Orth}(\mathbf{0})=\mathbf{0}$.
-Remark G.1.
-Following the practical Muon implementation [20] and Theorem 3.1, we write $\mathrm{NS5}$ as the five-step pre-scaled Newton–Schulz map.
-For any $\mathbf{A}\in\mathbb{R}^{m\times n}$, define $\alpha(\mathbf{A}):=\max\{1,\|\mathbf{A}\|_{F}\},\mathbf{Y}^{(0)}(\mathbf{A}):=\mathbf{A}/\alpha(\mathbf{A}),$ and, for $k=0,1,2,3,4$,
-$\mathbf{Y}^{(k+1)}(\mathbf{A})=p_{\mathrm{ns}}\bigl(\mathbf{Y}^{(k)}(\mathbf{A})\mathbf{Y}^{(k)}(\mathbf{A})^{\top}\bigr)\mathbf{Y}^{(k)}(\mathbf{A}),\qquad p_{\mathrm{ns}}(\mathbf{Z}):=a_{\mathrm{ns}}\mathbf{I}_{m}+b_{\mathrm{ns}}\mathbf{Z}+c_{\mathrm{ns}}\mathbf{Z}^{2}.$
-We then define
+H.2 Lemma H.2
+Lemma H.2.
+Let $\{\hat{\mathbf{M}}_{t}\}_{t=1}^{T}\subset\mathbb{R}^{m\times n}$ be any finite algorithmic trajectory, and set $\operatorname{Orth}(\mathbf{0}):=\mathbf{0}$.
+Consider the five-step pre-scaled Newton–Schulz map with the degree-two Taylor coefficients analyzed in [18, Definition 2].
+For any $\mathbf{A}\in\mathbb{R}^{m\times n}$, define
+$\alpha(\mathbf{A}):=\max\{1,\|\mathbf{A}\|_{F}\},\qquad\mathbf{Y}^{(0)}(\mathbf{A}):=\frac{\mathbf{A}}{\alpha(\mathbf{A})},$
+and, for $k=0,1,2,3,4$,
+$\mathbf{Y}^{(k+1)}(\mathbf{A})=p_{\mathrm{ns}}\bigl(\mathbf{Y}^{(k)}(\mathbf{A})\mathbf{Y}^{(k)}(\mathbf{A})^{\top}\bigr)\mathbf{Y}^{(k)}(\mathbf{A}),$
+where
+$p_{\mathrm{ns}}(\mathbf{Z}):=\frac{15}{8}\mathbf{I}_{m}-\frac{5}{4}\mathbf{Z}+\frac{3}{8}\mathbf{Z}^{2}.$
+We define
 $\mathrm{NS5}(\mathbf{A}):=\mathbf{Y}^{(5)}(\mathbf{A}).$
-Remark G.2.
-Following the trajectory-wise error formalism in [17, Definition 3 and Theorem 2], we work with the NS5 polar-approximation error along the algorithmic trajectory. We define
-$\varepsilon_{\mathrm{ns}}:=\sup_{t\geq 1}\bigl\|\mathrm{NS5}(\hat{\mathbf{M}}_{t})-\operatorname{Orth}(\hat{\mathbf{M}}_{t})\bigr\|_{2}.$
-In subsequent derivations, we assume $\varepsilon_{\mathrm{ns}}<1$.
-The quantity $\varepsilon_{\mathrm{ns}}$ plays the same role as the trajectory-wise polar-approximation error $\varepsilon_{q}$ in [17, Definition 3 and Theorem 2].
-Lemma G.2.
+For each $t$, let $\Pi_{t}$ be the orthogonal projector onto $\operatorname{range}(\hat{\mathbf{M}}_{t})$, and define
+$\delta_{t,0}:=\bigl\|\Pi_{t}-\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})^{\top}\bigr\|_{2},\qquad\delta_{0,T}:=\max_{1\leq t\leq T}\delta_{t,0}.$
+Then $\delta_{0,T}<1$. Moreover, the trajectory-wise NS5 polar-approximation error
+$\varepsilon_{\mathrm{ns},T}:=\max_{1\leq t\leq T}\bigl\|\mathrm{NS5}(\hat{\mathbf{M}}_{t})-\operatorname{Orth}(\hat{\mathbf{M}}_{t})\bigr\|_{2}<1.$
+Proof.
+If $\hat{\mathbf{M}}_{t}=\mathbf{0}$, then $\Pi_{t}=\mathbf{0}$ and $\delta_{t,0}=0$. Otherwise, the pre-scaling by $\alpha(\hat{\mathbf{M}}_{t})=\max\{1,\|\hat{\mathbf{M}}_{t}\|_{F}\}$ gives
+$0<\lambda_{\min}^{+}\bigl(\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})\mathbf{Y}^{(0)}(\hat{\mathbf{M}}_{t})^{\top}\bigr)\leq 1$
+on $\operatorname{range}(\hat{\mathbf{M}}_{t})$, and hence $\delta_{t,0}<1$. Since the horizon $T$ is finite, $\delta_{0,T}=\max_{1\leq t\leq T}\delta_{t,0}<1$.
+The polynomial above is the degree-two Taylor Newton–Schulz polynomial in [18, Definition 2]. Applying [18, Theorem 2] with $\kappa=2$ and $q=5$ yields
+$\varepsilon_{\mathrm{ns},T}\leq 1-\sqrt{1-\delta_{0,T}^{(\kappa+1)^{q}}}=1-\sqrt{1-\delta_{0,T}^{3^{5}}}.$
+Finally, since $1-\sqrt{1-x}\leq x$ for $x\in[0,1]$ and $\delta_{0,T}<1$, we obtain
+$\varepsilon_{\mathrm{ns},T}\leq\delta_{0,T}^{3^{5}}<1.$
+H.3 Lemma H.3
+Lemma H.3.
 Let $\mathbf{A}\in\mathbb{R}^{m\times n}$.
@@
 $\|\mathrm{NS5}(\mathbf{A})-\operatorname{Orth}(\mathbf{A})\|_{2}\leq\varepsilon_{\mathrm{ns}}.$
-Consequently, every diagonal entry $\tilde{\sigma}_{i}$ of $\widetilde{\boldsymbol{\Sigma}}$ satisfies
-$1-\varepsilon_{\mathrm{ns}}\leq\tilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}},\qquad i\in[\operatorname{rank}(\mathbf{A})].$
+Consequently, every diagonal entry $\widetilde{\sigma}_{i}$ of $\widetilde{\boldsymbol{\Sigma}}$ satisfies
+$1-\varepsilon_{\mathrm{ns}}\leq\widetilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}},\qquad i\in[\operatorname{rank}(\mathbf{A})].$
 In particular, whenever $\mathbf{A}=\hat{\mathbf{M}}_{t}\neq\mathbf{0}$, we have $\operatorname{Orth}(\mathrm{NS5}(\mathbf{A}))=\operatorname{Orth}(\mathbf{A}).$
@@
 Hence
-$\max_{i}|\tilde{\sigma}_{i}-1|\leq\varepsilon_{\mathrm{ns}},$
+$\max_{i}|\widetilde{\sigma}_{i}-1|\leq\varepsilon_{\mathrm{ns}},$
 which is equivalent to
-$1-\varepsilon_{\mathrm{ns}}\leq\tilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}}\qquad\text{for all }i.$
+$1-\varepsilon_{\mathrm{ns}}\leq\widetilde{\sigma}_{i}\leq 1+\varepsilon_{\mathrm{ns}}\qquad\text{for all }i.$
 Since $\varepsilon_{\mathrm{ns}}<1$, these diagonal entries are strictly positive. Therefore
@@
 whenever $\mathbf{A}=\hat{\mathbf{M}}_{t}\neq\mathbf{0}$. ∎
-G.3 Lemma G.3
-Lemma G.3.
-Under Lemma G.2, for all $t$,
+H.4 Lemma H.4
+Lemma H.4.
+Under Lemma H.3, for all $t$,
 $\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F},\qquad\|\mathbf{O}_{t}\|_{F}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n.$
@@
 $\hat{\mathbf{M}}_{t}=\mathbf{U}_{t}\boldsymbol{\Sigma}_{t}\mathbf{V}_{t}^{\top},\qquad\mathbf{H}_{t}:=(\hat{\mathbf{M}}_{t}\hat{\mathbf{M}}_{t}^{\top})^{1/2}=\mathbf{U}_{t}\boldsymbol{\Sigma}_{t}\mathbf{U}_{t}^{\top}.$
-By Lemma G.2,
+By Lemma H.3,
 $\mathbf{O}_{t}=\mathbf{U}_{t}\widetilde{\boldsymbol{\Sigma}}_{t}\mathbf{V}_{t}^{\top}$
@@
 $\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle\geq\frac{1-\varepsilon_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}.$
-For the norm bound, let $r_{t}:=\operatorname{rank}(\hat{\mathbf{M}}_{t})$. Again by Lemma G.2, every nonzero singular value of $\mathbf{O}_{t}$ is at most $1+\varepsilon_{\mathrm{ns}}$. Therefore
-$\|\mathbf{O}_{t}\|_{F}^{2}=\sum_{i=1}^{r_{t}}\tilde{\sigma}_{i,t}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}r_{t}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n.$
+For the norm bound, let $r_{t}:=\operatorname{rank}(\hat{\mathbf{M}}_{t})$. Again by Lemma H.3, every nonzero singular value of $\mathbf{O}_{t}$ is at most $1+\varepsilon_{\mathrm{ns}}$. Therefore
+$\|\mathbf{O}_{t}\|_{F}^{2}=\sum_{i=1}^{r_{t}}\widetilde{\sigma}_{i,t}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}r_{t}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n.$
 This completes the proof. ∎
-Appendix H Proofs of Theorem 3.5
+H.5 Lemma H.5
+Lemma H.5.
+Consider the row-normalized update with decoupled weight decay
+$\mathbf{X}_{t+1}=\mathbf{X}_{t}-\eta_{t}(a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}),$
+where $\mathbf{O}_{t}=\mathrm{NS5}(\hat{\mathbf{M}}_{t})$ and Lemma H.4 holds. Let $\gamma:=1-\varepsilon_{\mathrm{ns}},0\leq\rho<\frac{a\gamma}{\sqrt{m}},$ and assume
+$0\leq\lambda_{t}\leq\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}}t^{-1/4}.$ (7)
+Then, along every realization,
+$\lambda_{t}\|\mathbf{X}_{t}\|_{F}\leq\rho,\qquad\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}\leq(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho)\cdot\eta_{t}.$
 Proof.
-Let $\mathbf{S}_{t}:=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\gamma_{\mathrm{ns}}:=1-\varepsilon_{\mathrm{ns}},c_{m,n}^{\mathrm{ns}}:=\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}$.
-By Assumption 3.2 and the descent lemma, we have
-$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})+\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$
-$=f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{La^{2}\eta_{t}^{2}}{2}\|\mathbf{O}_{t}\|_{F}^{2}$
-$\leq f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2},$ (7)
-where the last step uses Lemma G.3.
-Again by Lemma G.3,
-$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle$ $=\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\langle\mathbf{S}_{t},\mathbf{O}_{t}\rangle$
-$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-\|\mathbf{S}_{t}\|_{F}\|\mathbf{O}_{t}\|_{F}$
-$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\|\mathbf{S}_{t}\|_{F}$
-$\geq\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}\|\nabla f(\mathbf{X}_{t})\|_{F}-\Bigl(\frac{\gamma_{\mathrm{ns}}}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)\|\mathbf{S}_{t}\|_{F}.$ (8)
-Substituting (H) into (H),
-$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+ac_{m,n}^{\mathrm{ns}}\eta_{t}\|\mathbf{S}_{t}\|_{F}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2}.$ (9)
+Let $\lambda_{0}:=\frac{\rho}{\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}}.$ Since $\rho<a\gamma_{\mathrm{ns}}/\sqrt{m}\leq a/\sqrt{m}$, we have
+$\lambda_{0}\leq\frac{1}{4(1+\varepsilon_{\mathrm{ns}})\sqrt{mn}}\leq 1.$
+Hence
+$0\leq\eta_{t}\lambda_{t}\leq\lambda_{0}t^{-1}\leq 1.$
+Using Lemma H.4,
+$\|\mathbf{O}_{t}\|_{F}\leq(1+\varepsilon_{\mathrm{ns}})\sqrt{n}.$
+Therefore,
+$\|\mathbf{X}_{t+1}\|_{F}$ $=\|(1-\eta_{t}\lambda_{t})\mathbf{X}_{t}-a\eta_{t}\mathbf{O}_{t}\|_{F}$
+$\leq(1-\eta_{t}\lambda_{t})\|\mathbf{X}_{t}\|_{F}+a\eta_{t}\|\mathbf{O}_{t}\|_{F}$
+$\leq\|\mathbf{X}_{t}\|_{F}+a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\eta_{t}.$
+By induction and $\sum_{s=1}^{t-1}s^{-3/4}\leq 4t^{1/4}$,
+$\|\mathbf{X}_{t}\|_{F}\leq\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}t^{1/4}.$
+Thus
+$\lambda_{t}\|\mathbf{X}_{t}\|_{F}\leq\lambda_{0}t^{-1/4}\Bigl(\|\mathbf{X}_{1}\|_{F}+4a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}t^{1/4}\Bigr)\leq\rho.$
+Finally,
+$\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}$ $=\eta_{t}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}$
+$\leq\eta_{t}\Bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\Bigr)$
+This completes the proof. ∎
+Appendix I Proofs of Theorem 3.5 and Corollary 3.6
+Proof.
+Let $\mathbf{S}_{t}:=\nabla f(\mathbf{X}_{t})-\mathbf{M}_{t},\gamma:=1-\varepsilon_{\mathrm{ns}},$ and define $c_{m,n}^{\mathrm{ns}}:=\frac{\gamma}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}.$
+By (7), we have $\frac{a\gamma}{\sqrt{m}}-\rho>0$. By Assumption 3.2 and the descent lemma,
+$f(\mathbf{X}_{t+1})$ $\leq f(\mathbf{X}_{t})+\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t+1}-\mathbf{X}_{t}\rangle+\frac{L}{2}\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}^{2}$ (8)
+$=f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle-\lambda_{t}\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{X}_{t}\rangle+\frac{L\eta_{t}^{2}}{2}\|a\mathbf{O}_{t}+\lambda_{t}\mathbf{X}_{t}\|_{F}^{2}$
+$\leq f(\mathbf{X}_{t})-a\eta_{t}\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle+\rho\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{L}{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}\eta_{t}^{2},$
+where the last step uses Lemma H.5 and Cauchy’s inequality.
+Again by Lemma H.4, we have
+$\langle\nabla f(\mathbf{X}_{t}),\mathbf{O}_{t}\rangle$ $=\langle\mathbf{M}_{t},\mathbf{O}_{t}\rangle+\langle\mathbf{S}_{t},\mathbf{O}_{t}\rangle$ (9)
+$\geq\frac{\gamma}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-\|\mathbf{S}_{t}\|_{F}\|\mathbf{O}_{t}\|_{F}$
+$\geq\frac{\gamma}{\sqrt{m}}\|\mathbf{M}_{t}\|_{F}-(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\|\mathbf{S}_{t}\|_{F}$
+$\geq\frac{\gamma}{\sqrt{m}}\|\nabla f(\mathbf{X}_{t})\|_{F}-\Bigl(\frac{\gamma}{\sqrt{m}}+(1+\varepsilon_{\mathrm{ns}})\sqrt{n}\Bigr)\|\mathbf{S}_{t}\|_{F}.$
+Substituting (9) into (8), we have
+$f(\mathbf{X}_{t+1})\leq f(\mathbf{X}_{t})-\left(\frac{a\gamma}{\sqrt{m}}-\rho\right)\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+ac_{m,n}^{\mathrm{ns}}\eta_{t}\|\mathbf{S}_{t}\|_{F}+\frac{L}{2}\bigl(a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho\bigr)^{2}\eta_{t}^{2}.$ (10)
 Apply Young’s inequality with parameter $h_{t+1}/L$:
 $ac_{m,n}^{\mathrm{ns}}\eta_{t}\|\mathbf{S}_{t}\|_{F}\leq\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}.$
-Hence
-$f(\mathbf{X}_{t+1})\leq$ $\ f(\mathbf{X}_{t})-\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}$
-$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\eta_{t}^{2}.$
+Let $d_{\rho}^{\mathrm{ns}}:=a(1+\varepsilon_{\mathrm{ns}})\sqrt{n}+\rho$ and $\kappa_{\rho}^{\mathrm{ns}}:=\frac{a\gamma}{\sqrt{m}}-\rho.$ Hence
+$f(\mathbf{X}_{t+1})\leq$ $f(\mathbf{X}_{t})-\kappa_{\rho}^{\mathrm{ns}}\eta_{t}\|\nabla f(\mathbf{X}_{t})\|_{F}+\frac{ah_{t+1}}{2L}\|\mathbf{S}_{t}\|_{F}^{2}$
+$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{L}{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}\eta_{t}^{2}.$
 Taking expectations and summing from $t=1$ to $T$, then using $f(\mathbf{X}_{T+1})\geq f^{\star}$, we obtain
-$\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $\leq f(\mathbf{X}_{1})-f^{\star}+\frac{a}{2L}\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$
-$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{La^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n}{2}\sum_{t=1}^{T}\eta_{t}^{2}.$ (10)
-A direct repetition of the proof of Lemma E.3, using $\|\mathbf{O}_{t}\|_{F}^{2}\leq(1+\varepsilon_{\mathrm{ns}})^{2}n$ from Lemma G.3, gives
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{(1-h_{t+1})^{2}}{h_{t+1}}L^{2}\eta_{t}^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+h_{t+1}^{2}\sigma^{2}.$
+$\kappa_{\rho}^{\mathrm{ns}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}$ $\leq f(\mathbf{X}_{1})-f^{\star}+\frac{a}{2L}\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ (11)
+$\quad+\frac{aL}{2}\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}+\frac{L}{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}\sum_{t=1}^{T}\eta_{t}^{2}.$
+By Lemma F.3, using $\|\mathbf{X}_{t+1}-\mathbf{X}_{t}\|_{F}\leq d_{\rho}^{\mathrm{ns}}\eta_{t}$ from Lemma H.5, we have
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+\frac{(1-h_{t+1})^{2}}{h_{t+1}}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}\eta_{t}^{2}+h_{t+1}^{2}\sigma^{2}.$
 Since $(1-h_{t+1})^{2}\leq 1$ and
@@
 we have
-$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr).$
-Applying Lemma E.2 with
-$E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},\qquad A_{t+1}=h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr),$
+$\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}\leq(1-h_{t+1})\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}+h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr).$
+Applying Lemma F.2 with
+$E_{t}=\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2},\qquad A_{t+1}=h_{t+1}^{2}\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr),$
 we get
@@
 $\sum_{t=1}^{T}h_{t}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}$ $\leq 2\sum_{t=1}^{T}\Bigl(\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}-\mathbb{E}\|\mathbf{S}_{t+1}\|_{F}^{2}+A_{t+1}\Bigr)$
-$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)\sum_{t=1}^{T}\frac{1}{t+1}$
-$\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)(1+\ln T).$
-Hence, since $h_{t+1}\leq h_{t}$,
-$\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}a^{2}(1+\varepsilon_{\mathrm{ns}})^{2}n+\sigma^{2}\Bigr)(1+\ln T).$
+$\leq 2\mathbb{E}\|\mathbf{S}_{1}\|_{F}^{2}+2\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)\sum_{t=1}^{T}\frac{1}{t+1}$
+$\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)(1+\ln T).$
+Hence, since $h_{t+1}\leq h_{t}$, we have
+$\sum_{t=1}^{T}h_{t+1}\mathbb{E}\|\mathbf{S}_{t}\|_{F}^{2}\leq 2\sigma^{2}+2\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)(1+\ln T).$
 Also,
 $\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{h_{t+1}}=\sum_{t=1}^{T}\frac{t^{-3/2}}{(t+1)^{-1/2}}\leq 2(1+\ln T),\qquad\sum_{t=1}^{T}\eta_{t}^{2}=\sum_{t=1}^{T}t^{-3/2}\leq 3.$
-Substituting these bounds into (H),
-$\frac{a\gamma_{\mathrm{ns}}}{\sqrt{m}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}.$
+Substituting the above bounds into (11),
+$\kappa_{\rho}^{\mathrm{ns}}\sum_{t=1}^{T}\eta_{t}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ns}}(1+\ln T)+C_{2,\rho}^{\mathrm{ns}}.$
+where
+$C_{1,\rho}^{\mathrm{ns}}:=\frac{a}{L}\Bigl(2\sqrt{2}L^{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}+\sigma^{2}\Bigr)+aL\bigl(c_{m,n}^{\mathrm{ns}}\bigr)^{2},C_{2,\rho}^{\mathrm{ns}}:=\frac{a\sigma^{2}}{L}+\frac{3L}{2}\bigl(d_{\rho}^{\mathrm{ns}}\bigr)^{2}.$
 Finally, since $\eta_{t}=t^{-3/4}\geq T^{-3/4}$ for all $1\leq t\leq T$,
@@
 Therefore,
-$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1}^{\mathrm{ns}}(1+\ln T)+C_{2}^{\mathrm{ns}}}{\bigl(a\gamma_{\mathrm{ns}}/\sqrt{m}\bigr)T^{1/4}}.$
-This completes the proof. ∎
-Appendix I More Results
+$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\mathbf{X}_{t})\|_{F}\leq\frac{f(\mathbf{X}_{1})-f^{\star}+C_{1,\rho}^{\mathrm{ns}}(1+\ln T)+C_{2,\rho}^{\mathrm{ns}}}{\kappa_{\rho}^{\mathrm{ns}}\cdot T^{1/4}}.$
+Setting $\varepsilon_{\mathrm{ns}}$ to zero yields the corresponding exact-polar result for Theorem 3.5 with decoupled weight decay.
+Setting additionally $\rho=0$ recovers the zero weight decay proof.
+This completes the proof.
+∎
+Appendix J More Results
+[…]
+((a))
+((b))
+Figure 5: Learning-rate sweeps for LLaMA2-130M and LLaMA2-350M trained on C4 for 2.6B and 7.5B tokens, respectively.
 […]
 […]
-(a) 130M
-(b) 350M
-Figure 5: Learning-rate sweeps for LLaMA2-130M and LLaMA2-350M trained on C4 for 2.6B and 7.5B tokens, respectively.
-(a) Step 1%
-(b) Step 10%
-(c) Step 50%
-(d) Step 100%
+((a))
+((b))
+((c))
+((d))
 Figure 6: Singular-value entropy and stable rank of Muon momentum matrices under different normalization schemes during LLaMA2-130M training on C4 for 2.6B tokens.
 RowColNorm consistently improves both metrics over Direct, ColNorm, and RowNorm.
-(a) Step 1%
-(b) Step 10%
-(c) Step 50%
-(d) Step 100%
+Appendix K Experimental Details
+K.1 Pretraining on C4
+[…]
+[…]
+[…]
+$\mathbf{M}_{t}^{\rm Nes}=\beta\mathbf{M}_{t-1}^{\rm Nes}+(1-\beta)\mathbf{G}_{t}+\gamma\beta\bigl(\mathbf{G}_{t}-\mathbf{G}_{t-1}\bigr),\qquad\mathbf{G}_{t-1}:=\nabla f(\mathbf{X}_{t-1};\xi_{t-1}).$
+The MuonEq-Nes convention used in Algorithm 1 corresponds to $\gamma=1-\beta$.
+Indeed, if the EMA buffer is $\mathbf{M}_{t}^{\rm EMA}=\beta\mathbf{M}_{t-1}^{\rm EMA}+(1-\beta)\mathbf{G}_{t}$ and the Nesterov input is $\widetilde{\mathbf{M}}_{t}=\beta\mathbf{M}_{t}^{\rm EMA}+(1-\beta)\mathbf{G}_{t}$, then $\widetilde{\mathbf{M}}_{t}$ satisfies the recurrence above with $\gamma=1-\beta$.
+Thus the reported Muon momentum $\beta=0.95$ corresponds to $\gamma=0.05$; the actual coefficient multiplying $\mathbf{G}_{t}-\mathbf{G}_{t-1}$ is $\gamma\beta=0.0475$.
+((a))
+((b))
+((c))
+((d))
 Figure 7: Layerwise Muon NS5 bias decomposition under different normalization schemes during LLaMA2-130M training on C4 for 2.6B tokens.
@@
 These plots reveal how the choice of normalization shifts the error budget between the finite-step approximation error term and the preconditioning bias term.
-Appendix J Experimental Details
-J.1 Pretraining on C4
+((a))
+((b))
+((c))
+((d))
+Figure 8: Validation perplexity heatmaps on GPT2-small/FineWeb (2.6B tokens) over learning rate and momentum at $K=4$ and $K=5$.
 Table 4: Hyperparameters used for training LLaMA2-130M on C4
-Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq
-Max Learning Rate 5e-3 3e-3 3e-3 3e-3 3e-3
+Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq MuonEq-Nes
+Max Learning Rate 5e-3 3e-3 3e-3 3e-3 3e-3 3e-3
 Warmup Steps 500
@@
 Muon-Momentum 0.95
-Gamma 0.025 0.05
+Gamma 0.025 0.05 0.05
 Table 5: Hyperparameters used for training LLaMA2-350M on C4
-Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq
-Max Learning Rate 1.5e-3 2e-3 1.5e-3 2e-3 2e-3
+Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq MuonEq-Nes
+Max Learning Rate 1.5e-3 2e-3 1.5e-3 2e-3 2e-3 2e-3
 Warmup Steps 500
@@
 Muon-Momentum 0.95
-Gamma 0.025 0.05
-[…]
-[…]
-J.2 Ablation Study Details
+Gamma 0.025 0.05 0.05
+Table 6: Hyperparameters used for training LLaMA2-1B on C4
+Hyper-parameter AdamW MARS-AdamW Muon Muon-Nes MuonEq MuonEq-Nes
+Max Learning Rate {5e-4, 8e-4}
+Warmup Steps 1000
+Batch Size 512
+Maximum Length 4096
+Weight Decay 0.1
+$(\beta_{1},\beta_{2})$ (0.9,0.95)
+Muon-Momentum 0.95
+Gamma 0.025 0.05 0.05
+K.2 Ablation Study Details
 Table 7: Hyperparameters used for training ResNet-18 on CIFAR10
-Hyper-parameter SGD AdamW Muon-Nes MuonEq (RC) MuonEq (C) MuonEq (R)
-Max Learning Rate 5e-2 1e-3 5e-2 5e-2 5e-2 5e-2
+Hyper-parameter SGD AdamW Muon Muon-Nes MuonEq-Nes (RC) MuonEq-Nes (C) MuonEq-Nes (R)
+Max Learning Rate 5e-2 1e-3 5e-2 5e-2 5e-2 5e-2 5e-2
 Warmup Steps 0
@@
 Table 8: Hyperparameters used for training GPT2-small on FineWeb
-Hyper-parameter AdamW Muon-Nes MuonEq (RC) MuonEq (C) MuonEq (R)
-Max Learning Rate 2e-3 2e-3 2e-3 2e-3 2e-3
+Hyper-parameter AdamW AdaMuon Muon+ (colrow) Mousse Muon-Nes MuonEq-Nes (RC) MuonEq-Nes (C) MuonEq-Nes (R)
+Max Learning Rate 2e-3 5e-3 5e-3 2e-3 2e-3 2e-3 2e-3 2e-3
 Warmup Steps 1000
@@
 […]
-We compare the three MuonEq variants RC, R, and C.
+We compare the three MuonEq-Nes variants RC, R, and C.
 RC applies two-sided row/column normalization throughout training, R applies row normalization throughout training, and C applies column normalization throughout training.
@@
 R and C test the two one-sided geometries under the same optimizer configuration.
-Since MuonEq targets hidden matrix weights, R is the more relevant default in our setting, while C is included as the column-sided companion form.
-J.3 A Note on Row/Column Terminology in Prior Work
+Since MuonEq/MuonEq-Nes targets hidden matrix weights, R is the more relevant default in our setting, while C is included as the column-sided companion form.
+K.3 A Note on Row/Column Terminology in Prior Work
 A careful comparison to prior row/column normalization papers requires fixing the matrix convention.
@@
 Their empirical observation that row-wise normalization can be unstable is also traced largely to the LM-head, where row-wise normalization produces extreme gradient values.
-This is not the same setting as MuonEq, which is aimed at hidden matrix weights.
+This is not the same setting as MuonEq/MuonEq-Nes, which is aimed at hidden matrix weights.
 By contrast, [36] use the standard convention $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$, so their ColNorm is genuinely column-wise in the same orientation as ours.
@@
 We therefore interpret our row-sided default choice as directly aligned with [55], compatible with [14] after accounting for layout conventions, and not in tension with [36] once first-layer/input geometry is separated from hidden-layer geometry.
+Appendix L Limitations
+MuonEq is designed as a lightweight pre-orthogonalization correction for matrix-valued hidden weights.
+The row-normalized variant is our default in this setting, but it should not be viewed as a universally optimal choice for every parameter block or architecture.
+Our analysis isolates the core geometry of the pre-balanced orthogonalized update, while broader coverage of implementation variants and training regimes remains an open direction.
+Our convergence analysis for the RC variant is restricted to an auxiliary exact-polar regime: Proposition 3.4 assumes Nesterov momentum is disabled, no decoupled weight decay is used, and the stabilization parameter $\varepsilon$ is sufficiently large.
+Therefore, this RC guarantee does not formally cover the practical RC ablations, which use finite-step Newton–Schulz iterations and the small numerical stabilization $\varepsilon=10^{-8}$.
+Extending the RC theory to this practical regime is left for future work.
+This limitation is specific to the auxiliary RC analysis; the main guarantee for the default R variant is given by Theorem 3.5 and Corollary 3.6.
+Empirically, our large-scale evaluation focuses on LLaMA-style pretraining on C4, complemented by smaller-scale ablations.
+Appendix M Broader Impacts
+This work studies an application-agnostic optimizer that may improve training efficiency for matrix-valued neural network parameters without introducing new data, data collection, or deployment pipelines.
+Its direct societal risks are therefore mainly inherited from the downstream models and datasets on which it is used.
+As with other efficiency improvements, MuonEq could lower the cost of both beneficial and harmful model training, so appropriate downstream safeguards remain necessary.
+Any environmental benefit depends on whether efficiency gains are used to reduce compute for a fixed objective rather than to scale up training.